### v1

# The Power of Graph Sparsification in the Continual Release Model

Abstract
The graph continual release model of differential privacy seeks to produce differentially private solutions to graph problems under a stream of updates where new private solutions are released after each update. Streaming graph algorithms in the non-private literature also produce (approximately) accurate solutions when provided updates in a stream, but they additionally try to achieve two other goals: 1) output vertex or edge subsets as approximate solutions to the problem (not just real-valued estimates) and 2) use space that is sublinear in the number of edges or the number of vertices. Thus far, all previously known edge-differentially private algorithms for graph problems in the continual release setting do not meet the above benchmarks. Instead, they require computing exact graph statistics on the input [FichtenbergerHO21, Song18]. In this paper, we leverage sparsification to address the above shortcomings. Our edge-differentially private algorithms use sublinear space with respect to the number of edges in the graph while some also achieve sublinear space in the number of vertices in the graph. In addition, for most of our problems, we also output differentially private vertex subsets.
We make novel use of assorted sparsification techniques from the non-private streaming and static graph algorithms literature and achieve new results in the sublinear space, continual release setting for a variety of problems including densest subgraph, $k$-core decomposition, maximum matching, and vertex cover. In addition to our edge-differential privacy results, we use graph sparsification based on arboricity to obtain a set of results in the node-differential privacy setting, illustrating a new connection between sparsification and privacy beyond minimizing space. Previous node-differentially private algorithms only provide utility guarantees for the restrictive class of (nearly) bounded degree graphs [FichtenbergerHO21, Song18, jain2024time]. Our connection between sparsification via arboricity and node-privacy potentially opens up the space for additional results in both the static differential privacy and continual release models. We conclude with polynomial additive error lower bounds for edge-privacy in the fully dynamic setting.
1 Introduction
Today’s data is not only massive in size but also rapidly growing. As the world becomes increasingly digitized, the sensitive associations between individuals are becoming a key focus for data analysis. Such a focus also comes with privacy risks and the question of how to safeguard privacy. Differential privacy [DMNS06] is the gold standard for privacy protection, with a rich body of work focusing on static graphs [AU19, blocki2022privately, dinitz2024tight, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, NRS07, RS16, raskhodnikova2016differentially, Upadhyay13]. However, today’s graphs are not static; they are rapidly expanding, with new connections between individuals forming every minute. Streaming graph algorithms, a well-studied area in the non-private setting [assadi2022semi, AG11, EHW16, Esfandiari2018, feigenbaum2005graph, FMU22, HuangP19, henzinger1998computing, muthukrishnan2005data, mcgregor2014graph, mcgregor2015densest, mcgregor2018simple], are designed to handle such massively evolving datasets. In this model, updates to the dataset are received in a stream. Online streaming algorithms continuously release accurate graph statistics after each update [chen2023sublinear, halldorsson2016streaming, cormode2018approximating, GS24]; but these graph statistics may reveal private information.
Our paper focuses on the well-known continual release model [DworkNPR10, CSS11] of differential privacy that promises a strong guarantee for online streaming outputs: the entire vector of all outputs satisfies differential privacy. The two notions of privacy we consider are (1) edge-differential privacy, where we require the output distributions of the algorithm on edge-neighboring stream inputs to be “close”, where edge-neighboring streams differ in one edge update, and, similarly, (2) node-differential privacy, where node-neighboring streams differ in all updates incident to one vertex. Despite the large body of differential privacy works on static graph algorithms, there are few works on graph continual release [FichtenbergerHO21, jain2024time, Song18] (and related graph models [upadhyay2021differentially]). While there has been substantial progress in graph continual release algorithms in recent years, there are some drawbacks to existing algorithms, including:
  • Being unable to release private (vertex subset) solutions in addition to their real-valued function outputs;
  • Requiring exact algorithms for computing graph statistics, thus necessitating linear space usage in the number of edges;
  • Only providing non-trivial node-privacy utility guarantees for the class of (nearly) bounded-degree graphs.
In this paper, we take the first steps to address these short-comings for insertion-only streams in the continual release model: we provide the first graph algorithms that return edge-differentially private vertex subset solutions (in addition to the value of the solutions) and use sublinear space for general graphs. The approximation guarantees of our algorithms nearly match the guarantees for their private static or non-private streaming counterparts. Moreover, as is standard in streaming, all of our results hold for a single pass over the stream. In the more challenging node-privacy setting, we derive algorithms with non-trivial utility guarantees for the class of bounded arboricity graphs, which are a strict super class of bounded degree graphs. [^1] Consider the star graph on $n$ vertices which has maximum degree $n-1$ but arboricity 1.
Resolving these issues makes continual release graph algorithms useful alternatives to their non-private streaming counterparts for many applications. While non-graph private streaming literature has explored sublinear space [ChanLSX12, epasto2023differentially, upadhyay2019sublinear], this topic has not been considered for general graphs in continual release. In most use cases, streaming algorithms are used for datasets that are so large that algorithm designers assume that the entire dataset cannot fit in memory. Moreover, real-life applications often call for vertex subset solutions instead of just the values; for example, advertisers may want the members of the current densest community. Finally, it is preferable to protect the privacy of all of the connections of any user instead of any one connection for problems for which we can achieve non-trivial node-private solutions.
1.1 Sparsification
The main algorithmic technique we use to resolve all three of the aforementioned drawbacks is sparsification, traditionally used in the streaming, static, and dynamic graph algorithms literature. The idea behind sparsification is to create a smaller subgraph of the input graph, using a carefully chosen subset of edges from the stream, which (approximately) preserves the desired property that we want to return for our problem. Such sparsifiers exist for a large number of problems in the non-private literature, including for problems such as densest subgraph, $k$-core decomposition, maximum matching, and vertex cover. In general, graph sparsifiers are problem-specific, where each sparsifier provides guarantees for only one or two closely related problems. We design a systematic method of creating problem-specific differentially private sparsifiers for our problems, which we describe in more detail in the technical overview (Section 2).
Although several projection techniques for graph sparsification exist in the static, one-shot, and sliding window differential privacy literature [BBS13, CZ13, Day16, kalemaj2023node, kasiviswanathan2013analyzing, Upadhyay13, upadhyay2021differentially], such techniques are harder to use in the continual release setting, as the decision to remove a single edge can propagate temporally to many additional changes down the road.
Example 1.1.
Consider the naive sparsifier that removes all edges adjacent to nodes with degree greater than a public bound $\tilde{D}$. On edge-neighboring streams $S\sim S^{\prime}$ node $v$ may have degree $\tilde{D}$ in $S$ and degree $\tilde{D}+1$ in $S^{\prime}$. The naive sparsifier would remove all edges adjacent to $v$ in $S^{\prime}$ but not in $S$. Whereas the previous sensitivity of the two streams is $1$, the sparsified streams now have sensitivity $\tilde{D}$. Every release after the insertion of the last edge to $v$ in $S^{\prime}$ must use sensitivity $\tilde{D}$.
Hence, our techniques must preserve the edge edit distance (Definition 3.5) between neighboring streams. The first time-aware projection algorithms [^2] Defined to be projections that project a stream into one that is (nearly) degree-bounded without significantly increasing the edge edit distance. were designed very recently in [jain2024time] that lead to $\operatorname*{poly}(\log n)$ approximations for node-differential privacy when the input graph stream is guaranteed to have (nearly) bounded degree [FichtenbergerHO21, Song18]. However, it does not guarantee good approximations in more general graphs when $\omega(\operatorname*{poly}(\log(n)))$ vertices have degree $\omega(\operatorname*{poly}(\log(n)))$. We design a number of novel sparsification techniques that are distance-preserving [^3] This means that the sparsified outputs differ by no more than a constant number of edges on edge-neighboring streams and no more than $\alpha$ edges in node-neighboring streams where $\alpha$ is the arboricity. This concept is also referred to as stability in [jain2024time]. and can also be applied in the continual release setting. Our sparsification techniques directly lead to space savings for both edge and node-privacy for a variety of problems including densest subgraph, $k$-core decomposition, maximum matching, and vertex cover; as a bonus, we return vertex subsets for the densest subgraph problem, approximate core numbers for every vertex in core decomposition, and an implicit solution for vertex cover.
Using our sparsification techniques, we also demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. The class of bounded arboricity [^4] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. We broaden the scope of node differentially private algorithms for graph problems in the continual release model beyond bounded degree [FichtenbergerHO21, Song18] and close to bounded degree graph inputs [jain2024time]. We develop algorithms for maximum matching and vertex cover, where, given a public bound $\tilde{\alpha}$ on the arboricity of the graph, our algorithm returns error bounded by $\tilde{\alpha}$ when $\tilde{\alpha}$ upper bounds the maximum (private) arboricity of the stream. Our algorithms are always private, return good approximations when the public bound upper bounds the private graph property we require, and use sublinear space.
Finally, to complement our results on insertion-only streams, we demonstrate the difficulty of obtaining algorithms in fully dynamic streams via improved polynomial error lower bounds for maximum cardinality matching, counting connected components, and triangle counting based on reductions from inner product queries. Similar to the recent continual release lower bounds for counting distinct elements given in [jain2023counting], our lower bounds support the observation from previous works that the fully dynamic setting results in significantly more error [Song18, FichtenbergerHO21] than the insertion-only setting within the continual release model.
1.2 Our Contributions
Our paper gives the following set of results, stated informally here and given in more precise terms in their respective cited sections. We group our results by graph problems. Our upper bounds consider insertion-only streams where edge additions are given in the stream. Our algorithms produce $(\beta,\zeta)$-bicriteria approximations where $\beta$ is the multiplicative factor in the approximation and $\zeta$ is the additive error. Throughout the paper, we assume $n\leq T=O(\operatorname*{poly}(n))$. This is reasonable as $O(n^{2})$ (non-empty) updates are sufficient to have a complete graph. [^5] Our results can be modified to handle streams of length $T=\omega(\operatorname*{poly}(n)).$ We sometimes compare our results with static DP lower bounds because such lower bounds also hold in the insertion-only continual release setting. [^6] Otherwise, if we obtain better error in the continual release setting, we can solve the static problem with better error by inserting all of the edges of the static graph and taking the solution released at the end of the stream. A more in-depth discussion of related works can be found in Appendix A.
Maximum Matching (Section 4)
We give the following results for estimating the size of a maximum matching in bounded arboricity graphs. In this setting, $\tilde{\alpha}$ is a given public bound on the maximum arboricity $\alpha$ of the graph given by the input stream.777 Such public bounds are commonly assumed in non-private streaming literature [CJMM17, mcgregor2018simple]. We also briefly sketch how to remove this assumption in Section E.2 at the cost of more space and a worse approximation guarantee. See Table 1 for a comparison of our algorithms with previous results. We remark that the additive error of our algorithms nearly match the information-theoretic lower bound in bounded arboricity graphs up to logarithmic factors.
Firstly, in the truly sublinear model, where space usage is sublinear in the number of nodes in the graph, we give an edge-DP algorithm for estimating the size of a maximum matching in bounded arboricity graphs.
Theorem 1.2 (See Theorem 4.1).
Given a public bound $\tilde{\alpha}$, we obtain an $\varepsilon$-edge differentially private approximate maximum cardinality matching algorithm that outputs an approximate matching size in $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon}\right)$ space in the insertion-only continual release model. If $\tilde{\alpha}\geq\alpha$, where $\alpha$ is the (private) maximum arboricity of the input graph, then our algorithm returns a $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$-approximation of the size of the maximum matching.
We also derive the first maximum matching algorithms for the stronger model of node-DP beyond (nearly) bounded degree graphs.
Theorem 1.3 (See Theorem 4.13).
[…]
Another version of our algorithm achieves a better multiplicative factor at the cost of slightly more space.
Theorem 1.4 (See Theorem 4.14).
Given a public bound $\tilde{\alpha}$, we obtain an $\varepsilon$-node differentially private approximate maximum matching algorithm in the insertion-only continual release model using $O(n\tilde{\alpha})$ space. When $\tilde{\alpha}\geq\alpha$ is an upper bound on the private arboricity $\alpha$ of the graph, our algorithm gives an $\left((1+\eta),O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\varepsilon}\right)\right)$-approximation of the size of the maximum matching.
Privacy Space Approximation Reference Notes
$\varepsilon$-node DP $\Theta(m)$ $\left(1+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ [FichtenbergerHO21]
$\varepsilon$-node DP $\left(1,\Omega\left(\log(n)\right)\right)$ [FichtenbergerHO21] Lower bound
$\varepsilon$-edge DP $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon}\right)$ $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ Theorem 4.1 Bounded arboricity
$\varepsilon$-node DP $O\left(n\tilde{\alpha}\right)$ $\left(1+\eta,O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\varepsilon}\right)\right)$ Theorem 4.14 Bounded arboricity
$\varepsilon$-node DP $O\left(n\log(\tilde{\alpha})+\frac{\tilde{\alpha}\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon}\right)$ $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\varepsilon}\right)\right)$ Theorem 4.13 Bounded arboricity
Table 1: Maximum matching results: $\tilde{\alpha}$ is a data-independent public bound; for bounds that use $\tilde{\alpha}$, our approximation guarantees hold with high probability when $\tilde{\alpha}$ upper bounds the maximum private arboricity of the stream.
An added advantage of our algorithms is that our node-DP algorithm filters the stream to produce a sparsified vertex set for matching where the maximum degree of any vertex is $\tilde{\alpha}$; the sparsifier may also be useful in the static setting where any edge-DP matching algorithm (that may even output an implicit/explicit solution) can be run on our sparsified graph with sensitivity $\tilde{\alpha}$ and obtain our error bounds.
Vertex Cover (Section 5)
We give the first algorithm for releasing an implicit vertex cover solution in the one-shot, semi-streaming node-DP setting [^9] Here, we define the one-shot semi-streaming setting as the setting where we use $\tilde{O}(n)$ space and the output is given only at the end of the stream. beyond (almost) bounded degree graphs. See Table 2 for a comparison against prior works. Note that any node-DP algorithm for estimating the size of a maximum matching immediately translates to a node-DP algorithm for estimating the size of a minimum vertex cover, but it may be desirable to output an actual implicit solution. The previous solution for vertex cover is only in the static edge-DP model. No prior node-DP solutions (with non-trivial utility bounds) to vertex cover existed, neither in the static or any variant of the private streaming models.
Similar to [gupta2010differentially], our node-DP vertex cover algorithm outputs an implicit solution in the form of a vertex ordering, indicating that an edge is covered by its earlier endpoint. However, our implicit solution differs in that we also output a degree threshold and an edge is first covered by an endpoint with degree exceeding the threshold, should such an endpoint exist, before the edge consults the vertex ordering.
Theorem 1.5 (See Theorem 5.1).
Given a public bound $\tilde{\alpha}$, we obtain an $\varepsilon$-node differentially private approximate implicit solution to the minimum vertex cover problem in the insertion-only semi-streaming setting using $O(n\tilde{\alpha})$ space. When $\tilde{\alpha}\geq\alpha$ is an upper bound on the private arboricity $\alpha$ of the graph, our implicit solution gives a $\left(3+\eta+O\left(\frac{\tilde{\alpha}}{\varepsilon}\right),O\left(\frac{\tilde{\alpha}\log(\nicefrac{{n}}{{\eta}})}{\varepsilon}\right)\right)$-approximation of the vertex cover.
Privacy Space Approximation Reference Note
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(2+\frac{16}{\varepsilon},0\right)$ [gupta2010differentially] Static
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(\Omega\left(\frac{1}{\varepsilon}\right),0\right)$ [gupta2010differentially] Static Lower bound
$\varepsilon$-node DP $O\left(\tilde{\alpha}n\right)$ $\left(3+\eta+O\left(\frac{\tilde{\alpha}}{\varepsilon}\right),O\left(\frac{\tilde{\alpha}\log(n)}{\varepsilon}\right)\right)$ Theorem 5.1 Bounded arboricity One-shot
Table 2: Implicit vertex cover results: $\tilde{\alpha}$ is a data-independent public bound; for bounds that use $\tilde{\alpha}$, our approximation guarantees hold with high probability when $\tilde{\alpha}$ upper bounds the maximum private arboricity of the stream.
Densest Subgraph (Section 6)
We obtain the first $\varepsilon$-edge DP densest subgraph algorithms that return subsets of vertices with near-optimal density in the continual release model while using $o(m)$ space. See Table 3 for a comparison of our results with prior works. We remark that the approximation guarantees nearly match, up to logarithmic terms, the guarantees of their private static or non-private streaming counterparts. In turn, these nearly match the information-theoretic lower bound up to logarithmic factors.
Theorem 1.6 (See Theorem 6.1).
We obtain a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$-approximate $\varepsilon$-edge differentially private densest subgraph algorithm that outputs a subset of vertices in the insertion-only continual release model using $\tilde{O}\left(n/\varepsilon\right)$ space.
At the cost of slightly more additive error and space usage, we can improve the multiplicative error.
Theorem 1.7 (See Theorem 6.2).
We obtain a $\left(1+\eta,O\left(\frac{\log^{5}(n)}{\varepsilon}\right)\right)$-approximate $\varepsilon$-edge differentially private densest subgraph algorithm that outputs a subset of vertices in the insertion-only continual release model using $\tilde{O}\left(n/\varepsilon\right)$ space.
Privacy Space Approximation Reference Note
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(1+\eta,O\left(\frac{\log^{4}(n)}{\varepsilon}\right)\right)$ [DLRSSY22] Static Vertex Subset
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(2,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$ [DLL23] Static Vertex Subset
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ [dinitz2024tight] Static Vertex Subset
$\varepsilon$-edge DP $\left(\beta,\Omega\left(\frac{1}{\beta}\sqrt{\frac{\log(n)}{\varepsilon}}\right)\right)$ [AHS21] Static Lower bound
$\varepsilon$-node DP $\Theta(m)$ $\left(1+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ [FichtenbergerHO21] Density only
$\varepsilon$-edge DP $\tilde{O}\left(\frac{n}{\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$ Theorem 6.1 Vertex Subset
$\varepsilon$-edge DP $\tilde{O}\left(\frac{n}{\varepsilon}\right)$ $\left(1+\eta,O\left(\frac{\log^{5}(n)}{\varepsilon}\right)\right)$ Theorem 6.2 Vertex Subset
Table 3: Densest subgraph results.
$k$-Core Decomposition (Section 7)
We also show the first $\varepsilon$-edge DP algorithm for the $k$-core decomposition problem in the continual release model, that returns approximate $k$-core numbers at every timestep while using $o(m)$ space. See Table 4 for a comparison against previous works. Note that the additive error of our algorithms nearly match the recent lower bound up to logarithmic factors. We also remark that since the $k$-core with largest $k$ is a 2-approximate densest subgraph, lower bounds for the densest subgraph also translate here.
Theorem 1.8 (See Theorem 7.1).
We obtain a $\left(2+\eta,O\left(\frac{\log^{3}(n)}{\varepsilon}\right)\right)$-approximate $\varepsilon$-edge differentially private $k$-core decomposition algorithm that outputs core number estimates in the insertion-only continual release model using $\tilde{O}(n/\varepsilon)$ space.
Privacy Space Approximation Reference Note
$\varepsilon$-edge DP $\Theta(m+n)$ $\left(1,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$ [DLL23] Static
$\varepsilon$-edge DP $\left(\beta,\Omega\left(\frac{\log n}{\varepsilon\beta}\right)\right)$ [HSZ24] Static Lower bound
$\varepsilon$-edge DP $\tilde{O}\left(\frac{n}{\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{3}(n)}{\varepsilon}\right)\right)$ Theorem 7.1
Table 4: $k$-core decomposition results.
Fully Dynamic Lower Bounds (Section 8)
Last but not least, we give improved lower bounds in the fully-dynamic setting in Section 8 for estimating the size of a maximum matching and even simpler problems like counting connected components and counting triangles. See Table 5 for a summary of our results and previous works. Note that previous lower bounds all show a $\Omega(\operatorname*{poly}\log(T))$ lower bound in the additive error while we strengthen these lower bounds to $\Omega(\operatorname*{poly}(T))$, yielding an exponential improvement. Comparing the upper bounds with our new lower bounds for the same problems yield polynomial separations in the error between the insertion-only and fully dynamic settings.
Privacy Approximation Reference Problem
$\varepsilon$-edge DP $\left(1,\Omega\left(\log(T)\right)\right)$ [FichtenbergerHO21] Matching size
$\varepsilon$-edge DP $\left(1,\Omega\left(\log(T)\right)\right)$ [FichtenbergerHO21] Triangle count
$\varepsilon$-edge DP $\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Theorem 8.5 Matching size
$\varepsilon$-edge DP $\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Theorem 8.8 Connected components
$\varepsilon$-edge DP $\left(1,\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right)\right)$ Section 8.4 Triangle count
Table 5: Fully dynamic lower bounds for edge-neighboring streams.
2 Technical Overview
2.1 High-Level Techniques
In this section, we outline our general approach and give an overview of the techniques we use to design our low space algorithms in the continual release setting. We defer the problem-specific details to the next sections. All of our algorithms use $\tilde{O}(n)$ space, i.e., near-linear in the number of vertices in the stream, and some are sublinear in the number of vertices. We present $\varepsilon$-differentially private algorithms in edge-neighboring streams (Definition 3.3), where the streams differ by one edge update, and for node-neighboring streams (Definition 3.4), where the streams differ by all edge updates adjacent to one vertex.
Previous work assumed the ability to compute exact solutions to the problem for each timestep [Song18, FichtenbergerHO21, jain2024time], essentially eliminating the possibility of using sublinear space in general (i.e. dense) graphs as one cannot guarantee exact solutions when using space that is sublinear in the input size. The central theme of our paper is the extension of sparsification techniques from non-private graph algorithms to differentially private algorithms. This extension is non-trivial for numerous reasons that we outline in the following paragraph. We also make a new connection between node-DP algorithms and bounded arboricity sparsifiers.
For our upper bounds, we focus on insertion-only streams, which means that edges can be dynamically added to the graph but not deleted. We also give new results in fully dynamic streams for our lower bounds.
Privacy via Distance-Preserving Sparsifiers
Although sparsification techniques (Section 1.1) are effective for a variety of non-DP graph algorithms, directly applying such methods in the differential privacy setting poses challenges. First, not all non-DP sparsification techniques maintain the same sensitivity bounds as the original stream (see e.g. Example 1.1). We use problem-specific techniques to ensure that the sparsified outputs differ by no more than a constant number of edges on edge-neighboring streams. [^10] This is also defined as stability in [jain2024time], i.e. edge-to-edge and node-to-edge stability.
For our node-privacy results, we demonstrate sparsification algorithms that, in fact, shrink the distance between our sparsified outputs, such that, the new distance is upper bounded by some public bound $\tilde{\alpha}$. While the original streams can differ by as many as $n-1$ edges, our sparsified outputs differ by at most $O(\tilde{\alpha})$ edges. When $\tilde{\alpha}$ upper bounds a private parameter known as the arboricity of the graph, we also obtain strong approximation guarantees for our problems in terms of the arboricity of the graph. This allows us to obtain better error in the node-privacy setting than algorithms that simply execute edge-private algorithms with noise calibrated using group privacy (Theorem 3.12). Another important characteristic of sparsifying node-neighboring streams in this fashion is that by minimizing the number of edges that differ, we can apply any edge-private algorithm on the product of the sparsification, which means our sparsified products can also be used in the static setting. In fact, we combine two of our streaming algorithms to obtain even better space guarantees in the node-privacy setting. This process is morally similar to the bounded-degree projection techniques [BBS13, Day16, jain2024time] in privacy literature in which an arbitrary graph is “projected” to a bounded degree graph, except for one main difference: sparsifiers in (non-private) graph algorithms literature are problem-specific and often it is very difficult to obtain a sparsifier (using small space) that works for many different graph problems.
Utility under Privacy
One crucial roadblock with respect to differentially private algorithms is that we must deal with additive error in our approximate solutions on the sparsified graph. As most graph algorithms that use sparsification require scaling up a solution in the sparsified graph into a solution in the original graph and most differentially private mechanisms have additive error, we must show that scaling up the approximate solution does not blow up the additive error.
In addition to our sparsification algorithms, another important building block in our algorithms is the use of exact or approximate algorithms for obtaining a differentially private solution in the static setting. Depending on the problem, we either apply these algorithm(s) in a black-box or white-box manner to our sparsified subgraph at carefully chosen points in the stream. We cannot afford to compute and release a new private solution at each point in the stream as that would result in privacy loss proportional to the length of the stream, $T$. Hence, another key theme in our algorithms is the concept of lazy updates where we do not update our private outputs until necessary. In this way, we update our differentially private output only a small, $O(\log T)$ times, resulting in smaller loss of privacy and hence smaller error.
2.2 Problem Specifics
We now apply the general strategies outlined above to a variety of problems in the next sections; each of these problems uses a unique sampling scheme as sparsifiers in graph literature are often problem-dependent. We prove the privacy of every algorithm as well as the utility of each of our algorithms within their individual sections. We summarize on a high level our technical contributions for each of the following problems.
Maximum Cardinality Matching Size (Sections 4.1 and 4.3Table 1)
Given a stream of updates $S$, the maximum cardinality matching problem asks for a real-valued integer at each timestep $t$ equal to the size of the maximum matching in $G_{t}$. We give two algorithms for this problem in this paper: one with edge-DP guarantees and one with node-DP guarantees. Both of our algorithms use a public bound $\tilde{\alpha}$. If this public bound upper bounds the private maximum arboricity of the input stream, then our algorithms give concentration bounds for the approximation in terms of $\tilde{\alpha}$.
Our first algorithm is a truly sublinear space, $\varepsilon$-edge DP algorithm in the continual release model, which uses space that is sublinear in the number of vertices in the graph. Given a public bound $\tilde{\alpha}$, if $\tilde{\alpha}$ upper bounds the private maximum arboricity of the input stream, then we obtain a $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}n}{\varepsilon}\right)\right)$-approximation of the matching size. For this algorithm, we take inspiration from the sublinear space non-private $((1+\eta)(2+\alpha))$-approximate maximum matching algorithm of [mcgregor2018simple] which also requires an upper bound $\alpha$ on the maximum arboricity of the stream; their approximation guarantees are given in terms of this upper bound.
The key observation that [mcgregor2018simple] makes is that maintaining $\Omega(\log n/\eta^{2})$ edge samples, sampled uniformly at random in the stream, and counting the number of edges in the sample which are adjacent to at most $\alpha+1$ edges occurring later in the stream is sufficient for providing a good estimate of the maximum matching size in terms of the arboricity upper bound. Edges which are adjacent to more than $\alpha+1$ edges that occur later in the stream are discarded from the sample. Unfortunately, in edge-neighboring streams, we cannot add noise to the counts of every edge in the sample since the sensitivity of such counts is $\Omega(\tilde{\alpha})$ (our public bound). Instead, we show through an intricate charging argument that the sensitivity of the size of the sample is bounded by $2$. Since the sensitivity of the size of the sample is bounded by $2$, we can use the sparse vector technique (SVT) to determine when to release a new estimate when the estimate changes by a large amount. Using SVT results in noisy sample sizes which requires a new analysis for the utility bounds. We can further remove the assumption of $\tilde{\alpha}$ using $\tilde{O}(n)$ space via the parallel guessing trick in Section E.2.
Our second algorithm is an $\varepsilon$-node DP algorithm in the continual release model which uses $O\left(n\tilde{\alpha}\right)$ space, given the public bound $\tilde{\alpha}$. When the public bound $\tilde{\alpha}$ upper bounds the private maximum arboricity of the input graph, then the algorithm is guaranteed to return a $O\left(1+\eta,O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\varepsilon}\right)\right)$-approximate solution. In this algorithm, we modify the static bounded arboricity sparsifier of [solomon2021local] to the streaming setting. Since we have a public bound $\tilde{\alpha}$ that could be much greater than the actual $\alpha$ of the graph, we provide a simple generalization of their bounds to any $\tilde{\alpha}\geq\alpha$ (beyond the precise bound of $\alpha$). Our streaming sparsifier consists of the first $O(\tilde{\alpha})$ edges adjacent to each vertex in the stream; we show that such a procedure has small sensitivity and is distance-preserving.
The main takeaway from this sparsification technique, similar to the bounded-degree projections in previous work [BBS13, Day16, jain2024time], is that we have now reduced the degree of each vertex to $O(\tilde{\alpha})$. This means that the distance between our neighboring sparsifiers is $O(\tilde{\alpha})$ compared to the distance of $n-1$ of the original streams. Finally, we use an exact matching algorithm on the sparsified graph to produce the maximum matching in the sparsifier. Since, by our sparsifier guarantees, the maximum matching size has sensitivity at most $O(\tilde{\alpha})$, we can combine this with an SVT to determine when to release a new approximation of our maximum matching size.
Alternatively, we can merge our edge-private streaming algorithm with our node-private sparsifier to obtain a node-DP algorithm that uses $\tilde{O}(n+\tilde{\alpha}/\varepsilon)$ space and has (almost) the same approximation guarantee as our edge-private algorithm. To merge our algorithms, we run our edge-DP algorithm in parallel with our node-DP algorithm by performing the sampling only on the edges that are part of the node-DP sparsifier. Such a merge removes the need for the use of the SVT on top of an exact matching algorithm and reduces the space used by a factor of $\tilde{\alpha}$.
Implicit Vertex Cover (Section 5Table 2)
In the one-shot, semi-streaming setting, we process edges in a stream using $\tilde{O}(n)$ space but only output an implicit solution at the end of the stream. The minimum vertex cover problem seeks to return a minimum sized set of vertices where every edge has an endpoint in the set. [gupta2010differentially] introduced the setting of implicit solutions for vertex cover where they release an $\varepsilon$-edge differentially private total ordering of the vertices in the static setting and each edge can determine which vertex covers it using the ordering. In our setting, we show that a variant of our matching sparsifier can be used as a vertex cover sparsifier since it contains the vertex cover sparsifier given in [solomon2021local]. Then, we make a simple adaptation of the implicit vertex cover algorithm of [gupta2010differentially] to the bounded-degree node-DP setting. Finally, we show that running the adapted algorithm on the sparsified stream yields the first one-shot, semi-streaming $\varepsilon$-node DP implicit vertex cover algorithm, where each edge can determine which of their endpoints covers it by using the public ordering and the degrees of their endpoints. This gives a $\left(3+\eta+O\left(\frac{\tilde{\alpha}}{\varepsilon}\right),O\left(\frac{\tilde{\alpha}\log(\nicefrac{{n}}{{\eta}})}{\varepsilon}\right)\right)$-approximation when $\tilde{\alpha}\geq\alpha$.
Densest Subgraph (Section 6, Table 3)
Given an edge-neighboring stream and timestamp $t$, the densest subgraph in $G_{t}$ is a subset of vertices $V_{\operatorname{OPT}}$ which maximizes the density of the induced subgraph, $V_{\operatorname{OPT}}=\arg\max_{V^{\prime}\subseteq V}\left(\frac{|E_{t}(V^{\prime})|}{|V^{\prime}|}\right)$. We return a subset of vertices $V_{approx}$ whose induced subgraph gives a $\left(1+\eta,O\left(\frac{\log^{5}(n)}{\varepsilon}\right)\right)$-approximation of the density of the densest subgraph (see Theorem 1.7 and Table 3) in $\tilde{O}(n/\varepsilon)$ space. Our algorithm calls the static $\varepsilon$-DP DSG algorithm of [DLRSSY22] in a black-box manner. Using a different static $\varepsilon$-edge DP DSG algorithm [DLL23], we can obtain a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\right)$-approximation while saving a polylogarithmic factor in space.
We use a modified version of the non-private $(1+\eta)$-approximate sparsification algorithms of [EHW16, mcgregor2015densest]. Previous works in the continual release model only released the value of the densest subgraph and not the vertex subset [FichtenbergerHO21, jain2024time]. Since the density of the densest subgraph has sensitivity $1$, previous works use the sparse vector technique [dwork2009complexity, roth2010interactive, hardt2010multiplicative, lyu2017understanding] and private prefix sums to release these values after adding appropriate Laplace noise. In our work, we also release a private set of vertices at timestep $t$ whose induced subgraph is an approximation of the densest subgraph in $G_{t}$. We use the sparse vector technique to determine when to release a new private set of vertices, and we use one of the many existing differentially private densest subgraph algorithms [AHS21, DLL23, DLRSSY22, NV21, dinitz2024tight] to return a vertex subset in the sparsified graph. However, there are several challenges in simply returning the subgraph obtained from these static algorithms. First, compared to the non-private setting, we cannot obtain an exact densest subgraph in the sparsified graph. Hence, we must ensure that neither the additive nor multiplicative error blows up when the subgraph obtained in the sparsified graph is scaled up to the original graph. To solve this issue, we ensure that the first subgraph we release has size (approximately) at least the additive error returned by the private static algorithms. Furthermore, we prove concentration bounds on our approximation in the original graph using a carefully constructed Chernoff bound argument.
We sparsify the stream by sampling each edge in the stream with probability $p_{t}$. Earlier on in the stream, our probability of sampling must be set high enough in order to sample enough edges to ensure our concentration bounds. However, as we see more edges, we must reduce the probability of sampling to ensure our sublinear space bounds. In edge-neighboring streams, such reductions in probability may not occur at the same time, thus leading to sparsified graphs that are not distance-preserving. Thus, we must also use the sparse vector technique to determine when to reduce our probability of sampling an edge. We ensure privacy based on a careful conditioning on these probabilities. Unlike the static setting where the privacy analysis of algorithms that rely on an SVT instance can apply simple composition, we must argue directly by definition since we must output a solution after every update but only lose privacy on timestamps where the SVT answers “above”.
$k$-Core Decomposition (Section 7Table 4)
Given a stream of updates $S$, the $k$-core decomposition problem returns a number for each vertex after every update indicating the $k$-core with the maximum value $k$ in which it is contained. A $k$-core is defined as a maximal set of vertices $V^{\prime}\subseteq V$ where the degree of every vertex in the induced graph consisting of $V^{\prime}$ is at least $k$. We return approximate core numbers for every vertex at every timestep $t$ that are $\left(2+\eta,O\left(\frac{\log^{3}(n)}{\varepsilon}\right)\right)$-approximations of the true core number (see Theorem 1.8 and Table 4) in $\tilde{O}(n/\varepsilon)$ space.
Our differentially private algorithm combines a non-private $(1+\eta)$-approximate sampling algorithm inspired by [Esfandiari2018] with the private static $\left(2+\eta,O\left(\frac{\operatorname*{poly}(\log n)}{\varepsilon}\right)\right)$-approximate level data structure algorithm of [DLRSSY22, DLL23] in a white-box manner. Unlike in the densest subgraph setting, we can no longer use the private algorithm in a black-box manner since we need to return the value of every vertex at each timestep. Every vertex may change its core number many times throughout the duration of the stream. Thus, using any static $k$-core decomposition algorithm in a black-box manner incurs a privacy loss of a factor of $n$ via composition.
In our algorithm, we use a version of the multidimensional sparse vector technique given in [DLL23] adapted to the continual release setting. In the static setting, [DLL23] uses the multidimensional sparse vector technique to determine when a vertex stops moving up levels; the level in which a vertex resides directly corresponds with an approximation of the core number of the vertex. In the static setting, once a vertex stops moving up levels, it will immediately output a new approximation; hence, each vertex fails the SVT check at most once. The main difference between our usage of the multidimensional sparse vector technique is that each vertex must output a value at every timestep. This means that the multidimensional sparse vector technique must be adapted to allow for conditioning on previous outputs.
As we are presenting a sparsified version of the level data structures used in previous works, we prove our approximation bounds via a new analysis that shows that vertices on a given level in our sparsified structure will be on (approximately) the same level in the non-sparsified structure, leading to our final bounds.
Fully Dynamic Lower Bounds (Section 8Table 5)
We show polynomial lower bounds for the maximum matching, triangle counting, and connected components problems in fully dynamic streams. Our lower bounds reduce each of these graph problems via unique graph constructions to answer inner product queries [Dinur2003Revealing, De2012Lower, Dwork2007Price, MMNW11]. The known lower bounds for inner product queries then apply to our problems. We encode the secret dataset given by the inner product query problem via $\Theta(n)$ insertions to construct our graph. Then, we map the value of the answers to each of our graph problems to the original inner product query via the inclusion-exclusion principle. Finally, we delete the insertion through an edge deletion and repeat the process for the next query. We remark that this technique is quite general and likely extends to many other natural graph problems. Roughly speaking, the only problem-specific requisite is that we can encode the bits of a secret dataset using the problem structure, and that we can easily flip the bits by adding and deleting edges. For example, we can encode each bit of a secret dataset as a (non-)edge in an induced matching so that flipping a bit equates to (adding/)deleting the (non-)edge.
3 Preliminaries
We introduce the notation we use in this paper as well as some standard privacy and graph theoretic definitions. The familiar reader can feel free to skip this section.
3.1 Setting & Notation
A graph $G=(V,E)$ consists of a set of vertices $V$ and a set of edges $E$ where edge $\{u,v\}\in E$ if and only if there is an edge between $u\in V$ and $v\in V$. We write $n:=|V|$ and $m:=|E|$.
We consider the insertion-only setting within the continual release model, where we begin with an empty graph $G_{0}=(V,E_{0})$ where $E_{0}=\varnothing$ and edge updates in the form of $\{e_{t},\perp\}$ arrive in a stream at every timestep $t\in[T]$. We are required to release an output for the problem of interest after every edge update. Our goal is to obtain algorithms that achieve sublinear space, either in the number of edges $\tilde{o}(m)$ (note that $T\geq m$ as we allow for empty updates), or the number of vertices $\tilde{o}(n)$. We assume $T\leq n^{c}$ for some absolute constant $c\geq 1$ to simplify our presentation. If there are no empty edge updates then it is sufficient to consider $c=2$.
Throughout the paper, we write $\eta\in(0,1]$ to denote a fixed multiplicative approximation parameter. We assume $\eta$ is a fixed constant that is ignored under asymptotic notation. Constants used throughout the paper may depend on $\eta$.
We use $\mathcal{M}$ to denote a mechanism, $M(\cdot)$ to denote the size of a maximum matching, and $\mu$ to denote the mean of a random variable.
3.2 Approximation Algorithms
We consider optimization problems in both the minimization and maximization setting. Let $\operatorname{OPT}\in\mathbb{R}$ be the optimum value for the problem. For a minimization problem, a $(\beta,\zeta)$-approximate algorithm outputs a solution with cost at most $\beta\cdot\operatorname{OPT}+\zeta$. For a maximization problem, $(\beta,\zeta)$-approximate algorithm outputs a solution of value at least $\frac{1}{\beta}\cdot\operatorname{OPT}-\zeta$. For estimation problems, a $(\beta,\zeta)$-approximate algorithm outputs an estimate $\tilde{R}\in\mathbb{R}$ of some quantity $R\in\mathbb{R}$ where $R-\zeta\leq\tilde{R}\leq\beta\cdot R+\zeta$ (one-sided multiplicative error). As a shorthand, we write $\beta$-approximation algorithm to indicate a $(\beta,0)$-approximation algorithm.
3.3 Concentration Inequalities
Below, we make use of the multiplicative Chernoff bound.
Theorem 3.1 (Multiplicative Chernoff Bound; Theorems 4.4, 4.5 in [mitzenmacher2015probability]).
Let $X=\sum_{i=1}^{n}X_{i}$ where each $X_{i}$ is a Bernoulli variable which takes value $1$ with probability $p_{i}$ and value 00 with probability $1-p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. Then, it holds:
  1. 1.
    Upper Tail: $\mathbf{P}(X\geq(1+\eta)\cdot\mu)\leq\exp\left(-\frac{\eta^{2}\mu}{2+\eta}\right)$ for all $\eta>0$;
  2. 2.
    Lower Tail: $\mathbf{P}(X\leq(1-\eta)\cdot\mu)\leq\exp\left(-\frac{\eta^{2}\mu}{3}\right)$ for all $10<\eta<$ 1.
3.4 Neighboring Streams
We use the phrasing of the below definitions as given in [jain2024time].
Definition 3.2 (Graph Stream [jain2024time]).
In the continual release model, a graph stream $S\in\mathcal{S}^{T}$ of length $T$ is a $T$-element vector where the $i$-th element is an edge update $u_{i}=\{v,w,insert\}$ indicating an edge insertion of edge $\{v,w\}$, $u_{i}=\{v,w,delete\}$ indicating an edge deletion of edge $\{v,w\}$, or $\bot$ (an empty operation).
We use $G_{t}$ and $E_{t}$ to denote the graph induced by the set of set of updates in the stream $S$ up to and including update $t$. Now, we define neighboring streams as follows. Intuitively, two graph streams are node-neighbors if one can be obtained from the other by removing all the edge updates incident to a single vertex. (For graph streams, the edges adjacent to the vertex are spread over many timesteps). Similarly, two graph streams are edge neighbors if one can be obtained from the other by removing one edge update (replacing the edge update by an empty update in a single timestep).
Definition 3.3 (Edge Neighboring Streams).
Two streams of edge updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are edge-neighboring if there exists exactly one timestamp $t^{*}\in[T]$ where $u_{t^{*}}\neq u^{\prime}_{t^{*}}$ and for all $t\neq t^{*}\in[T]$, it holds that $u_{t}=u^{\prime}_{t}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t^{*}}$ and $u^{\prime}_{t^{*}}$ that $u^{\prime}_{t^{*}}=\bot$ and $u_{t^{*}}=\pm e_{t^{*}}$ is an edge insertion or deletion.
Definition 3.4 (Node Neighboring Streams).
Two streams of edge updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are node-neighboring if there exists exactly one vertex $v^{*}\in V$ where for all $t\in[T]$, $u_{t}\neq u^{\prime}_{t}$ only if $u_{t}$ or $u_{t}^{\prime}$ is an edge addition or deletion of an edge adjacent to $v^{*}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t^{*}}\neq u^{\prime}_{t^{*}}$ that $u^{\prime}_{t^{*}}=\bot$ and $u_{t^{*}}=\pm e_{t}$ is an edge insertion or deletion of an edge adjacent to $v^{*}$.
We make a short note of comparison between this definition to the definition of node-neighboring graphs in the static setting. While node-neighboring graphs in the static setting are defined in terms of the difference of all edges adjacent to any one node, node-neighboring streams are defined in terms of all updates adjacent to any one node. This means that for fully dynamic streams where both edge insertions and deletions can occur, a particular edge can be inserted and deleted many times; this entire sequence of insertions and deletions on the same edge could differ between node-neighboring streams. For insertion-only streams, such a distinction is moot since any edge can be associated with at most one update. Note that our edge-neighboring streams are defined in terms of any one update event so such a distinction between insertion-only and fully dynamic streams also does not exist for edge-neighboring streams.
We now formalize the concepts of edge edit distance between streams and the concept of distance-preserving.
Definition 3.5 (Edge/Node Edit Distance).
Given two streams $S,S^{\prime}\in\mathcal{S}^{T}$, the edge edit distance between the two streams is the shortest chain of graph streams $S_{0},S_{1},\dots,S_{d}$ where $S_{0}=S$ and $S_{d}=S^{\prime}$ where every adjacent pair of streams in the chain are edge-neighboring. The edge edit distance is $d$, the length of the chain.
The node edit distance between two streams is the length of the shortest chain where every adjacent pair of streams in the chain are node-neighboring.
An algorithm that takes as input a stream $S$ and outputs a chosen set of updates from the stream is distance-preserving if on edge-neighboring streams $S$ and $S^{\prime}$, there exists a coupling between the randomness used in the algorithm on the inputs such that the edge distance between the output streams is $O(1)$. We say such an algorithm is distance-preserving on node-neighboring streams if given two node-neighboring streams $S$ and $S^{\prime}$, there exists a coupling such that the output streams have edge edit distance $O(\tilde{\alpha})$ where $\tilde{\alpha}$ is some public parameter passed into the algorithm.
3.5 Differential Privacy Tools
Here, we define the privacy tools commonly used in differential privacy in terms of the continual release model.
Definition 3.6 (Edge Differential Privacy).
Let $\varepsilon,\delta\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $(\varepsilon,\delta)$-edge differentially private (DP) if for any pair of edge-neighboring graph streams $S,S^{\prime}$ that differ by 1 edge update and for every $T$-sized vector of outcomes $Y\subseteq\operatorname{Range}(\mathcal{A})$,
$\mathbf{P}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathbf{P}\left[\mathcal{A}(S^{\prime})\in Y\right]+\delta.$
When $\delta=0$, we say that $\mathcal{A}$ is $\varepsilon$-edge DP.
Definition 3.7 (Node Differential Privacy).
Let $\varepsilon,\delta\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $(\varepsilon,\delta)$-node differentially private (DP) if for any pair of node-neighboring graph streams $S,S^{\prime}$, that differ in all updates adjacent to one vertex, and for every $T$-sized vector of outcomes $Y\subseteq\operatorname{Range}(\mathcal{A})$,
$\mathbf{P}\left[\mathcal{A}(S)\in Y\right]\leq e^{\varepsilon}\cdot\mathbf{P}\left[\mathcal{A}(S^{\prime})\in Y\right]+\delta.$
When $\delta=0$, we say that $\mathcal{A}$ is $\varepsilon$-node DP.
Throughout the paper, we use some standard privacy mechanisms as building blocks (see [dwork2014algorithmic] for a reference).
Definition 3.8 (Global sensitivity).
The global sensitivity of a function $f:{\mathcal{D}}\to\mathbb{R}^{d}$ is defined by
$\Delta_{f}=\max_{D,D^{\prime}\in{\mathcal{D}},D\sim D^{\prime}}\|f(D)-f(D^{\prime})\|_{1}.$
where $D\sim D^{\prime}$ are neighboring datasets and differ by an element.
Definition 3.9 (Laplace Distribution).
We say a random variable $X$ is drawn from a Laplace distribution with mean $\mu$ and scale $b>0$ if the probability density function of $X$ at $x$ is $\frac{1}{2b}\exp\left(-\frac{|x-\mu|}{b}\right)$. We use the notation $X\sim{\textsf{Lap}}(b)$ to denote that $X$ is drawn from the Laplace distribution with scale $b$ and mean $\mu=0$.
The Laplace mechanism for $f:X\to\mathbb{R}$ with global sensitivity $\sigma$ adds Laplace noise to the output of $f$ with scale $b=\nicefrac{{\sigma}}{{\varepsilon}}$ before releasing.
Proposition 3.10 ([dwork2014algorithmic]).
The Laplace mechanism is $\varepsilon$-DP.
Theorem 3.11 (Adaptive Composition; [DMNS06, DL09, DRV10]).
A sequence of DP algorithms, $(\mathcal{A}_{1},\dots,\mathcal{A}_{k})$, with privacy parameters $(\varepsilon_{1},\dots,\varepsilon_{k})$ form at worst an $\left(\varepsilon_{1}+\cdots+\varepsilon_{k}\right)$-DP algorithm under adaptive composition (where the adversary can adaptively select algorithms after seeing the output of previous algorithms).
Theorem 3.12 (Group Privacy; Theorem 2.2 in [dwork2014algorithmic]).
Given an $\varepsilon$-edge (node) differentially private algorithm, $\mathcal{A}$, for all pairs of input streams $S$ and $S^{\prime}$, it holds that for all possible outcomes $Y\in\operatorname{Range}(\mathcal{A})$,
$e^{-k\varepsilon}\leq\frac{\mathbf{P}[\mathcal{A}(S^{\prime})\in Y]}{\mathbf{P}[\mathcal{A}(S)\in Y]}\leq e^{k\varepsilon}$
where $k$ is the edge (node) edit distance between $S$ and $S^{\prime}$.
3.5.1 Sparse Vector Technique
Below, we define the sparse vector technique and give its privacy and approximation guarantees. The sparse vector technique is used to answer above threshold queries where an above threshold query checks whether the output of a function that operates on an input graph $G$ exceeds a threshold $T$.
We use the variant introduced by [lyu2017understanding] and used by [FichtenbergerHO21]. Let $D$ be an arbitrary (graph) dataset, $(f_{t},\tau_{t})$ a sequence of (possibly adaptive) query-threshold pairs, $\Delta$ an upper bound on the maximum sensitivity of all queries $f_{t}$, and an upper bound $c$ on the maximum number of queries to be answered “above”. Typically, the AboveThreshold algorithm stops running at the first instance of the input exceeding the threshold, but we use the variant where the input can exceed the threshold at most $c$ times where $c$ is a parameter passed into the function.
Throughout this paper, we use the class $\textsc{SVT}(\varepsilon,\Delta,c)$ (Algorithm 3.1) where $\varepsilon$ is our privacy parameter, $\Delta$ is an upper bound on the maximum sensitivity of incoming queries, and $c$ is the maximum number of “above” queries we can make. The class provides a $\textsc{ProcessQuery}(query,threshold)$ function where $query$ is the query to SVT and $threshold$ is the threshold that we wish to check whether the query exceeds.
Theorem 3.13 (Theorem 2 in [lyu2017understanding]).
Algorithm 3.1 is $\varepsilon$-DP.
We remark that the version of SVT we employ (Algorithm 3.1) does not require us to resample the noise for the thresholds (Algorithm 3.1) after each query but we do need to resample the noise (Algorithm 3.1) for the queries after each query.
1 Input: privacy budget $\varepsilon$, upper bound on query sensitivity $\Delta$, maximum allowed “above” answers $c$
2 Class SVT($\varepsilon,\Delta,c$)
3       $\varepsilon_{1},\varepsilon_{2}\leftarrow\nicefrac{{\varepsilon}}{{2}}$
4       $\rho\leftarrow{\textsf{Lap}}(\nicefrac{{\Delta}}{{\varepsilon_{1}}})$
5       $\operatorname{count}\leftarrow 0$
6
7       Function ProcessQuery($f_{t}(D),\tau_{t}$)
8             if $\operatorname{count}>c$ then
9                   return “abort”
10
11            if $f_{t}(D)+{\textsf{Lap}}({2c\Delta}/{\varepsilon_{2}})\geq\tau_{t}+\rho$ then
12                   return “above”
13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
14
15            else
16                   return “below”
17
18
19
Algorithm 3.1 Sparse Vector Technique
4 Maximum Matching and Arboricity Sparsification
We begin this section in Section 4.1 with an edge-DP algorithm for maximum matching that uses truly sublinear space. We also briefly sketch how to remove the assumption on $\tilde{\alpha}$ in Section E.2 at the cost of more space and a worse approximation guarantee.
Then, we derive space-efficient algorithms that satisfy the stronger node-differential privacy guarantee (cf. Definition 3.7) for graphs of bounded arboricity $\alpha$. The key idea is that a judicious choice of a sparsification algorithm reduces both the space complexity and the edge edit distance between node-neighboring graphs to some factor $\Lambda=O(\alpha)$. Such sparsification is useful since, in the worst case, the edge edit distance pre-sparsification can be $\Omega(n)$. Then, it suffices to run any edge-DP algorithm with privacy parameter $\nicefrac{{\varepsilon}}{{\Lambda}}$ after the sparsification to achieve the same privacy guarantee but reduced error.
By adapting techniques from [solomon2021local], we derive such sparsifiers for the maximum matching problem in Section 4.2 (and the minimum vertex cover problem in Section 5). Then, we design a node-DP maximum matching algorithm in the continual release setting in Section 4.3. Later, in Section 5, we explain how to release a node-DP implicit vertex cover in the semi-streaming setting.
4.1 Edge-DP Maximum Cardinality Matching
In this section, we adapt the algorithm of [mcgregor2018simple] to obtain a private approximate maximum cardinality algorithm in the continual release model using small, sublinear space in the number of vertices. As in their algorithm, we assume that we are provided with a public upper bound $\tilde{\alpha}$ on the maximum arboricity $\alpha$ of the graph at any point in the stream. The privacy of our algorithm is always guaranteed. However, we do not assume that $\tilde{\alpha}$ is guaranteed to upper bound $\alpha$; when $\tilde{\alpha}\geq\alpha$, then our approximation guarantees hold with high probability. Otherwise, our approximation guarantees do not necessarily hold. Note that the same type of guarantee holds for the original non-private streaming algorithm [mcgregor2018simple] where their utility guarantee is only given when their public estimate of $\alpha$ upper bounds the maximum arboricity of the input. We prove the following result in this section.
Theorem 4.1.
Fix $\eta\in(0,1]$. Given a public estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream [^11] We can eliminate this assumption with an additional pass of the stream or notify the observer when the utility guarantees no longer hold in the one-pass setting. See Appendix E., Algorithm 4.1 is an $\varepsilon$-edge DP algorithm for estimating the size of the maximum matching in the continual release model for insertion-only streams. If $\tilde{\alpha}\geq\alpha$, then with probability at least $1-1/\operatorname*{poly}(n)$, our algorithm returns a $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{\varepsilon\eta}\right)\right)$-approximation of the size of the maximum matching at every timestamp. Moreover, our algorithm uses $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon\eta^{2}}\right)$ space with probability $1-1/\operatorname*{poly}(n)$.
4.1.1 Algorithm Description
Algorithm Intuition
We revise the algorithm of [mcgregor2018simple] to the insertion-only continual release setting. On a high-level, [mcgregor2018simple] showed that the cardinality of a carefully chosen subset of edges $F\subseteq E$ is a good estimator for the size of the maximum matching for graphs of bounded arboricity $\alpha$. $F$ is obtained from $E$ by deleting edges $e$ adjacent to a vertex $v$ if more than $\alpha$ other edges adjacent to $v$ arrived after $e$. Then their algorithm maintains a small sample $S\subseteq F$ throughout the algorithm by down-sampling edges when the current sample exceeds some threshold. Our algorithm follows a similar approach with two main adjustments to satisfy privacy. First, we release powers of $(1+\eta)$ based on an SVT comparison against the current value of the estimator. Second, the decision to down-sample is also based on an SVT comparison against the threshold.
Detailed Algorithm Description
The pseudocode for our algorithm is given in Algorithm 4.1. Our algorithm works as follows. We set the probability of sampling edges from the stream initially to $1$. For each update $e_{t}$, we sample the update into $S$ with probability $p_{t}$ (Algorithm 4.1). If the sampled update is not $\bot$ (Algorithm 4.1), then we initialize counters for its endpoints if one or both of these counters do not already exist (Algorithm 4.1). A sampled $e_{t}$ that is $\bot$ does not increase the size of $S$ (it is a no-op). Then, we iterate through every other edge in $S$ (Algorithm 4.1) and if $e^{\prime}$ shares an endpoint with $e_{t}$ (Algorithm 4.1), then we increment the counter $c_{e^{\prime}}^{w}$ associated with $e^{\prime}$ (Algorithm 4.1). If the counter exceeds our cutoff $\tilde{\alpha}$ (Algorithm 4.1), then we remove $e^{\prime}$ and its corresponding counters from $S$ (Algorithm 4.1). After performing our edge sampling, we now check using SVT whether $|S|$ (the number of edges in $S$) exceeds $\frac{a_{3}\log^{2}n}{\varepsilon\eta^{2}}$ for some large enough constant $a_{3}>0$ (Algorithm 4.1). If the SVT AboveThreshold query is satisfied, then we halve the probability of sampling (Algorithm 4.1) and resample each edge in $S$ with $1/2$ probability (Algorithm 4.1). Finally, we output a new estimate if $|S|/p$ is greater than our previous estimate by a significantly large amount. Specifically, we check via SVT queries (Algorithm 4.1) whether $|S|$ exceeds $p\cdot(1+\eta)^{j_{t}}$. If so, we increment the counter $j_{t}$ and output our new estimate as $(1+\eta)^{j_{t}}$ (Algorithm 4.1); otherwise, we output our old estimate.
1 Class EdgePrivateSublinearMatching($\tilde{\alpha},\varepsilon,\eta,n,T$)
2       $S\leftarrow\varnothing$
3       $p_{1}\leftarrow 1$
4       $Q_{1}\leftarrow a_{1}\log(n),Q_{2}\leftarrow a_{2}\eta^{-1}\log(n)$
5       $j_{1}\leftarrow 0$
6       estimate $\leftarrow(1+\eta)^{j_{1}}$
7       Initialize class $\textsc{SubsampleSVT}\leftarrow\textsc{SVT}(\nicefrac{{\varepsilon}}{{2}},2,Q_{1})$   (Algorithm 3.1)
8       Initialize class $\textsc{EstimateSVT}\leftarrow\textsc{SVT}(\nicefrac{{\varepsilon}}{{2}},2,Q_{2})$
9       Function ProcessUpdate($e_{t}$)
10             if $e_{t}\not=\bot$ then
11                   With probability $p_{t}$ add $e_{t}$ to $S$ and initialize counters $c_{e}^{u}\leftarrow 0$ and $c_{e}^{v}\leftarrow 0$.
12                   for each edge $e^{\prime}\not=e_{t}\in S$ do
13                         $\triangleright$ even if $e_{t}$ is not sampled, it affects the counters of other edges
14                         if $e^{\prime}=\{w,w^{\prime}\}$ shares endpoint $w$ with $e_{t}$ then
15                               Increment $c_{e^{\prime}}^{w}$
16                               if $c_{e^{\prime}}^{w}>\tilde{\alpha}$ then
17                                     Remove $e^{\prime}$ from $S$ and delete $c_{e^{\prime}}^{w},c_{e^{\prime}}^{w^{\prime}}$
18
19
20
21
22            while EstimateSVT.ProcessQuery$(|S|,p_{t}\cdot(1+\eta)^{j_{t}})$ is “above” do
23                   $j_{t}\leftarrow j_{t}+1$
24                   estimate $\leftarrow(1+\eta)^{j_{t}}$
25
26            Output estimate.
27             $p_{t+1}\leftarrow p_{t}$, $j_{t+1}\leftarrow j_{t}$
28             if SubsampleSVT.ProcessQuery$(|S|,\frac{a_{3}\log^{2}(n)}{\varepsilon\eta^{2}})$ is “above” then
29                   $p_{t+1}\leftarrow p_{t}/2$
30                   for $e=\{u,v\}\in S$ do
31                         With probability $\nicefrac{{1}}{{2}}$ remove $e$ from $S$ and delete $c_{e}^{u},c_{e}^{v}$
32
33
34
35
Algorithm 4.1 Sampling $E_{t}^{*}$
4.1.2 Privacy Guarantee
We prove the privacy guarantees of our algorithm in this section.
Lemma 4.2.
Algorithm 4.1 is $\varepsilon$-edge differentially private.
Proof.
By the definition of edge-neighboring streams (Definition 3.3), let ${\mathcal{S}}=(e_{1},\ldots,e_{{T}}),{\mathcal{S}}^{\prime}=(e^{\prime}_{1},\ldots,e^{\prime}_{{T}})$ be neighboring streams of edges, i.e., there exists $t^{*}\in[{T}]$ such that $e_{t^{*}}\neq e^{\prime}_{t^{*}}$ and $e^{\prime}_{t^{*}}=\bot$. Note that $e_{t}$ for any $t\in[T]$ can either be an edge $\{u,v\}$ or $\bot$. We define the event ${\mathcal{P}}$ where the probability for sampling $e_{t}$ from the stream ${\mathcal{S}}$ is set to $p(e_{t})=p_{t}$ for every $t\in[{T}]$. Similarly define ${\mathcal{P}}^{\prime}$ where the probability for sampling $e^{\prime}_{t}$ from the stream ${\mathcal{S}}^{\prime}$ is set to $p(e^{\prime}_{t})=p_{t}$ for every $t\in[{T}]$.
Define $I_{-t^{*}}^{(t)}=(I^{(t)}_{1},\ldots,I^{(t)}_{t})$ for stream ${\mathcal{S}}$ where for $t^{*}\neq j\leq t$,
$I_{j}^{(t)}$ =$=\begin{cases}1&e_{j}\text{ is in the sample }S_{t}\\ 0&\text{otherwise}.\end{cases}$ (1)
Define ${I^{\prime}}_{-t^{*}}^{(t)}=({I^{\prime}}^{(t)}_{1},\ldots,{I^{\prime}}^{(t)}_{t})$ for stream ${\mathcal{S}}^{\prime}$ analogously. Let $\mathbf{b}^{(t)}$ be a $\{0,1\}$-vector with the same dimension as $I_{-t^{*}}^{(t)}$ and ${I^{\prime}}_{-t^{*}}^{(t)}$. Also let ${\mathcal{I}}=(I_{-t^{*}}^{(1)},\ldots,I_{-t^{*}}^{({T})})$ and ${\mathcal{I}}^{\prime}=(I_{-t^{*}}^{\prime(1)},\ldots,I_{-t^{*}}^{\prime({T})})$.
{adjustwidth}
3em0em
Claim 4.3.
For any $t\in[T]$, let $S_{t}$ be the set of sampled edges at timestep $t$. Conditioned on the events ${\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})}),{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})$, ${\mathcal{P}},{\mathcal{P}}^{\prime}$, the sensitivity of $|S_{t}|$ is at most 2.
Proof.
Let $S_{t}$ and $S^{\prime}_{t}$ denote the sampled edges of the neighboring streams at time $t$. By definition ${\mathcal{P}}$ and ${\mathcal{P}}^{\prime}$ define the sampling probabilities at every timestep $t\in[T]$ to be $p_{t}$. We consider the case when the differing edge at timestep $t^{*}$ is such that $e_{t^{*}}=\{u,v\}$ and $e^{\prime}_{t^{*}}=\perp$. If $t<t^{*}$, then the input stream so far is identical. Hence the sampled edges are the same, conditioned on the events in the lemma, and there is nothing to prove. Consider now the case that $t\geq t^{*}$. Processing $e_{t^{*}}$ may initially increase the sample size by $1$ but may also cause edges to be removed. We show that $e_{t^{*}}=\{u,v\}$ can at most cause $||S_{t}|-|S^{\prime}_{t}||\leq 2$ for all $t\geq t^{*}$. For every edge that is incident to $u$ or $v$ and is still in $S$, there is at most one edge $e^{\prime}$ incident to each $w\in\{u,v\}$ that has counter value $c_{e^{\prime}}^{w}=\tilde{\alpha}$. This is due to the fact that no two edges are inserted in the same timestep and the counters are incremented when a later edge adjacent to the vertex arrives. In other words, for each vertex $u$, the set of edges adjacent to $u$ always has distinct counts. Suppose that $e^{\prime}$ (adjacent to $u$ or $v$) is the edge removed from $G^{\prime}$ but not removed from $G$ due to the counter $c_{e^{\prime}}^{w}$ exceeding $\tilde{\alpha}$ ($w\in\{u,v\}$). Then, any future update that removes another edge from $G^{\prime}$ incident to $w$ must first cause $e^{\prime}$ to be removed. Hence, at any point in time $S^{\prime}_{t}$ contains at most two fewer edges incident to $w\in\{u,v\}$ compared to $S$, at most one per endpoint. $e_{t^{*}}$ does not affect the counters of any edges inserted after it. Hence, $||S_{t}|-|S^{\prime}_{t}||\leq 2$ for all $t\geq t^{*}$. ∎
{adjustwidth}
3em0em
Claim 4.4.
$\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon/2}$
Proof.
By 4.3, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2. The subsample SVT output is $\varepsilon/2$-DP by Theorem 3.13. Now, observe that the probability $p$ of sampling only changes whenever the subsample SVT (Algorithm 4.1) accepts. Since privacy is preserved after post-processing, this implies the statement in the claim. ∎
{adjustwidth}
3em0em
Claim 4.5.
$\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}=\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})}),{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})}),{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon/2}$ (2)
Proof.
By definition, ${\mathcal{P}}$ and ${\mathcal{P}}^{\prime}$ defines the sampling probability for all elements in ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ to be $p_{t}$. Hence, it holds that $\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}=1$. By 4.4, $\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon/2}$ and our claim follows. ∎
Let ${\mathcal{Q}}$ be the event that $({\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(T)}),{\mathcal{P}})$ and define ${\mathcal{Q}}^{\prime}$ analogously.
{adjustwidth}
3em0em
Claim 4.6.
Let $\mathbf{a}\in\{\text{``above''},\text{``below''},\text{``abort''}\}^{T}$ be a sequence of the estimate SVT answers in Algorithm 4.1. We have
$\frac{\mathbf{P}[SVT(|S|)=\mathbf{a}|{\mathcal{Q}}]}{\mathbf{P}[SVT(|S^{\prime}|)=\mathbf{a}|{\mathcal{Q}}^{\prime}]}$ $\leq e^{\varepsilon/2}$ (3)
Proof.
By 4.3, since the sensitivity of the sampled set of edges $S$ is upper bounded by 2, the estimate SVT output is $\varepsilon/2$-DP by Theorem 3.13. ∎
By the chain rule, putting the above claims together gives us
$\frac{\mathbf{P}[SVT(|S|)=\mathbf{a}|{\mathcal{Q}}]}{\mathbf{P}[SVT(|S^{\prime}|)=\mathbf{a}|{\mathcal{Q}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}$ $\leq e^{2\varepsilon/2}=e^{\varepsilon}.$ (4)
4.1.3 Approximation Guarantees
We now show the approximation guarantees of our algorithm. The below proofs modify the proofs of [mcgregor2018simple] in the non-private setting to the private setting. In particular, we modify their proof to hold for any $\tilde{\alpha}\geq\alpha$.
Let $M(G_{t})$ be the maximum size of a matching in input graph $G_{t}$ (consisting of all updates up to and including update $e_{t}$). We wish to compute an approximation of $M(G_{t})$. Let $B_{u}^{t}$ be the last $\tilde{\alpha}+1$ edges incident to $u\in V$ that appears in $G_{t}$. Let $E_{\tilde{\alpha}}^{t}$ be the set of edges $\{u,v\}\in G_{t}$ where $|B_{u}^{t}|\leq\tilde{\alpha}$ and $|B_{v}^{t}|\leq\tilde{\alpha}$. That is, $E_{\tilde{\alpha}}^{t}$ consists of the set of edges $\{u,v\}$ where the number of edges incident to $u$ and $v$ that appear in the stream after $\{u,v\}$ is at most $\tilde{\alpha}$. We say an edge $\{u,v\}$ is good if $\{u,v\}\in B_{u}^{t}\cap B_{v}^{t}$, and an edge is wasted if $\{u,v\}\in B_{u}^{t}\oplus B_{v}^{t}=(B_{u}^{t}\cup B_{v}^{t})\setminus(B_{u}^{t}\cap B_{v}^{t})$. Then, $E_{\tilde{\alpha}}^{t}$ is precisely the set of good edges in $G^{t}$. In other words, $E_{\tilde{\alpha}}^{t}=\bigcup_{u\neq v\in V}\left(B_{u}^{t}\cap B_{v}^{t}\right)$.
Our algorithm estimates the cardinality of $E^{t}_{\tilde{\alpha}}$ as an approximation of $M(G_{t})$. So, we first relate $|E^{t}_{\tilde{\alpha}}|$ to $M(G_{t})$.
Lemma 4.7.
$M(G^{t})\leq|E_{\tilde{\alpha}}^{t}|\leq(\tilde{\alpha}+2)\cdot M(G_{t})$.
Proof (Lemma 4.7).
We first prove the right-hand side of this expression. To do this, we define a fractional matching using $E_{\tilde{\alpha}}^{t}$. Let $Y_{e}=\frac{1}{\tilde{\alpha}+1}$ if $e\in E_{\tilde{\alpha}}^{t}$ and $Y_{e}=0$ otherwise. Then, $\{Y_{e}\}_{e\in E}$ is a fractional matching with maximum weight $\frac{1}{\tilde{\alpha}+1}$. [^12] A fractional matching is defined to be a set of weights $f(e)\geq 0$ on every edge $e$ where $\forall v\in V:\sum_{e\ni v}f(e)\leq 1$. We now show a corollary of Edmonds’ matching polytope theorem [edmonds1965maximum]. Edmonds’ matching polytope theorem implies that if the weight of a fractional matching on any induced subgraph $S\subseteq G$ is at most $\frac{|S|-1}{2}$, then the weight on the entire graph is at most $M(G)$. Now, we show the following proposition:
{adjustwidth}
3em0em
Proposition 4.8.
Let $\{Y_{e}\}_{e\in E}$ be a fractional matching where the maximum weight on any edge is $\psi$. Then, $\sum_{e\in E}Y_{e}\leq(1+\psi)\cdot M(G)$.
Proof (Proposition 4.8).
Let $S$ be an arbitrary subset of vertices, and let $E(S)$ be the edges in the induced subgraph of $S$. We know that $|E(S)|\leq\frac{|S|(|S|-1)}{2}$ and by the definition of fractional matching, $\sum_{e\in E(S)}Y_{e}=\frac{\sum_{u\in S}\sum_{v\in N(u)}Y_{\{u,v\}}}{2}\leq\frac{\sum_{u\in S}1}{2}=\frac{|S|}{2}$ where $\sum_{v\in N(u)}Y_{\{u,v\}}\leq 1$ by the constraints of the fractional matching. Thus, we know that (given the maximum weight on any edge is $\psi$)
$\sum_{e\in E(S)}Y_{e}$ $\leq\min\left(\frac{|S|}{2},\frac{\psi|S|(|S|-1)}{2}\right)$ (5)
$\leq\frac{|S|-1}{2}\cdot\min\left(\frac{|S|}{|S|-1},\psi|S|\right)$ (6)
$\leq\frac{|S|-1}{2}\cdot(1+\psi).$ (7)
We can show the last inequality by considering two cases:
  • If $\frac{|S|}{|S|-1}\leq\psi|S|$, then
    $\frac{1}{|S|-1}$ $\leq\psi$ (8)
    $1+\frac{1}{|S|-1}$ $\leq 1+\psi$ (9)
    $\frac{|S|}{|S|-1}$ $\leq 1+\psi.$ (10)
  • If $\psi|S|\leq\frac{|S|}{|S|-1}$, then
    $\psi$ $\leq\frac{1}{|S|-1}$ (11)
    $\psi|S|-\psi$ $\leq 1$ (12)
    $\psi|S|$ $\leq 1+\psi.$ (13)
Finally, let $Z_{e}=\frac{Y_{e}}{1+\psi}$. We can use Edmonds’ polytope theorem to show that $\sum_{e\in E}Z_{e}\leq\frac{|S|-1}{2}\leq M(G)$ and so $\sum_{e\in E}Y_{e}\leq(1+\psi)\sum_{e\in E}Z_{e}\leq(1+\psi)\cdot M(G)$. We have proven our corollary. ∎
Using the proposition above with maximum weight $\frac{1}{\tilde{\alpha}+1}$ implies that $\sum_{e\in E}Y_{e}=\frac{|E_{\tilde{\alpha}}^{t}|}{\tilde{\alpha}+1}\leq(1+\frac{1}{\tilde{\alpha}+1})M(G)=\frac{\tilde{\alpha}+2}{\tilde{\alpha}+1}\cdot M(G)$. Hence, we have that $|E_{\tilde{\alpha}}^{t}|\leq(\tilde{\alpha}+2)\cdot M(G)$.
Now, we prove the left inequality. To prove the left inequality, let $H$ be the set of vertices in $G_{t}$ with degree at least $\tilde{\alpha}+1$. These are the heavy vertices. We also define the following variables:
  • $w:=$ the number of good edges with no endpoints in $H$,
  • $x:=$ the number of good edges with exactly one endpoint in $H$,
  • $y:=$ the number of good edges with two endpoints in $H$,
  • $z:=$ the number of wasted edges with two endpoints in $H$.
First, $|E^{t}_{\tilde{\alpha}}|=w+x+y$. Below, we omit the superscript $t$ for clarity but everything holds with respect to $t$. Now, we show the following additional equalities. We will first calculate the number of edges in the $B_{u}$ of every $u\in H$ in terms of the variables we defined above. Since every vertex $u\in H$ is heavy, $|B_{u}|=\tilde{\alpha}+1$. Hence, $\sum_{u\in H}|B_{u}|=(\tilde{\alpha}+1)|H|$. Every edge $\{u,v\}$ in each of these $B_{u}$ must either be a good edge or a wasted edge by definition since $\{u,v\}\in B_{u}\cup B_{v}$. For each edge counted in $x$, it has one endpoint in $H$ so it is counted in exactly one $B_{u}$ in $H$. Furthermore, for each edge counted in $z$, it is also counted in the $B_{u}$ of exactly one of its two endpoints since it is wasted (i.e. the other endpoint doesn’t have at most $\tilde{\alpha}$ edges that come after it). This leaves every good edge counted in $y$ which is counted in exactly two $B_{u}$’s. Hence, $\sum_{u\in H}|B_{u}|=x+2y+z=(\tilde{\alpha}+1)|H|$.
Now, we can also compute $z+y\leq\alpha|H|$ since $z+y$ is a subset of the total number of edges in the induced subgraph consisting of $H$. Since we know that the graph has arboricity $\alpha$, we also know that (by the properties of arboricity) that the induced subgraph consisting of $H$ has at most $\alpha|H|$ edges. Hence, we can sum our inequalities: $x+2y+z=(\tilde{\alpha}+1)|H|$ and $-z-y\geq-\alpha|H|$ to obtain $x+y\geq(\tilde{\alpha}-\alpha+1)|H|$. Finally, let $E_{L}$ be the set of edges with no endpoints in $H$. Every edge $\{u,v\}\in E_{L}$ is good since $B_{u}=N(u)$ and $B_{v}=N(v)$ by assumption. Note that $w=|E_{L}|$. Therefore, $|E^{t}_{\tilde{\alpha}}|=w+x+y\geq|H|+|E_{L}|$ since $\tilde{\alpha}\geq\alpha$. We can show that $|H|+|E_{L}|\geq M(G)$ since we can partition the set of edges in the maximum matching to edges in $E_{L}$ and edges incident to $H$. All of the edges in $E_{L}$ can be in the matching and at most one edge incident to each vertex in $H$ is in the matching. We have successfully proven our lower bound: $|E^{t}_{\tilde{\alpha}}|\geq M(G)$. ∎
Now we have the following algorithm for estimating $|E_{\tilde{\alpha}}^{t}|$ for every $t\in[T]$. For every sampled edge, $e=\{u,v\}$, the algorithm also stores counters $c_{e}^{u}$ and $c_{e}^{v}$ for the degrees of $u$ and $v$ in the rest of the stream. This requires an additional factor of $O(\log(\tilde{\alpha}))$ space. Thus, the algorithm maintains the invariant that each edge stored in the sample is a good edge with respect to the current $G_{t}$.
Let $E^{*}_{t}=\max_{t^{\prime}\leq t}(|E_{\tilde{\alpha}}^{t^{\prime}}|)$, we now show that we get a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ with high probability. First, we note that $E^{*}_{t}$ satisfies $M(G^{t})\leq E^{*}_{t}\leq(2+\tilde{\alpha})M(G_{t})$ since $E^{*}_{t}\geq|E_{\tilde{\alpha}}^{t}|$, $M(G_{t^{\prime}})\leq M(G_{t})$, and Lemma 4.7.
We now show our main lemma for our algorithm.
Lemma 4.9.
Algorithm 4.1 returns a $\left(1+\eta,\frac{\log^{2}(n)}{\varepsilon\eta}\right)$-approximation of $E_{t}^{*}$ for every $t\in[T]$ with probability at least $1-1/\operatorname*{poly}(n)$.
Proof.
Fix some $t\in[T]$. We need to show at time $t$, we do not exceed the SVT “above” budgets as well as that sub-sampling yields sufficiently concentrated estimates with probability $1-1/\operatorname*{poly}(n)$. Then, taking a union bound over $T=\operatorname*{poly}(n)$ timestamps yields the desired result.
First, let $\tau=\frac{d\log^{2}(n)}{\varepsilon\eta^{2}}$ for some large enough constant $d>0$ and define level $i_{t}$ (starting with $i_{t}=2$) to be the level that contains $E_{t}^{*}$ if $2^{i_{t}-1}\cdot\tau<E^{*}_{t}\leq 2^{i_{t}}\cdot\tau$. Level $i_{t}=1$ is defined as $0\leq E^{*}_{t}\leq 2\cdot\tau$. Let $p_{t}$ denote the marginal probability that an edge is sampled from $E_{\alpha}^{t}$. Note that this probability is prior to sub-sampling on Algorithm 4.1. In a perfect situation, at time $t$, edge $e$ is sampled in level $i$ with probability $p_{t}^{\prime}:=2^{-i_{t}}$ when $i_{t}\geq 2$ and with probability $p_{t}^{\prime}=1$ when $i_{t}=1$, but this is not the case as our algorithm determines $p_{t}$ adaptively depending on the size of the sampled edges at time $t$.
SVT Budget.
We show that with probability at least $1-1/\operatorname*{poly}(n)$, neither of the SVT “above” budgets $Q_{1}$, $Q_{2}$ in Algorithm 4.1 are exceeded. First, for the sub-sampling SVT, we take $T=\operatorname*{poly}(n)$ noisy comparisons with ${\textsf{Lap}}(O(\varepsilon/\log(n)))$ noise. Hence with probability $1-1/\operatorname*{poly}(n)$, all Laplace realizations are of order $O(\varepsilon^{-1}\log^{2}(n))$.
$Q_{1}=a_{1}\log(n)$ bounds the number of times the subsampling SVT in Algorithm 4.1 is satisfied. Each time this SVT is satisfied, we must have
$\lvert S_{t}\rvert\geq a_{3}\varepsilon^{-1}\eta^{-2}\log^{2}(n)-O(\varepsilon^{-1}\log^{2}(n))\geq\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)$ (14)
for sufficiently large constant $a_{3}>0$, leading to the probability of sampling edges being halved. The expected number of edges that are sampled at any point is $p_{t}\cdot\lvert E_{\tilde{\alpha}}^{t}\rvert\leq p_{t}\cdot n^{2}$. Suppose $a_{1}$ is sufficiently large and we halved the sampling probability $\nicefrac{{Q_{1}}}{{2}}\geq\lceil\log(n^{2})\rceil$ times so that $p_{t}\leq\frac{1}{n^{2}}$. We can upper bound the probability that the SVT answers “above” by considering the experiment where we obtain a sample $\tilde{S}_{t}$ by sampling each edge with probability $\tilde{p}=\frac{1}{\lvert E_{\tilde{\alpha}}^{t}\rvert}\geq p_{t}$. By a multiplicative Chernoff bound (Theorem 3.1), for $\psi=\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)-1$,
$\mathbf{P}\left[\lvert S_{t}\rvert\geq\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)\right]$ $\leq\mathbf{P}\left[\lvert\tilde{S}_{t}\rvert\geq\frac{a_{3}}{2}\varepsilon^{-1}\log^{2}(n)\right]$ (15)
$=\mathbf{P}[\lvert\tilde{S}_{t}\rvert\geq(1+\psi)\cdot 1]$ (16)
$=\mathbf{P}[\lvert\tilde{S}_{t}\rvert\geq(1+\psi)\tilde{p}\cdot\lvert E_{\tilde{\alpha}}^{t}\rvert]$ (17)
$\leq\exp\left(-\frac{\psi^{2}\cdot 1}{2+\psi}\right)$ (18)
$\leq\exp\left(-\Omega(\log^{2}(n))\right)$ $n$ sufficiently large (19)
$\leq\frac{1}{\operatorname*{poly}(n)}.$ (20)
Hence, for large enough constants $a_{1},a_{3}$, we do not exceed the subsampling SVT “above” budget.
A similar argument shows that we do not exceed the estimate SVT “above” budget of $Q_{2}$ with probability at least $1-1/\operatorname*{poly}(n)$. In the worst case, the number of times $p$ is halved is $Q_{1}$ times. Thus, the maximum value of $|S|/p$ is $2^{Q_{1}}\cdot n^{2}$ since the maximum number of edges in the graph is upper bounded by $n^{2}$. Since we are increasing our threshold by a factor of $(1+\eta)$ each time, our threshold exceeds $2^{Q_{1}}n^{2}$ after $\log_{1+\eta}(2^{Q_{1}}n^{2})=O(\eta^{-1}Q_{1}+\eta^{-1}\log(n))=O(\eta^{-1}Q_{1})$ times. If we set $a_{2}$ to be a large enough constant, we do not exceed the “above” budget for the estimate SVT.
Concentration.
In a perfect situation, edge $e$ is sampled at time $t$ with probability $p^{\prime}_{t}=2^{-i_{t}}$ for $i_{t}\geq 2$ and with probability $p^{\prime}_{t}=1$ for $i_{t}=1$. In that case, we can use the multiplicative Chernoff bound to bound the probability that our estimate concentrates as our sampling probability would be sufficiently high enough to do so. However, it is not the case that $p_{t}$ is guaranteed to be $2^{-i_{t}}$ since it is determined adaptively by Algorithm 4.1. Hence, we need a slightly more sophisticated analysis than simply using the multiplicative Chernoff bound with $p^{\prime}_{t}$.
We consider two cases. If $p_{t}\geq p^{\prime}_{t}$, then we can lower bound the probability of success by what we would obtain using $p_{t}^{\prime}$ and we can use the multiplicative Chernoff bound in this case. Now, suppose that $p_{t}<p^{\prime}_{t}$; then, we claim that this case never occurs in Algorithm 4.1 with probability at least $1-1/\operatorname*{poly}(n)$. We now formally present these arguments.
If $p_{t}\geq p_{t}^{\prime}$, then we use a multiplicative Chernoff bound (Theorem 3.1) to bound our probability of success. Let $s_{t}$ be the number of edges we sampled for $E_{\tilde{\alpha}}^{t}$ prior to any sub-sampling. That is, we obtain $s_{t}$ by sampling with probability $p_{t}$ before Algorithm 4.1. For $\psi:=\eta\cdot E_{t}^{*}/\lvert E_{\tilde{\alpha}}^{t}\rvert$,
$\mathbf{P}\left[s_{t}-p_{t}\cdot|E_{\tilde{\alpha}}^{t}|\geq\eta\cdot p_{t}\cdot E_{t}^{*}\right]$ $=\mathbf{P}\left[s_{t}-p_{t}\cdot|E_{\tilde{\alpha}}^{t}|\geq\psi\cdot p_{t}\cdot|E_{\tilde{\alpha}}^{t}|\right]$ (21)
$\leq\exp\left(-\frac{\psi^{2}}{2+\psi}\cdot p_{t}\cdot|E_{\tilde{\alpha}}^{t}|\right)$ (22)
$\leq\exp\left(-\frac{\psi^{2}}{2+\psi}\cdot p_{t}^{\prime}\cdot|E_{\tilde{\alpha}}^{t}|\right).$ $p_{t}\geq p_{t}^{\prime}$ (23)
When $\psi\geq 1$, we have $\frac{\psi}{2+\psi}\geq\frac{1}{3}$ and this is at most
$\exp\left(-\frac{\psi}{3}\cdot p_{t}^{\prime}\cdot|E_{\tilde{\alpha}}^{t}|\right)$ $=\exp\left(-\frac{\eta E_{t}^{*}}{3}p_{t}^{\prime}\right)$ (24)
$\leq\exp\left(-\Omega(\varepsilon^{-1}\eta^{-1}\log^{2}(n))\right)$ definition of $p_{t}^{\prime}$, sufficiently large $a_{3}$ (25)
$=\frac{1}{\operatorname*{poly}(n)}.$ (26)
When $\psi\in(0,1)$, this is at most
$\exp\left(-\frac{\psi^{2}}{3}\cdot p_{t}^{\prime}\cdot E_{t}^{*}\right)$ $\leq\exp\left(-\eta^{2}E_{t}^{*}p_{t}^{\prime}\right)$ (27)
$\leq\exp\left(-\Omega(\varepsilon^{-1}\log^{2}(n))\right).$ definition of $p_{t}^{\prime}$, sufficiently large $a_{3}$ (28)
$=\frac{1}{\operatorname*{poly}(n)}$ (29)
Since we need only consider the case of $\psi\in(0,1)$ for the lower bound, an identical calculation yields the same concentration bound above but for the lower bound. All in all, $\frac{s_{t}}{p_{t}}\in[|E_{\tilde{\alpha}}^{t}|\pm\eta E_{t}^{*}]$ with high probability.
Now, we consider the case when $p_{t}<p_{t}^{\prime}$. We show that with probability at least $1-1/\operatorname*{poly}(n)$, this case does not occur. We prove this via induction on $t$. For $t=1$, $p_{t}=p_{t}^{\prime}$ trivially since both equal $1$. Now, we assume our claim holds for $t$ and show it holds for $t+1$. First, note that $E^{*}_{t}$ cannot decrease so $p_{t}^{\prime}$ cannot increase as $t$ increases. Then, by Equation 14, $p_{t+1}<p_{t+1}^{\prime}$ only if we sample more than $\frac{a_{3}^{\prime}\log^{2}(n)}{2\varepsilon\eta^{2}}$ elements of $E_{\tilde{\alpha}}^{t}$ at probability $p_{t}=p_{t}^{\prime}$ before Algorithm 4.1. However, we have just shown that with high probability,
$s_{t}\leq(1+\eta)p_{t}E_{t}^{*}\leq 2p_{t}^{\prime}E_{t}^{*}\leq 2\tau<\frac{a_{3}\log^{2}(n)}{2\varepsilon\eta^{2}}$ (30)
by the definition of $p_{t}^{\prime}$, $\tau$, and assuming a sufficiently large $a_{3}$. Hence this event does not occur with high probability.
Finally, the estimate SVT ensures that with high probability,
$s_{t}$ $\leq p_{t}(1+\eta)^{j_{t}}+O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))$ (31)
$s_{t}$ $>p_{t}(1+\eta)^{j_{t}-1}-O(\varepsilon^{-1}\eta^{-1}\log^{2}(n)).$ (32)
But we wish to approximate $s_{t}/p_{t}$, and dividing the inequalities by $p_{t}$ yield
$\frac{s_{t}}{p_{t}}$ $\leq(1+\eta)^{j_{t}}+\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}$ (33)
$\frac{s_{t}}{p_{t}}$ $>(1+\eta)^{j_{t}-1}-\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}.$ (34)
If $i_{t}=1$ then $p_{t}=1$ and we are done. Thus consider the case where $i_{t}\geq 2$. We have shown above that $p_{t}\geq p_{t}^{\prime}=2^{-i_{t}}$ and $2^{i_{t}-1}\tau\leq E_{t}^{*}$ for $i_{t}\geq 2$. Thus in this case,
$\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}$ $\leq\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))E_{t}^{*}}{\tau}$ (35)
$=O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))E_{t}^{*}\cdot\frac{\varepsilon\eta^{2}}{d\log^{2}(n)}$ (36)
$\leq\eta E_{t}^{*}$ (37)
assuming a sufficiently large constant $d$.
Thus, combining the case work above with our concentration bounds on $s_{t}/p_{t}$, we see that $(1+\eta)^{j_{t}}$ is a $\left(1+\eta,O\left(\varepsilon^{-1}\eta^{-1}\log^{2}n\right)\right)$-approximation of $E_{t}^{*}$.
Now, combining Lemma 4.7 with Lemma 4.9 gives us our final accuracy guarantee in Theorem 4.1. In terms of space complexity, we need $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon\eta^{2}}\right)$ space to store the sampled edges and the two counters per edge.
4.2 Bounded Arboricity Sparsifiers
For our bounded arboricity graphs, we take inspiration from the bounded arboricity sparsifier of Solomon [solomon2021local]. A closely related line of work is that of edge degree constrained subgraphs (EDCS) [assadi2019coresets, bernstein2015fully, bernstein2016faster]. We modify the sparsifiers from [solomon2021local] to show Proposition 4.11, which states that node-neighboring graphs have small edge edit distance post-sparsification.
Our proofs use the sensitivity of the arboricity of the graph which we prove below.
Proposition 4.10 (Node Sensitivity of Arboricity).
The node-sensitivity of the arboricity of a graph $G$ is $1$.
Proof.
By the Nash-Williams theorem [nash1964decomposition], the arboricity $\alpha$ of $G$ is equal to $\max_{S\subseteq G}\left\{\lceil\frac{m_{S}}{n_{S}-1}\rceil\right\}$. Suppose $S_{\max}\subseteq G$ is a subgraph that satisfies the maximization condition. Then, removing the edges adjacent to any vertex $v\in S_{max}$ decreases the number of edges $m_{S_{\max}}$ by at most $X$ where $n_{S_{\max}}\geq X+1$. Thus,
$\left\lceil\frac{m_{S_{\max}}-X}{n_{S_{\max}}-1}\right\rceil\geq\left\lceil\frac{m_{S_{\max}}}{n_{S_{\max}}-1}\right\rceil-\left\lceil\frac{X}{n_{S_{\max}}-1}\right\rceil\geq\left\lceil\frac{m_{S_{\max}}}{n_{S_{\max}}-1}\right\rceil-\left\lceil\frac{X}{X}\right\rceil=\left\lceil\frac{m_{S_{\max}}}{n_{S_{\max}}-1}\right\rceil-1.$
Hence, the arboricity of the graph with all edges adjacent to $v$ removed is at least $\alpha-1$ and the sensitivity is at most $1$. ∎
We first adopt some arguments from [solomon2021local] for orderings of vertices. Let $G=(V,E)$ be a graph. For a real number $\Lambda\geq 0$, define the following vertex sets and subgraphs.
$V_{>}^{\Lambda}$ :=$:=\{v\in V:\deg(v)>\Lambda\}$ (38)
$V_{\leq}^{\Lambda}$ :=$:=\{v\in V:\deg(v)\leq\Lambda\}$ (39)
$G_{>}^{\Lambda}$ $:=(V,E[V_{>}^{\Lambda}])$ (40)
$G_{\leq}^{\Lambda}$ $:=(V,E[V_{\leq}^{\Lambda}]).$ (41)
That is, $G_{\leq}^{\Lambda}$ is obtained from $G$ by deleting edges adjacent to high degree vertices and vice versa for $G_{>}^{\Lambda}$. When the context is clear, we may omit the superscript $\Lambda$.
[…]
Proposition 4.11.
Let $\pi\in P_{\binom{n}{2}}$ be a total ordering over unordered vertex pairs and $G\sim G^{\prime}$ be node neighboring graphs. Then the edge edit distance between $H:=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ and $H^{\prime}:=\textsc{ContractionSparsify}_{\pi}(G^{\prime},\Lambda)$ is at most $2\Lambda$.
Proof (Proposition 4.11).
Let $S_{G}$ and $S_{G^{\prime}}$ be the sparsified graphs of $G$ and $G^{\prime}$, respectively. Suppose without loss of generality that $G^{\prime}$ contains $E_{\operatorname{extra}}$ additional edges incident to vertex $v$ and $v$ has degree 00 in graph $G$ (by Definition 3.7). Then, for each edge $\{v,w\}\in E_{\operatorname{extra}}$, let $e^{w}_{\operatorname{last}}$ be the edge adjacent to $w$ in $G$ whose index in $\pi$ is the last among the edges incident to $w$ in $S_{G}$. If $i_{\pi}(e^{w}_{\operatorname{last}})>i_{\pi}(\{v,w\})$ (where $i_{\pi}(\{v,w\})$ is the index of edge $\{v,w\}$ in $\pi$), then $\{v,w\}$ replaces edge $e^{w}_{\operatorname{last}}$. Since both $G$ and $G^{\prime}$ are simple graphs, at most one edge incident to $w$ gets replaced by an edge in $E_{\operatorname{extra}}$ in $S_{G^{\prime}}$. This set of edge replacements leads to an edge edit distance of $2\Lambda$. ∎
Matching Sparsifier
The original sparsification algorithm in [solomon2021local] marks an arbitrary set of $\Lambda$ edges incident to every vertex and takes the subgraph consisting of all edges marked by both endpoints. In our setting, $\pi$ determines the arbitrary marking in our graphs. Hence, our sparsification procedure satisfies the below guarantee.
Theorem 4.12 (Theorem 3.3 in [solomon2021local]).
Let $\pi\in P_{\binom{n}{2}}$ be a total order over unordered vertex pairs, $G$ be a graph of arboricity at most $\alpha$, $\Lambda:=5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha$ for some $\eta\in(0,1]$, and $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$. Then if $\mu(\cdot)$ denotes the size of a maximum matching of the input graph,
$\mu(H)\leq\mu(G)\leq(1+\eta)\mu(H).$
In particular, any $(\beta,\zeta)$-approximate matching for $H$ is a $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$.
4.3 Node-DP Maximum Cardinality Matching
In this section, we design node-DP algorithms to estimate the size of a maximum matching (Algorithm 4.2) in the continual release model using the sparsification techniques derived in Section 4.2 and any edge DP algorithm to estimate the size of a matching in the insertion-only continual release model (e.g. Theorem 4.1). The below first theorem uses our truly sublinear $\varepsilon$-edge DP algorithm given in Section 4.1. It achieves better space bounds at a slightly worse multiplicative approximation guarantee.
Theorem 4.13.
Fix $\eta\in(0,1]$. Given a public estimate $\tilde{\alpha}$ on the maximum arboricity $\alpha$ over the stream [^13] See Appendix E for ways to address or remove this assumption., Algorithm 4.2 is an $\varepsilon$-node DP algorithm for estimating the size of a maximum matching in the insertion-only continual release model with the following guarantees:
  1. (i)
    If $\tilde{\alpha}\geq\alpha$, then with probability at least $1-1/\operatorname*{poly}(n)$, it outputs a $((1+\eta)(2+\tilde{\alpha}),O(\varepsilon^{-1}\eta^{-2}\tilde{\alpha}\log^{2}(n)))$-approximate estimation of the size of a maximum matching at every time stamp.
  2. (ii)
    The space complexity is $O(n\log(\eta^{-1}\tilde{\alpha})+\varepsilon^{-1}\eta^{-3}\tilde{\alpha}\log^{2}(n)\log(\tilde{\alpha}))$.
  3. (iii)
    The preprocessing and update times are polynomial.
Our second theorem uses the exact, static maximum matching algorithm of Micali-Vazirani (Theorem B.1) with a slightly worse space guarantee but better multiplicative error.
Theorem 4.14.
Fix $\eta\in(0,1]$. Given a public estimate $\tilde{\alpha}$ on the maximum arboricity $\alpha$ over the stream [^14] We can eliminate this assumption with an additional pass of the stream or notify the observer when the utility guarantees no longer hold in the one-pass setting. See Appendix E., there is is an $\varepsilon$-node DP algorithm for estimating the size of a maximum matching in the insertion-only continual release model with the following guarantees:
  1. (i)
    If $\tilde{\alpha}\geq\alpha$, then with probability at least $1-1/\operatorname*{poly}(n)$, it outputs a $((1+\eta),O(\varepsilon^{-1}\eta^{-2}\tilde{\alpha}\log^{2}(n)))$-approximate estimation of the size of a maximum matching at every time stamp.
  2. (ii)
    The space complexity is $O(\eta^{-1}n\tilde{\alpha})$.
  3. (iii)
    The preprocessing and update times are polynomial.
4.3.1 Algorithm Description & Guarantees
On a high level, we implement the matching sparsifier given in Theorem 4.12 in the insertion-only streaming setting and then feed the sparsified edge updates to Algorithm 4.1. This requires a public guess $\tilde{\alpha}$ for the maximum arboricity of all graphs within the stream. On top of the space usage of Algorithm 4.2, our algorithm uses an additional $O(n\log(\eta^{-1}\tilde{\alpha}))$ space to store a degree counter per node, which is sub-linear in the number of edges when $m=\omega(n)$. Our algorithms are always private but the utility is only guaranteed when $\tilde{\alpha}\geq\alpha$.
1 Function Alg($\tilde{\alpha},\varepsilon,\eta,n,T$)
2
3      $\Lambda\leftarrow 5\left(1+5\eta^{-1}\right)2\tilde{\alpha}$
4       Initialize counter $d_{v}\leftarrow 0$ for all $v\in V$
[…]
6       for incoming update $e_{t}$ at time $t=1,\dots,T$ do
7             if $\perp\neq e_{t}=\{u,v\}$ and $\max(d_{u},d_{v}))<\Lambda$ then
8                   Increment $d_{u},d_{v}$
9                   $M_{t}\leftarrow\textsc{EdgePrivateSublinearMatching.ProcessUpdate}(e_{t})$
10
11            else
[…]
13
14            Output $M_{t}$
15
16
Algorithm 4.2 Continual Release Node-DP Matching
Proof (Theorem 4.13).
Algorithm 4.2 first implements ContractionSparsify for insertion-only streams. Thus by Proposition 4.11, any node-neighboring graphs become $2\Lambda$ edge neighboring after sparsification. By a group privacy argument, the privacy and approximation guarantees follow directly from the guarantees of Algorithm 4.1 (Theorem 4.1) after adjusting the privacy parameter $\varepsilon\leftarrow\varepsilon/2\Lambda$. On top of the $O(\varepsilon^{-1}\eta^{-2}\Lambda\log^{2}(n)\log(\tilde{\alpha}))=O(\varepsilon^{-1}\eta^{-3}\tilde{\alpha}\log^{2}(n)\log(\tilde{\alpha}))$ space used by Algorithm 4.1, we also use an additional $O(n\log(\Lambda))=O(n\log(\eta^{-1}\tilde{\alpha}))$ space to store degree counters for each vertex.
The pseudocode for the proof of Theorem 4.14 is given in Algorithm 4.3.
1 Function Alg($\tilde{\alpha},\varepsilon,\eta,n,T$)
2       $E_{0}=\varnothing$
3
4      $\Lambda\leftarrow 5\left(1+5\eta^{-1}\right)2\tilde{\alpha}$
5       $k_{0}\leftarrow 0$
6       Initialize class SVT$(\varepsilon,2\Lambda,10\log_{1+\eta}(n))$   (Algorithm 3.1)
7       for incoming update $e_{t}$ at time $t=1,\dots,T$ do
8             $E_{t}\leftarrow E_{t-1}$
9             if $\perp\neq e=\{u,v\}$ and $\max(\deg_{t}(u),\deg_{t}(v))<\Lambda$ then
10                   $E_{t}\leftarrow E_{t}\cup\{e\}$
11            $\mu_{t}\leftarrow\mathtt{maximumMatchingSize}(V,E_{t})$   (Theorem B.1)
12             $\triangleright$ Compute exact matching size with any polynomial time, linear space algorithm
13             $k_{t}\leftarrow k_{t-1}$
14             while SVT.ProcessQuery$\left(\mu_{t},(1+\eta)^{k_{t}}\right)$ is “above” do
15                   $k_{t}\leftarrow k_{t}+1$
16
17            Output $(1+\eta)^{k_{t}}$
18
19
Algorithm 4.3 Continual Release Node-DP Matching
Proof (Theorem 4.14).
The algorithm which achieves the guarantees of Theorem 4.14 can be obtained from Algorithm 4.2 by replacing EdgePrivateSublinearMatching with a simple edge-DP insertion-only continual release algorithm as follows: At each timestamp $t$, compute the exact value of the sparsified graph using a static algorithm such as the Micali-Vazirani algorithm (Theorem B.1) which uses linear space. Then release an estimate as a power of $(1+\eta)$ using an SVT mechanism. Specifically, we release $(1+\eta)^{i}$ with the largest $i$ that the SVT query successfully returns that the matching size is above the threshold.
This algorithm is $\varepsilon$-edge DP and outputs a $\left(1+\eta,O(\varepsilon^{-1}\eta^{-2}\tilde{\alpha}\log^{2}(n))\right)$-approximation with high probability. By group privacy, adjusting the privacy parameter $\varepsilon^{\prime}\leftarrow\nicefrac{{\varepsilon}}{{2\Lambda}}$ yields the desired algorithm, and the error from the new setting of $\varepsilon^{\prime}$ appears in the additive error of the approximation (this error results from the SVT mechanism). We store at most $O(n\Lambda)=O(\eta^{-1}n\tilde{\alpha})$ edges, which corresponds to the space complexity. ∎
5 Implicit Vertex Cover in the Semi-Streaming Model
In this section, we design a node-DP vertex cover algorithm in the semi-streaming model that releases an implicit vertex cover (cf. Section 5.1) using the sparsification techniques derived in Section 4.2. We use the static edge DP implicit vertex cover algorithm (cf. Theorem 5.4) in [gupta2010differentially]. In order to obtain high probability guarantees, we use the private boosting framework derived in Appendix F that is adapted from the generalized private selection framework in [cohen2022generalized].
Theorem 5.1.
Let $\eta\in(0,1]$. Algorithm 5.1 is an $\varepsilon$-node DP vertex cover algorithm in the semi-streaming model with the following guarantees for an input graph $G$ with a given public bound $\tilde{\alpha}$. [^15] We can eliminate this assumption with an additional pass of the stream or notify the observer when the utility guarantees no longer hold in the one-pass setting. See Appendix E.
  1. (i)
    If $\tilde{\alpha}\geq\alpha$, where $\alpha$ is the maximum arboricity $\alpha$ over the stream, the algorithm outputs an implicit vertex cover (cf. Section 5.1) with cardinality at most
    $(1+\eta)\left(3+\frac{151\tilde{\alpha}}{\varepsilon}\right)\operatorname{OPT}(G)+O\left(\frac{\tilde{\alpha}\log(\nicefrac{{n}}{{\eta}})}{\varepsilon}\right)$
    with probability at least $1-1/\operatorname*{poly}(n)$.
  2. (ii)
    The space complexity is $O(\tilde{\alpha}n)$.
  3. (iii)
    The update, preprocessing, and postprocessing time complexities are polynomial.
5.1 Bounded Arboricity Sparsifiers for Vertex Cover
[solomon2021local] showed that any $\beta$-approximate vertex cover for $G_{\leq}^{\Lambda}$ plus the set of vertices $V_{>}^{\Lambda}$ is a $(\beta+\eta)$-approximate vertex cover for $G$ for any constant $\eta>0$. Recall that $G_{\leq}^{\Lambda}$ denotes the induced subgraph on all vertices with degree less than or equal to some bound $\Lambda$ (which we will set) and $V_{>}^{\Lambda}$ denotes the set of vertices with degree greater than $\Lambda$. We extend the ideas of [solomon2021local] to more general sparsifiers for vertex cover. Our argument holds in greater generality at the cost of an additional additive error of $\operatorname{OPT}(G)$ in the approximation guarantees.
Theorem 5.2.
Let $\eta>0$, $G$ be a graph of arboricity at most $\alpha$, and set
$\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha.$
Let $H\preceq G$ be any subgraph of $G$ such that $G_{\leq}^{\Lambda}\preceq H\preceq G$. Here $\preceq$ denotes the subgraph relation. If $C$ is a $(\beta,\zeta)$-approximate vertex cover of $H$, then $C\cup V_{>}^{\Lambda}$ is a cover of $G$ with cardinality at most
$(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$
Lemma 5.3 (Lemma 3.8 in [solomon2021local]).
Let $\eta>0$ and $G$ be a graph of arboricity at most $\alpha$ and set $\Lambda:=(1+\nicefrac{{1}}{{\eta}})\cdot 2\alpha$. If $C$ is any vertex cover for $G$, and $U_{>}:=V_{>}\setminus C$, then $\lvert U_{>}\rvert\leq\eta\lvert C\rvert$.
Remark.
Lemma 3.8 in [solomon2021local] holds as long as any induced subgraph of $G$ has average degree at most $2\alpha$, which is certainly the case for graphs of arboricity at most $\alpha$. Moreover, [solomon2021local] stated their result for
$V_{\geq}=\{v\in V:\deg(v)\geq\Lambda\},\qquad U_{\geq}=V_{\geq}\setminus C.$
But since $U_{>}\subseteq U_{\geq}$, the result clearly still holds. In order to remain consistent with our matching sparsifier, we state their result with $V_{>}$ and $U_{>}$.
Proof (Theorem 5.2).
Any edge in $H$ is covered by $C$, which is a $(\beta,\zeta)$-approximate vertex cover of $H$, and any edge that is not in $H$ is covered by $V_{>}^{\Lambda}$. Thus $C\cup V_{>}^{\Lambda}$ is indeed a vertex cover. By Lemma 5.3, $\lvert V_{>}^{\Lambda}\setminus C^{*}\rvert\leq\eta\lvert C^{*}\rvert$ where $C^{*}$ is a minimum vertex cover of $G$. Then,
$\lvert C\cup V_{>}^{\Lambda}\rvert$ $\leq\lvert C\rvert+\lvert V_{>}^{\Lambda}\rvert$ (42)
$\leq\beta\operatorname{OPT}(H)+\zeta+\lvert V_{>}^{\Lambda}\cap C^{*}\rvert+\lvert V_{>}^{\Lambda}\setminus C^{*}\rvert$ (43)
$\leq\beta\operatorname{OPT}(G)+\zeta+\operatorname{OPT}(G)+\eta\operatorname{OPT}(G)$ (44)
$=(1+\beta+\eta)\operatorname{OPT}(G)+\zeta.$ (45)
Equation 44 follows by Lemma 5.3 and since $\lvert V_{>}^{\Lambda}\cap C^{*}\rvert\leq\operatorname{OPT}(G)$ due to $C^{*}$ being an optimal cover for $G$. ∎
Using the above proof, we can set our vertex cover sparsifier $H$ to be the same as our matching sparsifier given in Section 4.3. It is guaranteed that $G_{\leq}^{\Lambda}\subseteq H$ since if an edge $\{u,w\}$ has $\deg(u)\leq\Lambda$ and $\deg(w)\leq\Lambda$ then the edge is guaranteed to be in the first $\Lambda$ edges of the ordering $\pi$ for both of its endpoints.
1 Function Alg($\tilde{\alpha},\eta,\varepsilon,n,T$)
2       $E_{0}\leftarrow\varnothing$
3       $\xi\leftarrow\sqrt{\frac{(2^{7}+2^{-2})\tilde{\alpha}}{\varepsilon}}$
4       $\Lambda\leftarrow\left(1+\xi^{-1}\right)2\tilde{\alpha}$
5       for incoming update $e_{t}$ at time $t=1,2,\dots,T$ do
6             $E_{t}\leftarrow E_{t-1}$
7             if $\perp\neq e_{t}=\{u,v\}$ and $\max(\deg_{t}(u),\deg_{t}(v))<\Lambda$ then
8                   $E_{t}\leftarrow E_{t}\cup\{e\}$
9
10
11      $\pi\leftarrow\textsc{NodePrivateVC}(V,E_{T},\tilde{\alpha},\eta,\varepsilon)$   (Theorem 5.6)
12       Return $(\pi,\Lambda)$
Algorithm 5.1 Continual Release Node DP Vertex Cover
The node-DP algorithm to estimate the value of a maximum matching immediately yields a node-DP algorithm for estimating the size of a minimum vertex cover. In fact, it suffices to compute the size of a greedy maximal matching, which has a coupled global sensitivity of 1 with respect to node neighboring graphs (cf. Theorem B.2). This is then a $2$-approximation for the size of a minimum vertex cover. However, it may be desirable to output actual solutions to the vertex cover problem. [gupta2010differentially] demonstrated how to output an implicit solution through an ordering of the vertices. It may be tempting to plug this into the same technique we adopt throughout this work to obtain an algorithm in the continual release model. However, it is unclear how the utility of the implicit cover degrades in between updates. Thus we present an algorithm for the semi-streaming model where we are allowed near-linear space and output a solution at the end of the stream.
The below theorem returns an implicit solution where every edge is considered an entity and can determine, using the released implicit solution, which of its endpoints covers it.
Theorem 5.4 (Theorem 5.1, Theorem 5.2 in [gupta2010differentially]).
Fix $\beta\in(0,1)$. There is a polynomial time $\varepsilon$-edge DP algorithm $\textsc{EdgePrivateVC}_{\beta}$ that outputs an implicit vertex cover with expected cardinality at most
$\left(2+\frac{16}{\varepsilon}\right)\operatorname{OPT}.$
By Markov’s inequality, running EdgePrivateVC guarantees a solution with cardinality at most
$(1+\eta)\left(2+\frac{16}{\varepsilon}\right)\operatorname{OPT}$
with probability at least $1-\frac{1}{1+\eta}=\frac{\eta}{1+\eta}$ for any $\eta\in(0,1]$. Then running the algorithm $O(\eta^{-1}\log(n))$ times ensures the algorithm outputs such a solution with probability at least $1-1/\operatorname*{poly}(n)$. In Appendix F, we show how we can perform this boosting process privately using the private selection algorithm from [cohen2022generalized].
Implicit Solutions for Sparsification
The implicit vertex cover from [gupta2010differentially] is an ordering of the vertices where each edge is covered by the earlier vertex in the ordering. Such an implicit ordering allows for each edge to determine which of its endpoints covers it given the public ordering.
In our setting, if we run EdgePrivateVC on a sparsified graph $H$, the ordering we output is with respect to $H$ and some edges that were deleted from $G$ to obtain $H$ may not be covered by the intended solution. Thus we also output the threshold we input into the sparsifier so that an edge is first covered by a high-degree endpoint if it has one, and otherwise by the earlier vertex in the ordering. More succinctly, our implicit solution is a triple
$\left(\pi,\Lambda\right)\in S_{n}\times\mathbb{Z}_{\geq 0}$
where $\pi$ is an ordering of the vertices and $\Lambda$ is the threshold used to perform the sparsification.
As shown in Lemma 5.3, adding all high-degree vertices of $G$ does not incur much overhead in the approximation guarantee. As in the setting given by [gupta2010differentially], all edges know which endpoint covers it using our implicit solution.
Computing an Implicit Solution in the Semi-Streaming Setting
We give the pseudocode for our algorithm in Algorithm 5.1. We assume a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the final graph $G_{T}$ is given to us. Our algorithm is always private but utility is only guaranteed when $\tilde{\alpha}\geq\alpha$. Algorithm 5.1 in Algorithm 5.1 runs a boosted version of the algorithm of [gupta2010differentially] on our sparsified graph with guarantees given in Theorem 5.4. We first prove the following lemma which we use in the proof of the privacy of our algorithm.
Lemma 5.5.
Let $G\sim G^{\prime}$ be edge-neighboring graphs and $\pi$ a fixed permutation of the vertices (in both graphs). Assign each edge to the endpoint that occurs earlier in the permutation. Then the number of vertices with at least one assigned edge differs by at most 1 between $G,G^{\prime}$.
This lemma shows that computing the size of a vertex cover with respect to a fixed permutation has edge-sensitivity 1.
Proof (Lemma 5.5).
Let $v_{1},\dots,v_{n}$ denote the vertex ordering and suppose we perform the assignment procedure in the statement. Then a vertex is assigned an edge if and only if it has a neighbor that is placed later in the ordering, i.e. its out-degree with respect to the ordering is non-zero. Let the differing edge be between some $v_{i},v_{j}$ where $i<j$. Then the out-degree of every vertex is identical except for $v_{i}$, whose out-degree differs by $1$. Hence the number of vertices with at least one assigned edge differs by at most $1$ between $G,G^{\prime}$. ∎
We now prove our main theorem.
Theorem 5.6.
Let $\eta,\gamma\in(0,1],\xi>0$ and $\varepsilon>0$. Let $\tilde{\alpha}$ be a public bound for the arboricity of the input graph $G$ and $H=\textsc{ContractionSparsify}(G,\Lambda)$ be the sparsified graph using a threshold
$\Lambda=(1+\xi^{-1})2\tilde{\alpha}.$
There is an algorithm NodePrivateVC for vertex cover such that given $H,\tilde{\alpha},\eta,\varepsilon$, outputs an implicit solution
$\left(\pi,\Lambda\right)\in S_{n}\times\mathbb{Z}_{\geq 0}$
with the following guarantees:
  1. (i)
    The algorithm is $\varepsilon$-node DP,
  2. (ii)
    $\pi$ induces an implicit vertex cover $C_{\pi}$ for $H$ with size at most
    $(1+\eta)\left(2+\frac{64(1+\gamma)}{\varepsilon}\Lambda\right)\operatorname{OPT}(H)+O(\varepsilon^{-1}\gamma^{-2}\Lambda\log(\eta^{-1}n)),$
    with probability at least $1-1/\operatorname*{poly}(n)$.
  3. (iii)
    $C_{\pi}\cup V_{>}^{\Lambda}(G)$ forms an implicit vertex cover of $G$ with size at most
    $(1+\eta)\left(3+\xi+\frac{64(1+\gamma)}{\varepsilon}\Lambda\right)\operatorname{OPT}(G)+O(\varepsilon^{-1}\gamma^{-2}\Lambda\log(\eta^{-1}n))$
    with probability at least $1-1/\operatorname*{poly}(n)$.
Proof (Theorem 5.6).
Fix $\eta>0$ and define $\Lambda=(1+\nicefrac{{1}}{{\eta}})2\tilde{\alpha}$. Let $G_{<}^{\Lambda}\preceq H\preceq G$ denote the sparsified graph obtained with inputs $G,\Lambda$.
Let NodePrivateVC be the algorithm that runs (cf. Theorem F.2)
[…]
where EdgePrivateVC is called [^16] We note the constants are arbitrarily chosen to obtain $\nicefrac{{\varepsilon}}{{(2+\gamma)}}$-node DP via composition and group privacy so that we can apply the boosting algorithm. with edge privacy parameter $\nicefrac{{\varepsilon}}{{(2+2\gamma)(2\Lambda)}}$ and augmented with a $(\nicefrac{{\gamma\varepsilon}}{{(2+\gamma)(2+2\gamma)(2\Lambda)}})$-edge DP estimate of the size of $C_{\pi}$
$\lvert C_{\pi}\rvert+{\textsf{Lap}}\left(\frac{\gamma\varepsilon}{(2+\gamma)(2+2\gamma)2\Lambda}\right).$
Note that by Lemma 5.5 and Proposition 4.11, any node-neighboring graphs $G^{\prime}\sim G$ differ in at most $2\Lambda$ edges after sparsification. Hence the augmented call to EdgePrivateVC is overall $(\nicefrac{{\varepsilon}}{{(2+\gamma)}})$-node DP.
Proof of (i)
The privacy guarantees follow from the edge-DP algorithm given in Theorem 5.4 and our private boosting framework given in Theorem F.2.
Proof of (ii)
By group privacy and Theorem 5.4, running EdgePrivateVC on $H$ as described above yields a ($\nicefrac{{\varepsilon}}{{(2+\gamma)}}$)-node DP implicit vertex cover $C_{\pi}$ with expected size at most
$\mu:=\mathop{{\mathbb{E}}}\left[\lvert C\rvert\right]\leq\left(2+64(1+\gamma)\varepsilon^{-1}\Lambda)\right)\operatorname{OPT}(H).$
An application of Markov’s inequality yields the guarantee that $\mathbf{P}[\lvert C\rvert>(1+\eta)\mu]\leq\frac{1}{1+\eta}$. Then the probability of $\lvert C\rvert\leq(1+\eta)\mu$ is at least $1-\frac{1}{1+\eta}=\frac{\eta}{1+\eta}=\Omega(\eta)$.
We incur additional additive error from the noisy estimates of $\lvert C_{\pi}\rvert$ in the private boosting framework. By a union bound, each of the $\tau$ estimates of the size of $\lvert C_{\pi}\rvert$ has additive error at most $O(\varepsilon^{-1}\Lambda\log(\tau n))$ with probability at least $1-1/\operatorname*{poly}(n)$.
Proof of (iii)
By Lemma 5.3, $\lvert V_{\geq}^{\Lambda}\setminus C\rvert\leq\xi\lvert C\rvert$ for any vertex cover $C$ of $G$. Hence we can simply take the union of all vertices with degree greater than $\Lambda$ with a cover for $H$ and obtain a vertex cover for $G$. By Theorem 5.2, adding all high-degree vertices and considering the resultant as a vertex cover of $G$ increases the size of any solution by at most $\left(1+\xi\right)\operatorname{OPT}(G)$. ∎
5.2 Putting it Together
Proof (Theorem 5.1).
Algorithm 5.1 implements ContractionSparsify (Proposition 4.11) for insertion-only streams. Thus the privacy and approximation guarantees follow directly from Theorem 5.6. The space usage follows since the degree of each stored vertex is at most $\Lambda=O(\tilde{\alpha})$. The approximation guarantees follow from the choice of
$\gamma=2^{-9}\text{ and }\xi=\sqrt{\frac{2^{7}(1+\gamma)\tilde{\alpha}}{\varepsilon}}<11.33\sqrt{\frac{\tilde{\alpha}}{\varepsilon}}.$
We get the final approximation of:
$(1+\eta)\left(3+\xi+\frac{64(1+\gamma)}{\varepsilon}\Lambda\right)\operatorname{OPT}(G)+O(\varepsilon^{-1}\gamma^{-2}\Lambda\log(\eta^{-1}n))$
$=(1+\eta)\left(3+\frac{2\cdot 64(1+\gamma)\tilde{\alpha}}{\varepsilon}+2\cdot\sqrt{\frac{2^{7}(1+\gamma)\tilde{\alpha}}{\varepsilon}}\right)\operatorname{OPT}(G)+O(\varepsilon^{-1}\tilde{\alpha}\log(\eta^{-1}n))$
$<(1+\eta)\left(3+\frac{(2^{7}+23)\tilde{\alpha}}{\varepsilon}\right)\operatorname{OPT}(G)+O(\varepsilon^{-1}\tilde{\alpha}\log(\eta^{-1}n)).$
6 Densest Subgraph
In this section, we focus on the densest subgraph problem and provide the first differentially private algorithm for densest subgraph in the continual release model using space sublinear in the total number of edges in the graph.
Theorem 6.1 (Sublinear Space Private Densest Subgraph).
Fix $\eta\in(0,1]$. Algorithm 6.2 is an $\varepsilon$-edge differentially private algorithm for the densest subgraph problem in the continual release model for insertion-only streams. The algorithm returns a set of vertices whose induced subgraph is a $\left(2+\eta,\varepsilon^{-1}\eta^{-1}\log^{2}(n)\right)$-approximation of the densest subgraph in $G_{t}$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$. The maximum space used is $O\left(\varepsilon^{-1}\eta^{-2}n\log^{2}(n)\right)$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$.
We can reduce the multiplicative error to $(1+\eta)$ at the cost of increasing the space usage by a $\operatorname*{poly}(\log(n))$ factor.
Theorem 6.2 (Sublinear Space Private Densest Subgraph).
Fix $\eta\in(0,1]$. There exists an $\varepsilon$-edge differentially private algorithm for the densest subgraph problem in the continual release model for insertion-only streams. The algorithm returns a set of vertices whose induced subgraph is a $\left(1+\eta,\varepsilon^{-1}\eta^{-1}\log^{5}(n)\right)$-approximation of the densest subgraph in $G_{t}$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$. The maximum space used is $O\left(\varepsilon^{-1}\eta^{-2}n\log^{5}(n)\right)$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$.
6.1 Algorithm Description
In this section, we begin with an $\varepsilon$-edge differentially private algorithm in the continual release model for insertion-only streams that releases, in addition to the approximate density of the densest subgraph, a differentially private set of vertices whose induced subgraph is an approximation of the maximum densest subgraph. We now introduce a sampling method that uses $O\left(\varepsilon^{-1}\eta^{-2}n\log n\right)$ total space [^17] Here, $\eta$ is the factor used in the multiplicative approximation. over the course of the stream allowing us to match the space bounds (up to factors of $\operatorname*{poly}(\log n)$ and $1/\varepsilon$) of the best non-private algorithms for approximate densest subgraphs while also matching the multiplicative approximation factor.
Algorithm Intuition.
We revise the algorithms of [mcgregor2015densest] and [EHW16] to the insertion-only continual release setting. On a high-level, our algorithm maintains a sample of edges over time and releases a DP set of vertices by running a black-box DP densest subgraph algorithm (e.g., Theorem C.2) on the subgraph induced by the sample. At every timestep $t\in[T]$, an edge $e_{t}$ is sampled with probability $p$ — the sampling probability is initialized to 1 as in the beginning we can afford to store every edge. We privately check whether the number of edges seen so far exceeds a certain threshold using a sparse vector technique (SVT) query and adjust the sampling probability $p$ accordingly. In order to avoid privacy loss that grows linearly in $T$, we do not invoke the black-box DP densest subgraph algorithm at every timestep and instead invoke it only if the current density of the sample exceeds a certain threshold using another SVT query. The pseudocode for our algorithm is given in Algorithm 6.1 and Algorithm 6.2.
1 Class PrivateDSG($\varepsilon$, $\eta$, $n$, $T$)
2       Maintain sampled edges $X\leftarrow\varnothing$.
3       Set $m^{\prime}\leftarrow\frac{c_{3}n\log^{2}(n)}{\varepsilon\eta^{2}}$ using a large enough constant $c_{3}>0$.
4       Initialize empty sampling hashmap $H$ to store edges and their associated random weight.
5       Initialize class SVT($\varepsilon,1,c_{5}\log_{1+\eta}(n)$).   (Algorithm 3.1)
6       Function SampleEdge($e_{t}$, $m$, $\varepsilon$)
7             if SVT.ProcessQuery$(m,m^{\prime})$ is “above” then
8                   $m^{\prime}\leftarrow(1+\eta)\cdot m^{\prime}$.
9
10                  $p\leftarrow\min\left(1,\frac{c_{4}n\log^{2}(n)}{\varepsilon\cdot\eta^{2}m^{\prime}}\right)$.
11                   for each edge $e\in H$ do
12                         if $H[e]\leq p$ then
13                               Keep $e$ in $X$.
14                        else
15                               Remove $e$ from $X$ and $H$.
16
17
18            if $e_{t}\neq\perp$ then
19                   Sample $h_{e_{t}}\sim U[0,1]$ uniformly at random in $[0,1]$.
20                   if $h_{e_{t}}\leq p$ then
21                         Store $X\leftarrow X\cup\{e_{t}\}$.
22                         Add $H[e_{t}]=h_{e_{t}}$.
23
24
25      Function GetPrivateApproxDensestSubgraph($\varepsilon$)
[…]
27      Function GetNonPrivateDensity()
28             Return $(\textsc{ExactDensity}(X),p)$.   (Theorem D.1)
29
30
Algorithm 6.1 Data Structure for Densest Subgraph in Adaptive Insertion-Only Streams
Detailed Algorithm Description.
We first define a data structure PrivateDSG (see Algorithm 6.1) for maintaining an approximate densest subgraph in insertion-only streams. This data structure takes as input an accuracy parameter $\eta$, the number of nodes $n$, and a bound on the stream size $T$ and uses as a black-box PrivateDensestSubgraph (an $\varepsilon$-DP static densest subgraph algorithm) and Densest-Subgraph (a non-private static densest subgraph algorithm). This data structure has three procedures: an update procedure (SampleEdge), a procedure for getting a private densest subgraph (GetPrivateApproxDensestSubgraph), and a procedure for returning the non-private exact density of the sampled subgraph (GetNonPrivateDensity).
Our main algorithm (see Algorithm 6.2) uses the data structure above and performs its initialization based on the input accuracy parameter $\eta$, the number of nodes $n$, the privacy parameter $\varepsilon$, and an upper bound $T$ on the total number of updates in the stream.
31 Initialize counter of number of edges $m\leftarrow 0$.
32 Initialize privacy parameters $\varepsilon_{1}\leftarrow\frac{\varepsilon}{3c_{2}\log_{1+\eta}(n)},\varepsilon_{2}\leftarrow\varepsilon/3$.
33 Initialize PrivateDSG($\varepsilon_{1},\eta$, $n$, $T$).
34 Initialize density threshold $L\leftarrow\frac{(1+\eta)c_{1}\cdot\log^{2}(n)}{\varepsilon\eta}$.
35 Initialize class SVT$(\varepsilon_{2},1,c_{2}\log_{1+\eta}(n))$.   (Algorithm 3.1)
36 for each new update $e_{t}$ do
37       if $e_{t}\neq\bot$ then
38             $m\leftarrow m+1$.
39      PrivateDSG.SampleEdge($e_{t},m,\varepsilon_{2}$).
[…]
41       $S\leftarrow V$.
42       if SVT.ProcessQuery$(D,p\cdot L)$ is “above” then
43             $L\leftarrow(1+\eta)\cdot L$.
44             $S\leftarrow\textsc{PrivateDSG.GetPrivateApproxDensestSubgraph}(\varepsilon_{1})$.
45      Release $S$.
Algorithm 6.2 Algorithm for Densest Subgraph in Adaptive Insertion-Only Streams
We perform the following initializations:
  1. (i)
    an instance of PrivateDSG($\varepsilon$, $\eta$, $n$, $T$) (Algorithm 6.2),
  2. (ii)
    a counter for the number of edge insertions we have seen so far in our stream (Algorithm 6.2),
  3. (iii)
    the privacy parameters for the different parts of our algorithm (Algorithm 6.2),
  4. (iv)
    and the initial density threshold for the density for returning a new private densest subgraph solution (Algorithm 6.2) where $c_{1}>0$ is a constant.
The initial cutoff for the density is equal to our additive error since we can just return the entire set of vertices as long as the density of the densest subgraph is (approximately) less than our additive error.
We now receive an online stream of updates (which can be empty $\bot$) one by one. The $t$-th update is denoted $e_{t}$ (Algorithm 6.2). For each update, we first check whether the update is an edge insertion or $\bot$ (Algorithm 6.2); if it is not $\bot$, then we increment $m$ (Algorithm 6.2). We then call PrivateDSG.SampleEdge($e_{t},m,\varepsilon_{2}$) which decides how to sample the edge (Algorithm 6.2).
The SampleEdge procedure is within the PrivateDSG class which maintains the following:
  1. (i)
    a set $X$ (Algorithm 6.1) of sampled edges in the stream,
  2. (ii)
    an estimate, $m^{\prime}$, of the number of edges seen so far in the stream (Algorithm 6.1),
  3. (iii)
    and the probability $p$ by which the current edge in the stream is sampled (Algorithm 6.1).
The initial probability that we sample an edge is $1$ since we have not seen many edges and can keep all of them within our space bounds.
In order to sample according to the desired (private) probability, we first privately check whether our current number of edges we have seen exceeds the threshold $m^{\prime}$ using the sparse vector technique (Algorithm 3.1). Our sparse vector technique function (Algorithm 6.1) is initialized with
  1. (i)
    the privacy parameter $\varepsilon_{2}$,
  2. (ii)
    the sensitivity of the query (which is $1$ in our setting),
  3. (iii)
    and the maximum number of successful queries (we perform queries until we’ve reached the end of the stream or we exceed the successful queries threshold).
Then, the SVT query is run with the query, $m$, (which is the current number of edges we’ve seen so far) and the threshold, $m^{\prime}$ (Algorithm 6.1). If the SVT query passes, then we update $m^{\prime}$ to a larger threshold (Algorithm 6.1) and also update the probability of sampling in terms of the new $m^{\prime}$ (Algorithm 6.1) (using constant $c_{4}$). We then sample the edges by sampling a value $h_{e_{t}}$ uniformly from $[0,1]$ (Algorithm 6.1). If $h_{e_{t}}\leq p$ (Algorithm 6.1), then we store $e_{t}$ in $X$ (Algorithm 6.1) and the value $h_{e_{t}}$ in our hashmap $H$ (Algorithm 6.1). Whenever $p$ changes, we also have to resample the edges in $X$; to do this, we keep the edge in $X$ if $H[e_{t}]\leq p$ and remove $e_{t}$ otherwise (Algorithms 6.1, 6.1 and 6.1). This ensures that the marginal probability of sampling an edge at any timestamp $t$ is $p$.
After we have sampled our edges, we now need to obtain the non-private density of our current subgraph (Algorithm 6.2). If the non-private density exceeds our density threshold $L$ via private SVT (Algorithm 6.2), then we increase our threshold (Algorithm 6.2) and use our private densest subgraph algorithm to return a private approximate densest subgraph (Algorithm 6.2). We release our stored private graph (Algorithm 6.2) for all new updates until we need to compute a new private densest subgraph.
The densest subgraph algorithm that we use to obtain a differentially private densest subgraph can be any existing static edge-DP algorithm for densest subgraph like Theorem C.1 which gives a $(2,O(\varepsilon_{1}^{-1}\log(n)))$-approximation.
6.2 Privacy Guarantee
In this section, we prove the privacy guarantees of our algorithm.
Lemma 6.3.
Algorithm 6.2 is $\varepsilon$-DP in the continual release model on edge-neighboring insertion-only streams.
Proof (Lemma 6.3).
By the definition of edge-neighboring streams (Definition 3.3), let ${\mathcal{S}}=(e_{1},\ldots,e_{{T}}),{\mathcal{S}}^{\prime}=(e^{\prime}_{1},\ldots,e^{\prime}_{{T}})$ be neighboring streams of edges, i.e., there exists $t^{*}\in[{T}]$ such that $e_{t^{*}}\neq e^{\prime}_{t^{*}}$ and $e^{\prime}_{t^{*}}=\bot$. Note that $e_{t}$ for any $t\in[T]$ can either be an edge $\{u,v\}$ or $\bot$. We define the event ${\mathcal{P}}$ where the probability for sampling $e_{t}$ from the stream ${\mathcal{S}}$ is set to $p(e_{t})=p_{t}$ for every $t\in[{T}]$. Similarly define ${\mathcal{P}}^{\prime}$ where the probability for sampling $e^{\prime}_{t}$ from the stream ${\mathcal{S}}^{\prime}$ is set to $p(e^{\prime}_{t})=p_{t}$ for every $t\in[{T}]$. Although we only use the probability to sample an update $e_{t}$ if $e_{t}\neq\bot$, our sampling of the update depends on our setting of $p_{t}$ which we show is the same across the two streams with roughly equal probability.
{adjustwidth}
3em0em
Claim 6.4.
$\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon_{2}}$
Proof.
Consider any output answer vector $\mathbf{a}\in\{\text{``above''},\text{``below''},\text{``abort''}\}^{T}$ to SVT. First, observe that the SVT (in SampleEdge of Algorithm 6.1) is $\varepsilon_{2}$-DP by Theorem 3.13, therefore $\frac{\mathbf{P}[SVT(m({\mathcal{S}}))=\mathbf{a}]}{\mathbf{P}[SVT(m({\mathcal{S}}^{\prime}))=\mathbf{a}]}\leq e^{\varepsilon_{2}}$ where $m(S)$ is the number of edges in $S$. Now, observe that the probability $p$ of sampling only changes whenever SVT accepts. Since privacy is preserved after post-processing, this implies the statement in the claim. ∎
{adjustwidth}
3em0em
Claim 6.5.
Define $I_{-t^{*}}^{(t)}=(I^{(t)}_{1},\ldots,I^{(t)}_{t})$ for stream ${\mathcal{S}}$ where for $t^{*}\neq j\leq t$,
$I_{j}^{(t)}$ =$=\begin{cases}1&e_{j}\text{ is in the sample }X^{(t)}\\ 0&\text{otherwise}.\end{cases}$ (46)
Define ${I^{\prime}}_{-t^{*}}^{(t)}=({I^{\prime}}^{(t)}_{1},\ldots,{I^{\prime}}^{(t)}_{t})$ for stream ${\mathcal{S}}^{\prime}$ analogously. Let $\mathbf{b}^{(t)}$ be a $\{0,1\}$-vector with the same dimension as $I_{-t^{*}}^{(t)}$ and ${I^{\prime}}_{-t^{*}}^{(t)}$. Also let ${\mathcal{I}}=(I_{-t^{*}}^{(1)},\ldots,I_{-t^{*}}^{({T})})$ and ${\mathcal{I}}^{\prime}=(I_{-t^{*}}^{\prime(1)},\ldots,I_{-t^{*}}^{\prime({T})})$. Then for any choice of $\mathbf{b}^{(t)}$’s,
$\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}=\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})}),{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})}),{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon_{2}}$ (47)
Proof.
By definition, ${\mathcal{P}}$ and ${\mathcal{P}}^{\prime}$ defines the sampling probability for all elements in ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ to be $p_{t}$. Hence, it holds that $\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}=1$. By 6.4, $\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}\leq e^{\varepsilon_{2}}$ and our claim follows. ∎
Let ${\mathcal{Q}}$ be the event that $({\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(T)}),{\mathcal{P}})$ and define ${\mathcal{Q}}^{\prime}$ analogously.
{adjustwidth}
3em0em
Claim 6.6.
Let $\mathbf{a}\in\{\text{``above''},\text{``below''},\text{``abort''}\}^{T}$ be a sequence of SVT answers, $D(\cdot)$ denote the function computing the sequence of densities, and $X(\cdot)$ be the produced samples from the stream. We have
$\frac{\mathbf{P}[SVT(D(X({\mathcal{S}})))=\mathbf{a}|{\mathcal{Q}}]}{\mathbf{P}[SVT(D(X({\mathcal{S}}^{\prime})))=\mathbf{a}|{\mathcal{Q}}^{\prime}]}$ $\leq e^{\varepsilon_{2}}$ (48)
Proof.
Note that conditioned on the events ${\mathcal{Q}}$ and ${\mathcal{Q}}^{\prime}$, the samples $X({\mathcal{S}})$ and $X({\mathcal{S}}^{\prime})$ differ by at most $1$ edge, specifically the edge $e_{t^{*}}$. This is due to the fact that the same $e_{t}$ are sampled from both ${\mathcal{S}}$ and ${\mathcal{S}}^{\prime}$ conditioned on ${\mathcal{Q}}$ and ${\mathcal{Q}}^{\prime}$ except for $e_{t^{*}}$, and by definition of ${\mathcal{S}}\sim{\mathcal{S}}^{\prime}$, for only one $t^{*}\in[{T}]$, it holds that $e_{t^{*}}\neq e_{t^{*}}^{\prime}$. Therefore the sensitivity of exactly computing the density on the sample is given by $|D(X({\mathcal{S}}))-D(X({\mathcal{S}}^{\prime}))|\leq 1$. And SVT is $\varepsilon_{2}$-DP conditioned on these events by Theorem 3.13. ∎
By the chain rule of conditional probability, putting the above claims together gives us
$\frac{\mathbf{P}[SVT(D(X({\mathcal{S}})))=\mathbf{a}|{\mathcal{Q}}]}{\mathbf{P}[SVT(D(X({\mathcal{S}})))=\mathbf{a}|{\mathcal{Q}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{I}}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}]}{\mathbf{P}[{\mathcal{I}}^{\prime}=(\mathbf{b}^{(1)},\ldots,\mathbf{b}^{({T})})|{\mathcal{P}}^{\prime}]}\cdot\frac{\mathbf{P}[{\mathcal{P}}]}{\mathbf{P}[{\mathcal{P}}^{\prime}]}$ $\leq e^{2\varepsilon_{2}}=e^{2\varepsilon/3}$ (49)
Lastly, we call on the private densest subgraph algorithm at most $c_{2}\log_{1+\eta}(n)$ times, and since the privacy budget assigned to this procedure is $\varepsilon_{1}=\frac{\varepsilon}{3c_{2}\log_{1+\eta}(n)}$, by sequential composition, this means that this operation is $(\nicefrac{{\varepsilon}}{{3}})$-DP.
Hence combining Equation 49 and the observation about the private densest subgraph algorithm operation being $(\nicefrac{{\varepsilon}}{{3}})$-DP gives us our main theorem statement by sequential composition. ∎
Remark.
We remark that although we are running a DP algorithm on a sample which may imply some privacy amplification bounds, these privacy savings only apply to our case if the sampling probability of the edges decreases per time step. Since we cannot guarantee that this will be the case for all input graphs, the privacy amplification bounds do not improve the worst-case privacy guarantees.
6.3 Approximation Guarantees
In this section, we prove the approximation factor of the approximate densest subgraph we obtain via our sampling procedure. We first prove that our sampling procedure produces, with high probability, a set of edges $X=X_{t}\subseteq E_{t}$ at time $t$ such that $|\mathrm{OPT}(G_{t})-\operatorname{den}_{G_{t}}(\mathrm{OPT}(X_{t}))|\leq\eta\cdot\mathrm{OPT}(G_{t})$ where $\mathrm{OPT}(G_{t})$ is the density of the densest subgraph in the dynamic graph $G_{t}$ at time $t$ and $\operatorname{den}_{G_{t}}(\mathrm{OPT}(X_{t}))$ is the density of the densest subgraph in the sampled graph $G[X_{t}]$ but with respect to the full graph $G_{t}$. Then, we show that the approximate solution returned by the private static densest subgraph subroutine applied to the sub-sampled graph translates to an approximate solution in the original graph. In the following, we assume full independence in each of the sampled values for each edge (which is guaranteed by our algorithm).
Lemma 6.7.
Fix $\eta\in(0,1]$. For every $t\in[T]$, Algorithm 6.2 returns a set of vertices $V^{*}\subseteq V$ where the induced subgraph $G_{t}[V^{*}]$ has density $\operatorname{den}_{G_{t}}(V^{*})\geq\frac{\operatorname{den}_{G_{t}}(V_{\mathrm{OPT}})}{(2+\eta)}-O\left(\frac{\log^{2}n}{\varepsilon\eta}\right)$ (where $V_{\mathrm{OPT}}$ is the set of vertices in a densest subgraph at time $t$ consisting of all edges seen in the stream so far), with high probability.
Proof.
For the first part of this proof, we take inspiration from the techniques given in [EHW16, Lemma 2.3], although our proof differs in several respects due to the noise resulting from our private mechanisms.
Sub-Sampling Concentration.
Let $x_{e}$ be the random variable indicating whether edge $e$ exists in $X$ and $p$ be the probability of sampling the edges. Fix an arbitrary non-empty subset of vertices $\varnothing\neq V^{\prime}\subseteq V$. The number of edges in $X[V^{\prime}]$ is given by $\lvert X[V^{\prime}]\rvert=\sum_{e\in X[V^{\prime}]}1=\sum_{e\in G_{t}[V^{\prime}]}x_{e}$. We use $G_{t}$ to denote the graph on edges inserted in updates $u_{1},\ldots,u_{t}$ at timestep $t$. Then, we know the expectation of $\operatorname{den}(X[V^{\prime}])$ is
$\mu=\mathbb{E}[\operatorname{den}(X[V^{\prime}])]$ $=\mathbb{E}\left[\frac{\sum_{e\in G_{t}[V^{\prime}]}x_{e}}{|V^{\prime}|}\right]=\frac{\mathbb{E}\left[\sum_{e\in G_{t}[V^{\prime}]}x_{e}\right]}{|V^{\prime}|}=\frac{\sum_{e\in G_{t}[V^{\prime}]}p}{|V^{\prime}|}$ (50)
$=p\cdot\frac{\sum_{e\in G_{t}[V^{\prime}]}1}{|V^{\prime}|}=p\cdot\operatorname{den}_{G_{t}}(V^{\prime}),$ (51)
and the expectation of $|X[V^{\prime}]|$ is also
$\mu_{|X[V^{\prime}]|}=\mathbb{E}[|X[V^{\prime}]|]=\mathbb{E}\left[\sum_{e\in G_{t}[V^{\prime}]}x_{e}\right]=p\cdot|G_{t}[V^{\prime}]|,$ (52)
where $|G_{t}[V^{\prime}]|$ denotes the number of edges in the induced subgraph of $G_{t}[V^{\prime}]$.
Since all of the variables $x_{e}$ are independent, we have by the multiplicative Chernoff bound (Theorem 3.1),
$\mathbf{P}\left[|X[V^{\prime}]|\geq(1+\psi)\cdot\mu_{|X[V^{\prime}]|}\right]$ $\leq\exp\left(-\frac{\psi^{2}\mu_{|X[V^{\prime}]|}}{2+\psi}\right)$ (53)
$\leq\exp\left(-\frac{\psi^{2}\mu_{|X[V^{\prime}]|}}{14}\right)$ When $\psi\in(0,12]$ (54)
$\leq\exp\left(-\frac{\psi^{2}\cdot p\cdot|G_{t}(V^{\prime})|}{14}\right).$ (55)
Consider the private estimate $m_{t}^{\prime}$ of $m_{t}$ stored in Algorithm 6.1 of Algorithm 6.1, where $m_{t}$ is the number of true edges in the graph. We add $O(T)$ instances of ${\textsf{Lap}}(O(\varepsilon/\log_{1+\eta}(n)))$ noise in the SVT subroutine (Algorithm 6.2). By a union bound over $T=\operatorname*{poly}(n)$, all such noises are of order $O(\varepsilon^{-1}\log_{1+\eta}(n)\log(n))=O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))$ with probability $1-1/\operatorname*{poly}(n)$. Hence $m_{t}-O\left(\frac{\log^{2}n}{\varepsilon\eta}\right)\leq m^{\prime}_{t}\leq(1+\eta)\cdot m_{t}+O\left(\frac{\log^{2}n}{\varepsilon\eta}\right)$, with high probability. Note that by the choice of initialization of $m^{\prime}$ (Algorithm 6.1), we sub-sample edges only once $m_{t}\geq\Omega(\varepsilon^{-1}\eta^{-2}n\log^{2}(n))-O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))=\omega(\varepsilon^{-1}\eta^{-1}\log^{2}(n))$. Thus, we also know that (by Algorithm 6.1 of Algorithm 6.1), for a constant $c>0$,
$1\geq p$ $=\frac{cn\log^{2}(n)}{\varepsilon\cdot\eta^{2}\cdot m_{t}^{\prime}}$ (56)
$\geq\frac{cn\log^{2}(n)}{\varepsilon\cdot\eta^{2}\cdot\left((1+\eta)\cdot m_{t}+O\left(\frac{\log^{2}n}{\varepsilon\eta}\right)\right)}$ (57)
$\geq\frac{cn\log^{2}(n)}{c^{\prime}\left(\varepsilon\cdot\eta^{2}\cdot m_{t}\right)}$ since $m_{t}=\omega(\varepsilon^{-1}\eta^{-1}\log^{2}n)$ (58)
for an appropriately large constant $c^{\prime}>0$.
[…]
$\mathbf{P}\left[|X[V^{\prime}]|\geq(1+\psi)\cdot\mu_{|X[V^{\prime}]|}\right]$ (59)
$\leq\exp\left(-\frac{p^{2}k^{2}\eta^{2}\mathrm{OPT}(G_{t})^{2}\cdot\mu_{|X[V^{\prime}]|}}{14\cdot 4\mu_{|X[V^{\prime}]|}^{2}}\right)$ (60)
$\leq\exp\left(-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right),$ (61)
where the last expression follows since $p\cdot k\cdot\mathrm{OPT}(G_{t})\geq\mu_{|X[V^{\prime}]|}$ due to the fact that by Equation 52, $p\cdot k\cdot\mathrm{OPT}(G_{t})\geq p\cdot k\cdot\operatorname{den}(G[V^{\prime}])=p\cdot(|G[V^{\prime}]|)=\mu_{|X[V^{\prime}]|}$. Then, suppose that $\psi>12$; substituting this $\psi$ into Theorem 3.1 gives the following probability expression: $\mathbf{P}[X\geq(1+\psi)\mu]\leq\exp\left(-\frac{\psi\mu}{2}\right)$ since $\frac{\psi}{2+\psi}\geq\frac{1}{2}$ for the given bound on $\psi$. Thus, using this value of $\psi$ in Equation 53 yields the following bound
$\mathbf{P}\left[|X[V^{\prime}]|\geq(1+\psi)\cdot\mu_{|X[V^{\prime}]|}\right]$ (62)
$\leq\exp\left(-\frac{pk\eta\mathrm{OPT}(G_{t})\cdot\mu_{|X[V^{\prime}]|}}{2\cdot 2\mu_{|X[V^{\prime}]|}}\right)$ (63)
$\leq\exp\left(-\frac{pk\eta\mathrm{OPT}(G_{t})}{4}\right),$ (64)
Altogether, for both cases, our expression is bounded by
$\mathbf{P}\left[|X[V^{\prime}]|\geq(1+\psi)\cdot\mu_{|X[V^{\prime}]|}\right]\leq\exp\left(-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right).$ (65)
We can simplify the LHS of the above inequality by:
$\mathbf{P}\left[|X[V^{\prime}]|\geq\mu_{|X[V^{\prime}]|}+\frac{pk\eta\mathrm{OPT}(G_{t})}{2}\right]$ $=\mathbf{P}\left[\frac{|X[V^{\prime}]|}{pk}\geq\frac{\mu_{|X[V^{\prime}]|}}{pk}+\frac{\eta\mathrm{OPT}(G_{t})}{2}\right]$ (66)
$=\mathbf{P}\left[\frac{1}{p}\cdot\operatorname{den}(X[V^{\prime}])\geq\frac{1}{p}\cdot p\cdot\operatorname{den}_{G_{t}}(V^{\prime})+\frac{\eta\mathrm{OPT}(G_{t})}{2}\right]$ (67)
$=\mathbf{P}\left[\frac{1}{p}\cdot\operatorname{den}(X[V^{\prime}])\geq\operatorname{den}_{G_{t}}(V^{\prime})+\frac{\eta\mathrm{OPT}(G_{t})}{2}\right].$ (68)
We now consider all subsets of vertices of size $k$. There are at most $\binom{n}{k}$ such sets. There are also $n$ possible values of $k$ since there are at most $n$ vertices. Hence, the total probability that the bound holds for all $V^{\prime}\subseteq V$ (using Equation 65) is
$\sum_{k=1}^{n}\binom{n}{k}\cdot\exp\left(-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right)$ $\leq\sum_{k=1}^{n}n^{k}\cdot\exp\left(-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right)$ (69)
$=\sum_{k=1}^{n}\exp\left(k\cdot\ln(n)-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right).$ (70)
Note that the densest subgraph has density at least the density of the entire graph, $\frac{m_{t}}{n}$. Combined with the substitution of Equation 58 into the above, we obtain
$\sum_{k=1}^{n}\exp\left(k\cdot\ln(n)-\frac{pk\eta^{2}\mathrm{OPT}(G_{t})}{56}\right)$ (71)
$\leq\sum_{k=1}^{n}\exp\left(k\cdot\ln(n)-\frac{cn\log^{2}(n)}{c^{\prime}\left(\varepsilon\cdot\eta^{2}\cdot m_{t}\right)}\cdot\frac{k\eta^{2}m_{t}}{56n}\right)$ by Equation 58 (72)
$\leq\sum_{k=1}^{n}\exp\left(k\cdot 2\log^{2}(n)-\frac{c\log^{2}(n)}{c^{\prime}\varepsilon}\cdot\frac{k}{56}\right)$ (73)
$=\sum_{k=1}^{n}\exp\left(-\left(\frac{c}{56c^{\prime}\varepsilon}-2\right)\cdot k\cdot\log^{2}(n)\right)$ (74)
$\leq\sum_{k=1}^{n}\exp\left(-\left(\frac{c}{56c^{\prime}\varepsilon}-2\right)\cdot\log^{2}(n)\right)$ since $k\geq 1$ (75)
$=n\cdot\exp\left(-\left(\frac{c}{56c^{\prime}\varepsilon}-2\right)\cdot\log^{2}(n)\right)$ (76)
$\leq\exp\left(-\left(\frac{c}{56c^{\prime}\varepsilon}-4\right)\cdot\log^{2}(n)\right)$ (77)
where Equation 75 holds when $\left(\frac{c}{56c^{\prime}\varepsilon}-2\right)\geq 1$. To obtain the high probability bound of $1/\operatorname*{poly}(n)$ over all time stamps $t$, it suffices to set $c$ to be a sufficiently large absolute constant and take a union bound over $T=\operatorname*{poly}(n)$ events. Hence, we have proven that with high probability, the density of any subgraph $G_{t}[V^{\prime}]$ is at least $\frac{1}{p}\operatorname{den}_{X}(V^{\prime})-\frac{\eta}{2}\operatorname{OPT}(G_{t})$ for every $t$. In particular, no induced subgraph in $X$ has estimated density greater than $(1+\frac{\eta}{2})p\cdot\mathrm{OPT}(G_{t})$, with high probability.
Now, we show that $\mathrm{OPT}(G_{t})$ has large enough induced density in $X$ such that the returned subgraph is a $(1+\eta)$-approximate densest subgraph in $G_{t}$. We do this by setting $V^{\prime}=V(\mathrm{OPT}(G_{t}))$, i.e. to the set of vertices of the densest subgraph in $G_{t}$. By applying a Chernoff bound (Theorem 3.1) with $\psi=\eta\in(0,1]$ and using the fact that $\operatorname{OPT}(G_{t})\geq\nicefrac{{m_{t}}}{{n}}$, we have that
$\mathbf{P}\left[\frac{1}{p}\operatorname{den}(X[V^{\prime}])\leq(1-\psi)\cdot\operatorname{OPT}(G_{t})\right]$ (78)
$=\mathbf{P}\left[\operatorname{den}(X[V^{\prime}])\leq(1-\psi)\cdot\mu\right]$ (79)
$\leq\exp\left(-\frac{\psi^{2}\mu}{2}\right)=\exp\left(-\frac{\psi^{2}\cdot p\cdot\operatorname{OPT}(G_{t})}{2}\right)$ (80)
$\leq\exp\left(-\frac{\psi^{2}}{2}\cdot\frac{cn\log^{2}(n)}{c^{\prime}\left(\varepsilon\cdot\eta^{2}\cdot m_{t}\right)}\cdot\frac{m_{t}}{n}\right)$ by Equation 58 (81)
$=\exp\left(-\frac{c\log^{2}(n)}{2c^{\prime}\varepsilon}\right)$ by setting $\text{by setting }\psi=\eta.$ (82)
By setting a large enough constant $c>0$ and together with what we showed above, we obtain that with high probability, the densest subgraph in $X$ has induced density in $G_{t}$ that is a $(1+\eta)$-approximation of $\mathrm{OPT}(G_{t})$. We take the union bound over $T=\operatorname*{poly}(n)$ to show that for each time step $t$, our bound holds.
Approximation.
By Theorem C.1, the vertex set $S$ we output is a $(2,O(\varepsilon_{1}^{-1}\log n))$-approximate densest subgraph of $X$ after each call to the private static densest subgraph subroutine. If we have yet to begin sub-sampling, i.e. $X=G_{t}$, then we are done.
Suppose the algorithm started to sub-sample. Then with high probability, after each call to the private static densest subgraph algorithm,
$\operatorname{den}_{G_{t}}(S)$ (83)
$\geq\frac{1}{p}\operatorname{den}_{X}(S)-\frac{\eta\operatorname{OPT}(G_{t})}{2}$ by Equation 68 (84)
$\geq\frac{1}{p}\left(\frac{\operatorname{OPT}(X)}{2}-O(\varepsilon_{1}^{-1}\log n)\right)-\frac{\eta\operatorname{OPT}(G_{t})}{2}$ by Theorem C.1 (85)
$\geq\frac{(1-\eta)\operatorname{OPT}(G_{t})}{2}-\frac{c^{\prime}\left(\varepsilon\cdot\eta^{2}\cdot m_{t}\right)}{cn\log^{2}(n)}\cdot\frac{c^{\prime\prime}\log^{2}(n)}{\varepsilon\eta}-\frac{\eta\operatorname{OPT}(G_{t})}{2}$ since $\text{since }\varepsilon_{1}=\Theta\left(\frac{\varepsilon}{\log_{1+\eta}(n)}\right)$ (86)
$=\frac{(1-2\eta)\operatorname{OPT}(G_{t})}{2}-\frac{c^{\prime}c^{\prime\prime}\eta m_{t}}{cn}$ (87)
$\geq\frac{(1-3\eta)\operatorname{OPT}(G_{t})}{2},$ for $\text{for }c\geq 2c^{\prime}c^{\prime\prime},\operatorname{OPT}(G_{t})\geq\frac{m_{t}}{n}$ (88)
where Equation 86 follows from Equation 78.
Finally, we account for the additional error due to not updating $S$ when the SVT (Algorithm 6.2) does not output “above”. Let $D_{t}$ denote the exact density of the subgraph and $p_{t}$ the sampling probability at time $t$ as returned on Algorithm 6.2. If the SVT did not output “above”, we know that with high probability,
$D_{t}$ $\leq p_{t}\cdot L+O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))$ (89)
$\frac{D_{t}}{p_{t}}$ $\leq L+\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}$ (90)
$D_{t}$ $>p_{t}\cdot\frac{L}{1+\eta}-O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))$ (91)
$\frac{D_{t}}{p_{t}}$ $>\frac{L}{1+\eta}-\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}.$ (92)
Similar to the calculation we made regarding $\operatorname{OPT}(X)$, either $p=1$ and there is nothing to prove or $p\geq\frac{cn\log^{2}(n)}{c^{\prime}\varepsilon\eta^{2}m_{t}}$ and
$\frac{O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))}{p_{t}}$ $=O(\varepsilon^{-1}\eta^{-1}\log^{2}(n))\cdot\frac{c^{\prime}\varepsilon\eta^{2}m_{t}}{cn\log^{2}(n)}$ (93)
$\leq\frac{\eta m_{t}}{n}$ (94)
$\leq\eta\operatorname{OPT}(G_{t}).$ (95)
In other words, assuming a sufficiently large constant $c$, we incur an additional $(1+2\eta)$ multiplicative error and $O(\varepsilon^{-1}\eta^{-1}\log^{2}n)$ additive error due to the SVT. Hence, we obtain a $(2+O(\eta),O(\varepsilon^{-1}\eta^{-1}\log^{2}n))$-approximate densest subgraph at every timestamp $t$. ∎ Combining Lemma 6.7 and Lemma 6.3 yields the proof of Theorem 6.1.
Reducing the Multiplicative Factor
We remark that the only reason we require $1/p=\Omega(\varepsilon^{-1}\eta^{-2}n\log^{2}n)$ and not $\Omega(\varepsilon^{-1}\eta^{-2}n\log n)$ is to absorb the additive error of the private static densest subgraph subroutine (Theorem C.1) into the multiplicative error (see Equation 86). If we instead use an alternative private static densest subgraph subroutine, say Theorem C.2, which yields a $O(1+\eta,\varepsilon_{1}^{-1}\log^{4}(n))$-approximation, we would need to set $1/p=\Omega(\varepsilon^{-1}\eta^{-2}n\log^{5}(n))$ (Algorithm 6.1) since $\varepsilon_{1}=\varepsilon/\log_{1+\eta}(n)$. Suppose we further adjust the initial edge threshold to $m^{\prime}=\Theta(\varepsilon^{-1}\eta^{-2}n\log^{5}(n))$ (Algorithm 6.1), the initial density threshold to $L=\Theta(\varepsilon^{-1}\eta^{-1}n\log^{5}(n))$, and repeat the proof steps of Lemma 6.7 for this version of our algorithm. This yields a proof of Theorem 6.2.
7 $k$-Core Decomposition
In this section, we introduce a semi-streaming sampling algorithm that preserves the $k$-cores in an input graph $G=(V,E)$ while ensuring privacy. Specifically, we prove the following theorem.
Theorem 7.1 (Sublinear Space Private $k$-Core Decomposition).
Fix $\eta\in(0,1]$. Algorithm 7.2 is an $\varepsilon$-DP algorithm for the $k$-core decomposition problem in the continual release model for insertion-only streams. At every $t\in[T]$, the algorithm returns a value for each vertex, such that every value is a $\left(2+\eta,\frac{\log^{3}(n)}{\eta^{2}\varepsilon}\right)$-approximation of the corresponding vertex’s true core value, with probability $1-1/\operatorname*{poly}(n)$. The maximum space used is $O\left(\varepsilon^{-1}n\log^{5}n\right)$, with probability $1-1/\operatorname*{poly}(n)$.
Our algorithm is similar in spirit to our densest subgraph algorithm although it differs in several crucial and non-trivial respects as outlined below, making our algorithm conceptually more complicated than our densest subgraph algorithm.
  • We cannot use a differentially private $k$-core decomposition algorithm as a black-box because all vertices may change their core numbers at some point in the stream, resulting in $\Omega(n)$ changes from the black-box private $k$-core decomposition algorithm. Such a black-box usage of the private algorithm is difficult to analyze without losing a factor of $\varepsilon\cdot n$ in the pure DP setting resulting from composition.
  • We cannot produce one uniform sample of edges from the graph since this results in an uneven distribution of edges among the cores. Suppose we sample each edge with probability $p$, if $p$ is set to be too small, then vertices with smaller core numbers will not have enough adjacent edges to produce a good concentration bound.
  • We must also be able to deal with vertices which have large degree and very small cores (consider a star graph). Sampling edges uniformly from the original graph, without maintaining some form of a data structure on the sampled edges, will not allow us to distinguish between vertices with large degree and large core numbers from vertices with large degree but small core numbers.
  • We require a new composition theorem that allows us to not lose privacy for each release of a new core number among the $n$ vertices. Intuitively, we should not lose privacy for each release because for edge-neighboring insertion-only streams, only one edge differs between the two streams. This theorem uses the recently introduced multidimensional sparse vector technique [DLL23].
We give the pseudocode for our data structure and algorithm in Algorithm 7.1 and Algorithm 7.2, respectively. Below, our algorithm is inspired by the insertion-only sketching algorithm of [Esfandiari2018] to maintain our distance preserving sparsified graph from which we find our approximate and private $k$-core values; however, it is simpler in nature and uses an entirely new analysis. We combine a form of their algorithm with a version of the insertion-only level data structure of [DLRSSY22] to obtain our final algorithm. However, since [Esfandiari2018] is not private and [DLRSSY22] cannot use sublinear (in the number of edges) space, our algorithm non-trivially combines ideas in both. Furthermore, since the degeneracy of the graph is equal to the maximum core number of any node in the input graph, our algorithm also gives a private approximation of the degeneracy of the input graph.
1 Class PrivateCore($\varepsilon,\eta,L,n,T$)
2       Maintain sampled edges $X_{j}\leftarrow\varnothing$ for each $j\in Q$ where $Q=\{\lceil\log_{(1+\eta)}(L)\rceil,\dots,\lceil 2\log_{(1+\eta)}(n)\rceil\}$
3       Initialize $F\leftarrow\lceil 2\log_{(1+\eta)}(n)\rceil$triangleright $F-1$ denotes the topmost level of any level data structure
4
5      Initialize $\varepsilon_{1}\leftarrow\varepsilon/(6|Q|F)$
6       for $j\in Q$ do
7             Initialize $p_{j}\leftarrow\frac{c_{1}\log^{3}(n)}{\varepsilon(1+\eta)^{j}}$
8             for $v\in V$ do
9                   Initialize class $\textsc{SVT}^{j,v}(\varepsilon_{1},1,c_{2}\log_{1+\eta}(n))$   (Algorithm 3.1)
10                   Initialize $levels[j][v]\leftarrow 0$
11
12
13      Function SampleEdge($e_{t}$)
14             for $j\in Q$ do
15                   for $w\in e_{t}=\{u,v\}$ do
16                         if $levels[j][w]<F-1$ then
17                               Sample $e_{t}$ into $X_{j}$ with probability $p_{j}$.
18
19
20
21
22
23      Function UpdateLevels()
24             for $j\in Q$ do
25                   for level $\ell\in\{0,\dots,F-2\}$ do
26                         for $v\in V$ do
27                               if $levels[j][v]\neq\ell$ then
28                                     go to next iteration
29
30                              if $\textsc{SVT}^{j,v}\textsc{.ProcessQuery}(\deg^{+}_{X_{j}}(v),p_{j}\cdot(1+\eta)^{j-1})$ is “above” then
31                                     $levels[j][v]\leftarrow levels[j][v]+1$
32
33
34
35
36
37      Function GetPrivateLevels()
38             return $levels$
39
40
Algorithm 7.1 Data Structure for $k$-Core Decomposition for Adaptive Insertion-Only Streams
41 Initialize initial core threshold $L\leftarrow\frac{c_{3}\log^{3}(n)}{\varepsilon}$
42 Initialize PrivateCore($\varepsilon,\eta,L$)
[…]
44 for each new update $e_{t}$ do
45       if $e_{t}\neq\bot$ then
46             PrivateCore.SampleEdge($e_{t}$)
47
48
[…]
51       for every vertex $v\in V$ do
52             $j_{now}\leftarrow\max\{j:levels[j][v]=F-1\}$
53             if $(1+\eta)^{j_{now}}>L$ then
54                   Release $v$’s core as $(2+\eta)\cdot(1+\eta)^{j_{now}}$
55
56            else
57                   Release $v$’s core as 1
58
59
60
Algorithm 7.2 Algorithm for $k$-Core Decomposition for Adaptive Insertion-Only Streams
7.1 Algorithm Description
Algorithm Intuition
First, we give some intuition for our algorithm. Our algorithm essentially performs a sampling version of the classic peeling algorithm for $k$-core decomposition. The classic peeling algorithm successively peels (removes) vertices with the minimum degree until all vertices are removed from the graph. A core that is formed during the peeling process is the induced subgraph consisting of the remaining vertices after a vertex is peeled and the value of such a core is the minimum induced degree within the subgraph. The core number for each vertex $v$ is equal to the maximum valued core that $v$ is a part of during any stage of the peeling. A dynamic version of this algorithm can be obtained by maintaining a level data structure where a vertex is moved up a level if its induced degree among vertices in the same or higher levels is larger than a cutoff $C$. One can show that having $O(\log n)$ levels of the structure and appropriately setting $C$ among $O(\log n)$ duplicates of the structure gives a $(2+\eta)$-approximation of the core numbers of the nodes in the non-private, insertion-only setting [SCS20, LSYDS22, DLRSSY22].
When sparsifying the graph, we cannot simply take a uniform sample of the edges adjacent to each vertex. An easy example to consider is a vertex $v$ which is part of a $10$-clique and also adjacent to $n/2$ degree one vertices. A uniform sample of the edges adjacent to $v$ will most likely not discover the $9$ edges connecting it to the $10$-clique (for large $n$). We call the edges connecting $v$ to the $10$-clique the set of important edges. Thus, we must take a smarter sample of edges adjacent to $v$. To maintain a sparsified, sampling-based version of the level data structure, we maintain samples of large enough size of the up-edges adjacent to each vertex. The up-edges adjacent to each vertex are the edges connecting each vertex to neighbors in the same or higher levels. Once we see enough sampled up-edges, we move the vertex up one level and continue sampling edges until we either reach the topmost level or the vertex is adjacent to only a very small sample of up-edges. Such a sampling method allows us to keep enough of the important edges which connect to other vertices in higher valued cores.
Finally, to make the above algorithm differentially private, we use SVT to determine when to move the vertex up a level. We show that although many vertices may move up levels, we only lose privacy when the vertices that are adjacent to the edge that differs between neighboring streams, move up. Since our total number of levels and duplicates is bounded by $O(\log^{2}n)$, the privacy loss from SVT is also bounded by $O(\log^{2}n)$.
Detailed Algorithm Description
We now give the detailed description of our algorithm. Our algorithm proceeds as follows. We maintain $O\left(\log_{(1+\eta)}(n)\right)$ subgraphs where for each subgraph, the probability we use to sample edges into the subgraph is different. Let $Q=[\lceil\log_{1+\eta}(L)\rceil,\dots,\lceil 2\log_{(1+\eta)}(n)\rceil]$. Specifically, we consider all subgraphs $j\in Q$ (Algorithm 7.1) where the integer $j\geq\log_{(1+\eta)}(L)$ and we set $L=\frac{c_{3}\log^{3}(n)}{\varepsilon}$ (Algorithm 7.2). For each update we receive in the stream (Algorithm 7.2), we first determine which subgraphs to sample the edge into using the procedure PrivateCore.SampleEdge($e_{t}$) (Algorithm 7.2). The procedure determines whether to sample an edge by iterating through all of the subgraphs $X_{j}$ for $j\in Q$ (Algorithm 7.1). We decide to sample $e_{t}=\{u,v\}$ into $X_{j}$ by looking at both endpoints of the edge update (Algorithm 7.1). If for either endpoint $w\in\{u,v\}$, vertex $w$ is not on level $F-1$ of $X_{j}$ (Algorithm 7.1), then we sample the edge using probability $p_{j}$ (Algorithm 7.1).
Each $X_{j}$ is organized into levels where vertices are moved up levels if they have induced degree among vertices at the same or higher level approximately greater than $(p_{j})\cdot(1+\eta)^{j-1}$. We require the degree to be approximately greater rather than exactly greater to preserve privacy.
[…]
Let $\deg^{+}_{X_{j}}(v)$ be the degree of $v$ in the induced subgraph of its neighbors at the same level or higher. To see whether we should move that vertex up a level, we first determine whether the current level of that vertex is the same level that we are iterating (Algorithm 7.1). If it is the same level and we pass the SVT check that $\deg^{+}_{X_{j}}(v)$ exceeds $p_{j}\cdot(1+\eta)^{j}$ (Algorithm 7.1), then we move the vertex up a level by incrementing $levels[j][v]$ (Algorithm 7.1). The particular threshold of $p_{j}\cdot(1+\eta)^{j}$ that we use will become apparent once we discuss our analysis of the approximation factor. Intuitively, the levels mimic the traditional peeling algorithm for the static $k$-core decomposition problem. The threshold $(1+\eta)^{j}$ is used in previous works (e.g. [DLRSSY22]) and because we are sampling edges instead of using the entire graph, the threshold is multiplied by $p_{j}$.
Finally, in the last part of the algorithm, we iterate through each vertex (Algorithm 7.2) and release an estimate if the vertex $v$ is in the topmost level of a subgraph $X_{j_{now}}$ where $j_{now}$ is maximized. If this is the case, we release the new estimate $(2+\eta)\cdot(1+\eta)^{j_{now}}$ for vertex $v$ (Algorithm 7.2), assuming $(1+\eta)^{j_{now}}$ exceeds some data-oblivious lower bound $L$.
7.2 Privacy Guarantee
We now prove the privacy guarantees of our algorithm. Specifically, we show that our algorithm maintains $\varepsilon$-differential privacy on edge-neighboring insertion-only streams. Note that although we are using intuition from the multidimensional sparse vector technique given in [DLL23], our proof is different in that each vertex can satisfy the SVT query multiple times, adaptively, as we receive more edges in the stream. This is a subtle difference since the original mechanism only works in the static setting. Hence, we provide the full proof below.
Lemma 7.2.
Algorithm 7.2 is $\varepsilon$-differentially private on edge-neighboring insertion-only streams as defined in Definition 3.3.
Proof.
We first observe that the only information that depends on private data is the level of each node in each $X_{j}$ for $j\in Q$. The level of any vertex $v$ depends on the edges that are sampled from the stream for each graph as well as their $\deg^{+}_{X_{j}}(v)$ values. This proof is a variant of the multidimensional AboveThreshold (MAT) technique used in [DLL23] but we state a version of the proof that directly proves the privacy of the algorithm in our setting; we follow the proof style of [lyu2017understanding] below. Let $e_{t^{*}}=\{x,y\}$ be the fixed edge that differs between the two neighboring streams $G$ and $G^{\prime}$.
SVT is called in Algorithm 7.1 of Algorithm 7.1 to determine whether a vertex moves up a level. In fact, vertices can change levels if and only if this SVT passes (and when the current level of the vertex is $\ell$). We make the observation that for each vertex $v$ which is not incident to $e_{t^{*}}$, the distributions of the outputs of the SVT queries are the same in $G_{t}$ and $G^{\prime}_{t}$ when conditioned on the levels of the vertices in each $X_{j}$. Let $E_{1}$ be the event where prior to the SVT call in Algorithm 7.1 at time $t$ all vertices are in the levels given in $E_{1}$. Furthermore, $E_{1}$ has a sample of edges, $e_{t}$’s, which may or may not contain $e_{t^{*}}$. We now show the probability of the next $A$ failed queries to Algorithm 7.1 by vertex $w$ in graph $X_{j}$; a failed query is one that does not return “above”. We denote a failed SVT query (returns “below”) by $fail$ and a successful SVT query (returns “above”) by $success$. We denote the output of Algorithm 7.1 by $\mathcal{S}_{j,v}(\mathcal{X})$ where $\mathcal{X}$ is the state of our subgraphs in $G$ and $\mathcal{X}^{\prime}$ is the state of our subgraphs in $G^{\prime}$. First, for every $w\not\in\{x,y\}$, it holds that $\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X})=\{fail\}^{A}\mid E_{1}]=\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=\{fail\}^{A}\mid E_{1}]$ since $\deg_{X_{j}}^{+}(w)=\deg_{X^{\prime}_{j}}^{+}(w)$ when conditioned on $E_{1}$. Then, we show that for every $w\in\{x,y\}$,
$\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X})=\{fail\}^{A}\mid E_{1}]$ $\leq e^{\varepsilon_{1}}\cdot\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=\{fail\}^{A}\mid E_{1}].$ (96)
We show Equation 96 as follows where $f$ is the probability density function for picking a threshold noise and $g_{\mathcal{\mathcal{X}}}$ is the probability density function for failing Algorithm 7.2 using $\mathcal{X}$. Let $\nu^{i}_{j,w}$ be the individual noises that are picked each time a query is made to Algorithm 7.2. For simplicity of expression, we do not write the conditioning on $E_{1}$ on the RHS but all expressions below are conditioned on $E_{1}$.
[…]
$=\int_{-\infty}^{\infty}f(z)\cdot\prod_{i\in[A]}\mathbf{P}\left[\deg^{+}_{i,X_{j}}(w)+\nu^{i}_{j,w}<\frac{c_{1}\log^{3}(n)}{\varepsilon}+z\right]dz$ (98)
$=\int_{-\infty}^{\infty}f(z)\cdot\prod_{i\in[A]}\mathbf{P}\left[\nu^{i}_{w,j}<\frac{c_{1}\log^{3}(n)}{\varepsilon}-\deg^{+}_{i,X_{j}}(w)+z\right]dz$ (99)
$\leq\int_{-\infty}^{\infty}f(z)\cdot\prod_{i\in[A]}\mathbf{P}\left[\nu^{i}_{w,j}<\frac{c_{1}\log^{3}(n)}{\varepsilon}-\deg^{+}_{i,X^{\prime}_{j}}(w)+z+1\right]dz$ (100)
[…]
$=e^{\varepsilon_{1}}\cdot\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=\{fail\}^{A}\mid E_{1}].$ (104)
We now prove the privacy characteristics of the cases where the queries succeed. As in the case above, for all $w\not\in\{x,y\}$, it holds that $\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X})=success\mid E_{2}]=\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=success\mid E_{2}]$ conditioned on $E_{2}$, the event where we fix the sampled edges and levels of vertices. Now, we consider the case where $w\in\{x,y\}$. As before, we show that
$\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X})=success\mid E_{2}]$ $\leq e^{\varepsilon_{1}}\cdot\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=success\mid E_{2}].$ (105)
Using $f$ as the probability density function for the noise picked for the threshold, $h_{\mathcal{\mathcal{X}}}$ is the probability density function for success, and conditioning on $E_{2}$ (for simplicity of expression, we do not write the conditioning on the RHS):
[…]
$=e^{2\varepsilon_{1}}\cdot\mathbf{P}[\mathcal{S}_{j,w}(\mathcal{X}^{\prime})=success\mid E_{2}].$ (113)
Using the above proofs, the threshold in Algorithm 7.1 is exceeded at most $|Q|\cdot F$ total times and so Algorithm 7.1 is satisfied at most $|Q|\cdot F$ times. Because there is a deterministic mapping between the SVT answers and the levels, i.e., the levels only change when the SVT accepts, it is sufficient to show that the levels can be published differentially privately as the rest of the computation in Algorithm 7.2 can be achieved through postprocessing.
Below, our algorithm is denoted $\mathcal{M}$ and $\mathcal{M}$ outputs a set of levels, one for each vertex, subgraph, and timestamp, $\ell_{j,v}^{t}$ for each $v\in V,t\in\{1,\dots,T\}$, and $j\in\{1,\dots,\lceil 2\log_{(1+\eta)}(n)\rceil\}$, which directly determines the approximate core number for each vertex. Each $\ell_{j,v}^{t}$ is determine solely by the success and failure of the threshold queries. Let $\mathcal{X}_{t}$ be the state of our subgraphs at time $t$ (where $\mathcal{X}_{0}$ is the initial state where all vertices are on level 00 for each subgraph) and we fix the set of outputs of $\mathcal{S}_{j,w}(\mathcal{X}_{t})$ by $\textbf{b}^{t}=\left[b_{1,1}^{t},\dots,b_{\lceil 2\log_{(1+\eta)}(n)\rceil,n}^{t}\right]$. We use $\mathcal{S}(\mathcal{X}_{t})=\textbf{b}^{t}$ as shorthand for $\mathcal{S}_{j,w}(\mathcal{X}_{t})=b^{t}_{j,w}$ for all $j,w$. Then, since $\varepsilon_{1}=\varepsilon/(6|Q|F)$, we can use the chain rule and the above expressions to show the following:
$\mathbf{P}[\mathcal{M}(G)=(\ell_{1,v_{1}}^{1},\dots,\ell_{j,v}^{t},\dots,\ell_{\lceil 2\log_{(1+\eta)}(n)\rceil,v_{n}}^{T})]$ (114)
$\leq\prod_{t\in\{1,\dots,T\}}\mathbf{P}\left[\bigcap_{j,w}\mathcal{S}_{j,w}(\mathcal{X}_{t})=b_{j,w}^{t}\mid\mathcal{S}(\mathcal{X}_{t-1})=\textbf{b}^{t-1}\cap\cdots\cap\mathcal{S}(\mathcal{X}_{0})=\textbf{b}^{0}\right]$ (115)
$\leq\prod_{t\in\{1,\dots,T\}}\prod_{j,w}\mathbf{P}\left[\mathcal{S}_{j,w}(\mathcal{X}_{t})=b_{j,w}^{t}\mid\mathcal{S}(\mathcal{X}_{t-1})=\textbf{b}^{t-1}\cap\cdots\cap\mathcal{S}(\mathcal{X}_{0})=\textbf{b}^{0}\right]$ (116)
$\leq\exp(\varepsilon_{1})^{2|Q|F}\cdot\exp(2\varepsilon_{1})^{2|Q|F}\cdot\prod_{t\in\{1,\dots,T\}}\prod_{j,w}\mathbf{P}\left[\mathcal{S}_{j,w}(\mathcal{X}^{\prime}_{t})=b_{j,w}^{t}\mid\mathcal{S}(\mathcal{X}_{t-1}^{\prime})=\textbf{b}^{t-1}\cap\cdots\cap\mathcal{S}(\mathcal{X}_{0}^{\prime})=\textbf{b}^{0}\right]$ (117)
$\leq\left(\exp(\varepsilon_{1})^{|Q|\cdot F}\right)^{2}\cdot\left(\exp(\varepsilon_{1})^{2|Q|\cdot F}\right)^{2}\cdot\mathbf{P}[\mathcal{M}(G^{\prime})=(\ell_{1,v_{1}}^{1},\dots,\ell_{j,v}^{t},\dots,\ell_{\lceil 2\log_{(1+\eta)}(n)\rceil,v_{n}}^{T})]$ (118)
$\leq\exp(\varepsilon)\cdot\mathbf{P}[\mathcal{M}(G^{\prime})=(\ell_{1,v_{1}}^{1},\dots,\ell_{j,v}^{t},\dots,\ell_{\lceil 2\log_{(1+\eta)}(n)\rceil,v_{n}}^{T})].$ (119)
Equation 115 follows from the chain rule and since the levels are determined by the success and failures of the SVT threshold queries. Equation 116 follows since each output of the threshold query at $t$ is independent conditioned on the states of the vertices from $t-1$. Equation 117 follows from Equation 104 and Equation 113 for vertices $x$ and $y$ and the fact that there are at most $|Q|\cdot F$ successes for each of $x$ and $y$ (where $e_{t^{*}}=\{x,y\}$) and hence also $|Q|\cdot F$ consecutive runs of failures. Equation 118 follows since the levels of the vertices in $G^{\prime}$ are determined by the successes and failures of the queries on $G^{\prime}$. Finally, Equation 119 follows because we set $\varepsilon_{1}=\varepsilon/(6|Q|F)$. This concludes the proof of the privacy of our algorithm since only the levels are used (via post-processing) to obtain the approximations. ∎
7.3 Approximation Guarantees
Given our privacy guarantees, we now prove our approximation guarantees in this section. We first prove our concentration bound on the samples we obtain in our procedure. Specifically, we show that with high probability, the sampled edges give an approximately accurate estimate of $\deg^{+}_{X_{j}}(v)$. Our proof strategy is as follows. We use the proof of the approximation given in [DLRSSY22, Theorem 4.7]. In order to use their theorem, we consider a hypothetical set of level data structures where we keep all of the edges in the graph. Note that we do not maintain these level data structures in our algorithm but only use them for the sake of analysis. Let this set of level data structures be denoted as $\mathcal{K}$. We place each vertex $v$ in $\mathcal{K}$ on the exact same level as $v$ in the level data structures within Algorithm 7.1. Then, we show that the vertices in $\mathcal{K}$ satisfy modified versions of Invariant 3 (Lemma 7.3) and Invariant 4 (Lemma 7.4) in [DLRSSY22], which directly gives our approximation factor by a modified version of [DLRSSY22, Theorem 4.7] which we present in Appendix G.
In the below proofs, let $\deg_{\mathcal{K}}(\ell,j,v)$ be the induced degree of $v$ in the $j$-th level data structure within $\mathcal{K}$ consisting of all neighbors of $v$ that are on level $\ell$ and higher. We prove the following lemmas for graph $G_{t}$ and $G^{\prime}_{t}$ which are formed after the first $t\in[T]$ updates.
Lemma 7.3.
If vertex $v$ is in level $\ell<F-1$, with high probability, in subgraph $X_{j}$ after the levels are updated by Algorithm 7.2, then, $\deg_{\mathcal{K}}(\ell,j,v)\leq(1+\eta)^{j}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$, with high probability.
Proof.
We prove this statement via contradiction. Suppose $v$ is in level $\ell<F-1$ within subgraph $X_{j}$ with high probability and $\deg_{\mathcal{K}}(\ell,j,v)>(1+\eta)^{j}+\frac{a_{1}\log^{3}n}{\varepsilon}$ with probability at least $n^{-a_{2}}$ for some fixed constants $a_{1},a_{2}\geq 1$. We are in graph $X_{j}$, hence the probability we used to sample is $p_{j}=\frac{c_{1}\log^{3}(n)}{\varepsilon(1+\eta)^{j}}$. The expected number of edges we sample out of the $\deg_{\mathcal{K}}(\ell,j,v)$ edges using $p_{j}$ is at least
$\mu\geq\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{j}}\cdot\left((1+\eta)^{j}+\frac{a_{1}\log^{3}n}{\varepsilon}\right)=\frac{c_{1}\log^{3}n}{\varepsilon}+\frac{a_{1}c_{1}\log^{6}n}{\varepsilon^{2}(1+\eta)^{j}}>\frac{c_{1}\log^{3}n}{\varepsilon}.$ (120)
Let $S_{\ell,j,v}$ be the set of edges we sampled. Using a multiplicative Chernoff bound (Theorem 3.1), we have that the probability that our sample has size smaller than $\frac{(1-\psi)c_{1}\log^{3}n}{\varepsilon}$ is as follows, for $\psi\in(0,1)$:
$\mathbf{P}\left[|S_{\ell,j,v}|\leq(1-\psi)\mu\right]\leq\exp\left(-\frac{\mu\psi^{2}}{3}\right)\leq\exp\left(-\frac{\psi^{2}c_{1}\log^{3}n}{3\varepsilon}\right).$ (121)
The SVT introduces at most $\frac{a_{3}\log^{3}(n)}{\varepsilon}$ additive error for some constant $a_{3}$, with high probability, and so the value we are comparing against the threshold of $p_{j}\cdot(1+\eta)^{j-1}$ is at least $|S_{\ell,j,v}|-\frac{2a_{3}\log^{3}(n)}{\varepsilon}$, with high probability. We proved above that the probability that $|S_{\ell,j,v}|>\frac{(1-\psi)c_{1}\log^{3}(n)}{\varepsilon}$ is at least $1-\exp\left(-\frac{\psi^{2}c_{1}\log^{3}(n)}{3\varepsilon}\right)$. Thus conditioned on the SVT error bound and sampling bound above, the SVT comparison succeeds if
$|S_{\ell,j,v}|-\frac{2a_{3}\log^{3}(n)}{\varepsilon}$ $\geq\frac{(1-\psi)c_{1}\log^{3}(n)}{\varepsilon}-\frac{2a_{3}\log^{3}(n)}{\varepsilon}$ (122)
$\geq p_{j}\cdot(1+\eta)^{j-1}$ (123)
$=\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)}.$ (124)
Note that this inequality always holds when
$(1-\psi)c_{1}-2a_{3}\geq\frac{c_{1}}{1+\eta}.$ (125)
To satisfy the above expression, we require $\psi<\frac{\eta}{\eta+1}$ and $\eta>0$, which is easily satisfied by the constraints of our problem, and also $c_{1}\geq-\frac{(\eta+1)2a_{3}}{\eta(\psi-1)+\psi}$. We set $c_{1}$ to be an appropriately large enough constant in terms of $\eta,a_{3},\psi$ to amplify the probability of success to satisfy with high probability.
For an appropriate setting of $c_{1},\psi,a_{3}$, we have that the SVT succeeds with probability at least $1-n^{-c}$ for any constant $c\geq 1$. Taking the union bound over all levels $\ell^{\prime}\leq\ell$ (such that SVT outputs succeed for all such $\ell^{\prime}$), this contradicts with the fact that $v$ is at level $\ell$ with high probability. ∎
Lemma 7.3 directly shows that Invariant 3 is satisfied in [DLRSSY22]. Now, we prove that Invariant $4$ is also satisfied. The proof follows a similar structure to the proof of Lemma 7.3.
Lemma 7.4.
If vertex $v$ is in level $\ell>0$, with high probability, in subgraph $X_{j}$ after the levels are updated by Algorithm 7.2, then, $\deg_{\mathcal{K}}(\ell-1,j,v)\geq(1+\eta)^{j-2}-O\left(\frac{\log^{3}n}{\varepsilon}\right)$, with high probability.
Proof.
As in the proof of Lemma 7.3, we prove this lemma by contradiction. Suppose $v$ is in level $\ell>0$, with high probability, in subgraph $X_{j}$ and $\deg_{\mathcal{K}}(\ell,j,v)<(1+\eta)^{j-2}-\frac{a_{1}\log^{3}n}{\varepsilon}$ with probability at least $n^{-a_{2}}$ for some fixed constants $a_{1},a_{2}\geq 1$. Suppose that $\deg_{\mathcal{K}}(\ell-1,j,v)=(1+\eta)^{j-1}-\frac{a_{1}\log^{3}n}{\varepsilon}-1$ since this is the worst case. Here, worst case means the probability that the SVT is satisfied and hence the vertex moves up a level is maximized. We are in graph $X_{j}$, hence the probability we used to sample is $\frac{c_{1}\log^{3}(n)}{\varepsilon(1+\eta)^{j}}$. The expected number of edges we sample out of the $\deg_{\mathcal{K}}(\ell,j,v)$ edges using $p_{j}$ is at least
$\mu\geq\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{j}}\cdot\left((1+\eta)^{j-2}-\frac{a_{1}\log^{3}n}{\varepsilon}-1\right)$ $=\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{2}}-\frac{a_{1}c_{1}\log^{6}n}{\varepsilon^{2}(1+\eta)^{j}}-p_{j}$ (126)
$\geq\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{3}}-\frac{2a_{1}c_{1}\log^{6}n}{\varepsilon^{2}(1+\eta)^{j}}.$ (127)
By Algorithm 7.2, we set $j$ such that $(1+\eta)^{j}\geq\frac{c_{3}\log^{3}(n)}{\varepsilon}$. Hence, using this setting, we can further simplify Equation 127 as follows:
$\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{2}}-\frac{2a_{1}c_{1}\log^{6}n}{\varepsilon^{2}(1+\eta)^{j}}\geq\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{2}}-\frac{2a_{1}c_{1}\log^{3}(n)}{c_{3}\varepsilon}=\left(\frac{1}{(1+\eta)^{2}}-\frac{2a_{1}}{c_{3}}\right)\frac{c_{1}\log^{3}(n)}{\varepsilon}.$ (128)
Furthermore, we can upper bound $\mu$ by
$\mu\leq\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{2}}.$ (129)
Let $S_{\ell,j,v}$ be the set of edges we sampled. Using a multiplicative Chernoff bound (Theorem 3.1), we have that the probability that our sample has size larger than $(1+\psi)\mu$ is as follows, for $\psi\in(0,1)$:
$\mathbf{P}\left[|S_{\ell,j,v}|\geq(1+\psi)\mu\right]\leq\exp\left(-\frac{\mu\psi^{2}}{3}\right)\leq\exp\left(-\frac{\psi^{2}\left(\frac{1}{(1+\eta)}-\frac{2a_{1}}{c_{3}}\right)c_{1}\log^{3}n}{3\varepsilon}\right).$ (130)
The SVT introduces at most $\frac{a_{3}\log^{3}(n)}{\varepsilon}$ additive error for some constant $a_{3}$, with high probability, and so the value we are comparing against the threshold of $p_{j}\cdot(1+\eta)^{j-1}$ is at most $|S_{\ell,j,v}|+\frac{2a_{3}\log^{3}(n)}{\varepsilon}$, with high probability. We proved above the probability that $|S_{\ell,j,v}|\leq\frac{(1+\psi)c_{1}\log^{3}(n)}{\varepsilon(1+\eta)}$ is at least $1-\exp\left(-\frac{\psi^{2}\left(\frac{1}{(1+\eta)}-\frac{2a_{1}}{c_{3}}\right)c_{1}\log^{3}(n)}{3\varepsilon}\right)$. Conditioning on the SVT error bound and the sampling bound above, the SVT fails the check at level $\ell-1$ if
$|S_{\ell,j,v}|+\frac{2a_{3}\log^{3}(n)}{\varepsilon}$ $\leq\frac{(1+\psi)c_{1}\log^{3}(n)}{\varepsilon(1+\eta)^{3}}+\frac{2a_{3}\log^{3}(n)}{\varepsilon}$ (131)
$<p_{j}\cdot(1+\eta)^{j-1}$ (132)
$=\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)}.$ (133)
Note that the inequality is satisfied as long as we have
$\frac{(1+\psi)c_{1}}{(1+\eta)^{2}}+2a_{3}<\frac{c_{1}}{1+\eta}.$ (134)
To satisfy the above expression, we require $\psi<\eta$ and $\eta>0$, which is easily satisfied by the constraints of our problem, and also $c_{1}>-\frac{(\eta+1)^{2}2a_{3}}{\psi-\eta}$. We set $c_{1}$ and $c_{3}$ to be appropriately large enough constants in terms of $\eta,a_{1},a_{3},\psi$ to amplify the probability of success to satisfy with high probability.
For an appropriate setting of $c_{1},c_{3},\psi,a_{1},a_{3}$, we have that the SVT failed the check at level $\ell-1$ with probability at least $1-n^{-c}$ for any constant $c\geq 1$. This contradicts the fact that $v$ is at level $\ell$, with high probability. ∎
Lemma 7.3 and Lemma 7.4 together with the proof of Theorem 4.7 of [DLRSSY22] gives our final accuracy and space guarantees for Theorem 7.1 below.
Lemma 7.5.
Algorithm 7.2 returns a $\left(2+\eta,O\left(\frac{\log^{3}n}{\varepsilon}\right)\right)$-approximate $k$-core decomposition, with high probability. The algorithm uses $O\left(\frac{n\log^{5}n}{\varepsilon}\right)$ space, with high probability.
Proof.
Lemma 7.3 and Lemma 7.4 shows that a vertex that is on level $\ell$ satisfies Invariant 3 and Invariant 4 of [DLRSSY22] in $\mathcal{K}$ and hence the approximation bound given by [DLRSSY22, Theorem 4.7] holds. The only change is that for Invariant $4$, the exponent for $(1+\eta)$ is $j-2$ instead of $j$. Carrying this exponent through the proof of [DLRSSY22, Theorem 4.7] results in an additional multiplicative factor of $(1+\eta)^{2}$ in the approximation. Because this proof is nearly verbatim as the proof in the original paper, we relegate the slightly modified proof that uses Lemma 7.3 and Lemma 7.4 to Appendix G. Thus, we give a $(2+\eta)(1+\eta)^{3}=(2+O(\eta))$-approximation, with high probability. Finally, by the Chernoff bound and since we set the probability of sampling to $\frac{c_{1}\log^{3}n}{\varepsilon(1+\eta)^{j}}$, we sample $O\left(\frac{\log^{3}n}{\varepsilon}\right)$ edges per level, per node, for each of $F$ levels in each of $|Q|$ graphs. Hence, in total over $F\cdot|Q|$ levels, we use $O\left(\frac{\log^{5}n}{\varepsilon}\right)$ space per node, with high probability. ∎
Finally, combining Lemma 7.2 and Lemma 7.5 yields the proof of Theorem 7.1.
8 Lower Bounds for Fully Dynamic Streams
In this section, we establish lower bounds on the additive error of differentially private algorithms for estimating the size of a maximum matching and the number of connected components in the continual release model. Similar to the lower bound for counting distinct elements in the continual release model [jain2023counting], we reduce the problem of answering matching queries and connected component queries to answering inner product queries. Then, we leverage known lower bounds for the inner product problem [Dinur2003Revealing, Dwork2007Price, MMNW11, De2012Lower] to obtain our lower bounds.
In Section 8.1, we review the key problem which we reduce to matching and connected components. The lower bound for maximum matching is proven in Section 8.2 and similarly for connected components in Section 8.3.
8.1 Inner Product Queries
Given a private database $y\in\{0,1\}^{n}$ and a set of $k$ linear queries $q^{(j)}\in\{0,1\}^{n},j\in[d]$, the inner product problem asks for the values
$\langle y,q^{(j)}\rangle\in\mathbb{Z}_{\geq 0},\qquad\forall j\in[k].$
We use the following lower bound stated in [jain2023counting] that is based on [Dinur2003Revealing, Dwork2007Price, MMNW11, De2012Lower].
Theorem 8.1 (Theorem 4.5 in [jain2023counting]).
There are constants $c_{1},c_{2}>0$ such that, for sufficiently large $n>0$: if an algorithm $\mathcal{A}$ answers $c_{1}n$ inner product queries within additive error $c_{2}\sqrt{n}$ with probability at least $0.99$, then $\mathcal{A}$ is not $(1,\nicefrac{{1}}{{3}})$-DP.
8.2 Maximum Cardinality Matching
We reduce to the inner product query problem similar to [jain2023counting]. Let $n$ be the dimension of a secret dataset $y\in\{0,1\}^{n}$. Consider $3n$ vertices $V=\{u_{i},v_{i},w_{i}:i\in[n]\}$. We use the first $\Theta(n)$ updates to encode the non-zero bits of $y$ as a matching between the vertices $u_{i},v_{i}$, i.e. we add the edge $\{u_{i},v_{i}\}$ if and only if $y_{i}=1$. Note that the size of the maximum matching is precisely $\lVert y\rVert_{0}$ by construction.
Given linear queries $q^{(1)}\in\{0,1\}^{n}$, we make the following updates: For each $i$ such that $q_{i}^{(1)}=1$, add the edge $\{v_{i},w_{i}\}$. Then the size of the maximum matching is $\lVert y\mid q^{(1)}\rVert_{0}$, where $a\mid b$ denotes the bitwise OR of $a,b\in\{0,1\}^{n}$. The inner product can then be recovered using the inclusion-exclusion principle:
$\langle y,q^{(1)}\rangle=\lVert y\rVert_{0}+\lVert q^{(1)}\rVert_{0}-\lVert y\mid q^{(1)}\rVert_{0}.$
We can then delete the added edges and process the next query $q^{(2)}$ and so on.
Since emulating each query requires $\Theta(n)$ updates, we can answer $\Theta(n)$ inner product queries after $T=\Theta(n^{2})$ updates. At a high level, since the error lower bound for $O(n)$ inner product queries is $\Omega(\sqrt{n})$, we have a lower bound of $\Omega(\min(n^{1/2},T^{1/4}))$.
1 Function Alg(private dataset $y\in\{0,1\}^{n}$, public queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$, DP mechanism for matchings $\mathcal{M}$)
2       $V\leftarrow\{u_{i},v_{i},w_{i}:i\in[n]\}$ ($3n$ vertices)
3       $E_{0}\leftarrow\varnothing$ empty edge set
4       $S^{(0)}\leftarrow$ empty update stream of length $n$
5       $S^{(1)},\dots,S^{(k)}\leftarrow$ $k$ empty update streams each of length $2n$
6       for $i=1,2,\dots,n$ do
7             if $y_{i}=1$ then
8                   $S_{i}^{(0)}\leftarrow+\{u_{i},v_{i}\}$ (insert edge $\{u_{i},v_{i}\}$)
9
10
11      for $j=1,2,\dots,k$ do
12             for $i=1,2,\dots,n$ do
13                   if $q^{(j)}_{i}=1$ then
14                         $S_{i}^{(j)}\leftarrow+\{v_{i},w_{i}\}$
15                         $S_{n+i}^{(j)}\leftarrow-\{v_{i},w_{i}\}$ (delete edge $\{v_{i},w_{i}\}$)
16
17
18
19      $S\leftarrow S^{(0)}+S^{(1)}+\dots+S^{(k)}$ concatenation of all streams
20       $r^{(0)},r^{(1)},\dots,r^{(k)}\leftarrow\mathcal{A}(S)$ answers to queries
21       for $j=1,2,\dots,k$ do
22             Output $\lVert q^{(j)}\rVert_{0}+r_{n}^{(0)}-r_{n}^{(j)}$ as approximate answer to $\langle q^{(j)},y\rangle$
23             $\triangleright$ $\langle q^{(j)},y\rangle=\lVert q\rVert_{0}+\lVert y\rVert_{0}-\lVert q^{(j)}\mid y\rVert_{0}$
24
25
Algorithm 8.1 Reduction from Inner Product Queries to Maximum Matching
Lemma 8.2.
Given an $\varepsilon$-DP mechanism for maximum matching in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq n+2nk$, Algorithm 8.1 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma 8.2).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm 8.1 and
$(r^{(0)},r^{(1)},\dots,r^{(k)})=\mathcal{M}(S^{(0)}+S^{(1)}+\dots+S^{(k)})$
be the corresponding stream of outputs when we run $\mathcal{M}$ on the concatenation of partial streams. By assumption, $r^{{(0)}}_{n}$ is an estimate of $\lVert y\rVert_{0}$ within additive error $\zeta$ and $r^{(j)}_{n}$ is an estimate of $\lVert y\mid q^{(j)}\rVert_{0}$ within additive error $\zeta$. Hence the value of
$\lVert q^{(j)}\rVert_{0}+r_{n}^{(0)}-r_{n}^{(j)}$
is an estimate of $\langle y,q^{(j)}\rangle$ with additive error within $2\zeta$. ∎
Theorem 8.3.
If $\mathcal{A}$ is a $1$-DP mechanism that answers maximum matching queries on graphs with $n$ vertices in the continual release model within additive error $\zeta$ with probability at least $0.99$ for fully dynamic streams of length $T$, then
$\zeta=\Omega(\min(\sqrt{n},T^{1/4})).$
Moreover, we may assume the graph is bipartite, has maximum degree 2, and arboricity 1.
Proof (Theorem 8.3).
Let $c_{1},c_{2}$ be the constants from Theorem 8.1. Fix $n\in\mathbb{N}$ and suppose $T\geq n+2c_{1}n^{2}$. By Lemma 8.2, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem 8.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
Now consider the regime of $T<n+2c_{1}n^{2}$. The same argument holds for $\tilde{n}=\Theta(\sqrt{T})$ such that $\tilde{n}+2c_{1}\tilde{n}^{2}\leq T$ to yield a lower bound of $\zeta\geq c_{2}\sqrt{\tilde{n}}=\Omega(T^{1/4})$. ∎
Lemma 8.4.
Let $\varepsilon\in(0,1)$, $\ell:=\lfloor\nicefrac{{1}}{{\varepsilon}}\rfloor$, $\tilde{T}\geq\ell$, $\zeta:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be an increasing error function. Suppose there is an $\varepsilon$-DP mechanism $\mathcal{A}$ for answering maximum matching queries on graphs of $\tilde{n}$ vertices in the continual release model that outputs estimates within additive error at most
$\ell\cdot\zeta(\nicefrac{{\tilde{n}}}{{\ell}},\nicefrac{{\tilde{T}}}{{\ell}})=O\left(\frac{\zeta(\varepsilon\tilde{n},\varepsilon\tilde{T})}{\varepsilon}\right)$
with probability $0.99$ for fully-dynamic streams of length $\tilde{T}$.
Let $T=\nicefrac{{\tilde{T}}}{{\ell}}$, and $n=\nicefrac{{\tilde{n}}}{{\ell}}$. There is a $1$-DP mechanism for answering maximum matching queries on graphs of $n$ vertices in the continual release model that outputs estimates within additive error at most $\zeta(n,T)$ with probability $0.99$ for fully-dynamic streams of length $T$.
Proof (Lemma 8.4).
Let $S$ be an edge update stream of length $T$ from a graph on $n$ vertices.
We construct an edge update stream of length $\tilde{T}=\ell T$ from a graph on $\tilde{n}=\ell n$ vertices as follows: Duplicate each vertex $v$ in the graph $\ell$ times into $v_{1},\dots,v_{\ell}$. For each update $S_{i}\in\{+\{u,v\},-\{u,v\},\perp\}$ from $S$, define the length $\ell$ partial stream
$S^{(i)}:=\begin{cases}+\{u_{1},v_{1}\},+\{u_{2},v_{2}\},\dots,+\{u_{\ell},v_{\ell}\},&S_{i}=+\{u,v\}\\ -\{u_{1},v_{1}\},-\{u_{2},v_{2}\},\dots,-\{u_{\ell},v_{\ell}\},&S_{i}=-\{u,v\}\\ \perp,\perp,\dots,\perp,&S_{i}=\perp.\end{cases}$
Here $u_{i},v_{i}$ are the copies of the original vertices $u,v$. Then we take the new stream $\tilde{S}:=S^{(1)}+\dots+S^{(T)}$ to be the concatenation of all partial streams.
Let $\tilde{r}=\mathcal{A}(\tilde{S})$ be the result of the algorithm on $\tilde{S}$ and output $r=(\nicefrac{{\tilde{r}_{\ell}}}{{\ell}},\nicefrac{{\tilde{r}_{2\ell}}}{{\ell}},\dots,\nicefrac{{\tilde{r}_{T\ell}}}{{\ell}})$ as the query answers to the original stream $T$.
By construction, the size of the maximum matching $\mu(\tilde{S}_{i\ell})$ in the new stream at time $i\ell,i\in[T]$ is exactly $\ell\cdot\mu(S_{i})$ in the original stream at time $i$. Thus by assumption, our output has additive error within
$\frac{\zeta(\varepsilon\cdot\ell n,\varepsilon\cdot\ell T)}{\varepsilon\ell}=O(\zeta(n,T))$
with probability $0.99$. Any neighboring streams from $S$ becomes $\ell$-neighboring streams from $\tilde{S}$ and thus by group privacy, our output is $1$-DP. ∎
Theorem 8.5.
Fix $\varepsilon\in(0,1)$. If $\mathcal{A}$ is an $\varepsilon$-DP mechanism that answers maximum matching queries on graphs with $n$ vertices in the continual release model within additive error $\zeta$ with probability at least $0.99$ for fully dynamic streams of length $T$, then
$\zeta=\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right).$
Moreover, we may assume the graph is bipartite, has maximum degree 2, and arboricity 1.
Proof (Theorem 8.5).
In the regime $T\geq\nicefrac{{1}}{{\varepsilon}}$, combining Lemma 8.4 and Theorem 8.3 implies a lower bound of
$\frac{1}{\varepsilon}\Omega(\min(\sqrt{\varepsilon n},(\varepsilon T)^{1/4}))=\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}}\right)\right).$
In the second regime $T<\nicefrac{{1}}{{\varepsilon}},\nicefrac{{n}}{{2}}$, we prove a lower bound of $\Omega(T)$. Suppose towards a contradiction that there is an $\varepsilon$-DP mechanism in this regime with additive error within $\zeta=\nicefrac{{T}}{{4}}$. Let $r,\tilde{r}$ be the outputs of $\mathcal{A}$ on the empty update stream of length $T$ and the stream that adds an edge of a fixed perfect matching at every time step, respectively. By the accuracy of $\mathcal{A}$, we have $\mathbf{P}[r_{T}\leq\nicefrac{{T}}{{4}}]\geq 0.99$. By $\varepsilon$-DP and group privacy, we have
$\mathbf{P}[\tilde{r}_{T}>\nicefrac{{T}}{{4}}]\leq e^{\varepsilon T}\mathbf{P}[r_{T}>\nicefrac{{T}}{{4}}]\leq 0.01e<0.99.$
But then since the true matching size in the second stream at time $T$ is $T$, $\mathcal{A}$ does not have additive error at most $\zeta=\nicefrac{{T}}{{4}}$ with probability at least $0.99$. By contradiction. it follows that we must have $\zeta=\Omega(T)$.
In the final regime $\nicefrac{{n}}{{2}}\leq T<\nicefrac{{1}}{{\varepsilon}}$, we show a lower bound of $\Omega(n)$. Simply apply our second argument with $\tilde{T}=\nicefrac{{n}}{{2}}$ to arrive at a lower bound of $\Omega(\tilde{T})=\Omega(n)$. ∎
8.3 Connected Components
Using a similar technique, we show a lower bound for the problem of privately estimating the number of connected components. In particular, we encode a single bit of a private database $y\in\{0,1\}^{n}$ using a subgraph on 4 vertices $u,v,a,b$. The edges $\{u,a\},\{v,b\}$ always exist but $\{a,b\}$ exists if and only if the bit is non-zero. Then notice that adding the edge $\{u,v\}$ decreases the number of connected components if and only if the bit is zero.
1 Function Alg(private dataset $y\in\{0,1\}^{n}$, public queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$, DP mechanism for connected components $\mathcal{M}$)
2       $V\leftarrow\{u_{i},v_{i},a_{i},b_{i}:i\in[n]\}$ ($4n$ vertices)
3       $E_{0}\leftarrow\varnothing$ empty edge set
4       $S^{(0)}\leftarrow$ empty update stream of length $3n$
5       $S^{(1)},\dots,S^{(k)}\leftarrow$ $k$ empty update streams each of length $2n$
6       for $i=1,2,\dots,n$ do
7             $S_{i}^{(0)}\leftarrow+\{u_{i},a_{i}\}$ (insert edge $\{u_{i},a_{i}\}$)
8             $S_{n+i}^{(0)}\leftarrow+\{v_{i},b_{i}\}$
9             if $y_{i}=1$ then
10                   $S_{2n+i}^{(0)}\leftarrow+\{a_{i},b_{i}\}$
11
12
13      for $j=1,2,\dots,k$ do
14             for $i=1,2,\dots,n$ do
15                   if $q^{(j)}_{i}=1$ then
16                         $S_{i}^{(j)}\leftarrow+\{u_{i},v_{i}\}$
17                         $S_{n+i}^{(j)}\leftarrow-\{u_{i},v_{i}\}$ (delete edge $\{u_{i},v_{i}\}$)
18
19
20
21      $S\leftarrow S^{(0)}+S^{(1)}+\dots+S^{(k)}$ concatenation of all streams
22       $r^{(0)},r^{(1)},\dots,r^{(k)}\leftarrow\mathcal{A}(S)$ answers to queries
23       for $j=1,2,\dots,k$ do
24             Output $\lVert q^{(j)}\rVert_{0}+r^{{(j)}}_{n}-r^{(0)}_{3n}$
25
26
Algorithm 8.2 Reduction from Inner Product Queries to Connected Components
Lemma 8.6.
Given an $\varepsilon$-DP mechanism for connected components in the continual release setting that outputs estimates within additive error at most $\zeta$ with probability $0.99$ for a fully-dynamic stream of length $T\geq 3n+2nk$, Algorithm 8.2 is an $\varepsilon$-DP mechanism for answering $k$ inner product queries within additive error $2\zeta$ with probability $0.99$.
Proof (Lemma 8.6).
Let $\mathcal{M}$ be such a $\varepsilon$-DP mechanism and suppose we are given a private dataset $y\in\{0,1\}^{n}$ as well as public linear queries $q^{(1)},\dots,q^{(k)}$. Let $S^{(0)},S^{(1)},\dots,S^{(k)}$ be the partial update streams as in Algorithm 8.2 and
$(r^{(0)},r^{(1)},\dots,r^{(k)})=\mathcal{M}(S^{(0)}+S^{(1)}+\dots+S^{(k)})$
be the corresponding stream of outputs when we run $\mathcal{M}$ on the concatenation of partial streams. By assumption, $(2n-r^{{(0)}}_{3n})$ is an estimate of $\lVert y\rVert_{0}$ within additive error $\zeta$ and $(2n-r^{(j)}_{n})$ is an estimate of $\lVert y\mid q^{(j)}\rVert_{0}$ within additive error $\zeta$. Hence the value of
$\lVert q^{(j)}\rVert_{0}+(2n-r^{{(0)}}_{3n})-(2n-r^{(j)}_{n})=\lVert q^{(j)}\rVert_{0}+r^{{(j)}}_{n}-r^{(0)}_{3n}$
is an estimate of $\langle y,q^{(j)}\rangle$ with additive error within $2\zeta$. ∎
Theorem 8.7.
If $\mathcal{A}$ is a $1$-DP mechanism that answers connected component queries on graphs with $n$ vertices in the continual release model within additive error $\zeta$ with probability at least $0.99$ for fully dynamic streams of length $T$, then
$\zeta=\Omega(\min(\sqrt{n},T^{1/4})).$
Moreover, we may assume the graph is bipartite, has maximum degree 2, and arboricity 2.
Proof (Theorem 8.7).
Let $c_{1},c_{2}$ be the constants from Theorem 8.1. Fix $n\in\mathbb{N}$ and suppose $T\geq 3n+2c_{1}n^{2}$. By Lemma 8.6, $\mathcal{A}$ implies a $1$-DP mechanism for answering $k=c_{1}n$ inner product queries within additive error $2\zeta$ with probability $0.99$. But then by Theorem 8.1, we must have $\zeta\geq c_{2}\sqrt{n}/2$.
Now consider the regime of $T<3n+2c_{1}n^{2}$. The same argument holds for $\tilde{n}=\Theta(\sqrt{T})$ such that $3\tilde{n}+2c_{1}\tilde{n}^{2}\leq T$ to yield a lower bound of $\zeta\geq c_{2}\sqrt{\tilde{n}}=\Omega(T^{1/4})$. ∎
By following an identical argument to Lemma 8.4, i.e. creating duplicate instances of the lower bound construction within a longer stream, we derive a similar lower bound to Theorem 8.5 that accounts for the privacy parameter $\varepsilon$.
Theorem 8.8.
Fix $\varepsilon\in(0,1)$. If $\mathcal{A}$ is an $\varepsilon$-DP mechanism that answers connected component queries on graphs with $n$ vertices in the continual release model within additive error $\zeta$ with probability at least $0.99$ for fully dynamic streams of length $T$, then
$\zeta=\Omega\left(\min\left(\sqrt{\frac{n}{\varepsilon}},\frac{T^{1/4}}{\varepsilon^{3/4}},n,T\right)\right).$
Moreover, we may assume the graph is bipartite, has maximum degree 2, and arboricity 2.
8.4 Further Graph Statistics
We remark that the underlying idea for the basic maximum matching and connected components lower bounds (Theorem 8.3, Theorem 8.7) is that we can encode the bits of a secret database $y\in\{0,1\}^{n}$ within the structure of a sparse graph. Privately answering an inner product query on this database then reduces to answering a “bitwise OR” query by the inclusion-exclusion principle. The general bound which accounts for the privacy parameter $\varepsilon$ (Theorem 8.5, Theorem 8.8) is obtained through a folklore reduction to the basic lower bound.
It is not hard to see that this technique can extend to $k$-edge-connected component queries, $k$-vertex-connected component queries, and triangle counting queries for sparse graphs.
9 Conclusion & Future Work
In this paper, we initiated the study of low-space continual release algorithms for general graph problems. Using techniques from the non-private graph sparsification literature, we provided continual release algorithms for a variety of general graph problems, achieving, for the first time, nearly the same space and approximation guarantees of their non-private streaming counterparts. The improved space bounds are especially relevant for enabling computations on massive datasets, which are the core motivation of the field of online streaming algorithms. In addition to our space gains, we provide the first bounded arboricity sparsifiers for node-DP algorithms in the continual release model. We hope such sparsifiers have implications in the static DP model.
For our upper bounds, we mostly focused on the insertion-only setting of continual release. As the area of fully-dynamic algorithms in the continual release model is largely unexplored, we believe that an interesting future research direction is closing the gap in our theoretical understanding of this model. As observed in prior work, and hinted by our hardness results, the fully dynamic setting is significantly harder in continual release, with even basic graph problems requiring $\tilde{\Omega}(\operatorname*{poly}(n))$ additive error for dynamic streams while admitting $\tilde{O}(\operatorname*{poly}(\log(n))/\varepsilon)$ additive error in the insertion-only case. In this context, it would be especially interesting to deepen our understanding of the interplay between the dynamicity of the continual release setting (insertion-only vs fully-dynamic) and the space lower bounds (as opposed to error lower bounds) imposed by privacy. This is an area that has only recently received attention [dinur2023differential] and is an interesting future direction to explore.
[…]
Appendix A Related Works
Our work is related to multiple areas of algorithms including private graph algorithms, private streaming algorithms, and related non-private variants. We now cover the most relevant work in each area.
Streaming Algorithms
The streaming model of computation is a prominent model for large-scale data analysis that has been studied for multiple decades [morris1978counting]. In this model, one usually seeks space- and time-efficient algorithms that can process the stream on-the-fly without the need to store all data. A long line of work in this area includes classical results such as the Flajolet-Martin algorithm for counting distinct elements [flajolet1985probabilistic] and many others [flajolet2007hyperloglog, alon1996space]. In the context of streaming graph algorithms, ideally one would like to obtain space sublinear in the number of vertices [mcgregor2018simple, HuangP19]; but for many graph problems it is necessary to work in the semi-streaming model settling on space near-linear in the number of vertices [assadi2022semi, AG11, Esfandiari2018, feigenbaum2005graph, FMU22, mcgregor2015densest]. Online streaming algorithms release graph statistics after every update while maintaining the low space bounds [chen2023sublinear, halldorsson2016streaming, cormode2018approximating, GS24].  In the non-private streaming setting, there exists many works on graph algorithms including approximating the size of the maximum matching [AG11, ALT21, Assadi22, FMU22, mcgregor2018simple], vertex cover [assadi2019coresets], densest subgraph [EHW16, mcgregor2015densest], $k$-core decomposition [Esfandiari2018, KingTY23], number of connected components [BerenbrinkKM14], and others [assadi2022semi, AG11, feigenbaum2005graph, FMU22, HuangP19, henzinger1998computing, muthukrishnan2005data, mcgregor2014graph]
Continual Release Model
In the context of streaming computation, the DP model of reference is the continual release model [DworkNPR10, CSS11] where we require algorithms to abide by a strong privacy notion: an observer obtaining all future outputs of the algorithm must in essence learn almost nothing about the existence of any single input. Since its introduction, this research area has received vast attention outside of graphs, including many recent works (see e.g. [ChanLSX12, henzinger2024unifying, henzinger2023differentially, fichtenberger2023constant, jain2023counting, JRSS21]).
Among the insertion-only continual release work, prior research has tackled classical estimation problems [CSS11, henzinger2024unifying, henzinger2023differentially], as well as heavy hitters-related problems [ChanLSX12, epasto2023differentially]. More recent works tackle the fully-dynamic continual release setting [jain2023counting, dupre2023differential, FichtenbergerHO21]. Particularly relevant is [FichtenbergerHO21] that shows hardness results for graph estimation in fully-dynamic streams. Our work contributes new hardness results in this new emerging area as well.
Most relevant to our paper is the literature on continual release algorithms in graphs [upadhyay2021differentially, FichtenbergerHO21, Song18, jain2024time]. In this area, [Song18] studied graph statistics (degree distribution, subgraph counts, etc.) on bounded degree graphs. [FichtenbergerHO21] focused on estimating a number of graph statistics like maximum matching, triangle counting, and the density of the densest subgraph for both edge and node privacy; they provide approximation guarantees in terms of the maximum degree in the graph as well as lower bounds in insertion-only and fully dynamic streams. [jain2024time] focuses on counting problems in graphs (counting edges, triangles, stars, connected components), for node-privacy and where privacy must hold for arbitrary graphs (e.g., graphs with arbitrary degrees). They give time-aware projection algorithms that can transform any continual release algorithm that gives approximation guarantees for bounded degree graphs into a truly private algorithm on nearly bounded degree graphs. Like in [jain2024time] for our algorithms that assume a public bound, say on the stream arboricity, this bound affects only the approximation guarantees but not the privacy claims, hence our algorithms satisfy their notion of truly private [jain2024time].
Private Graph Algorithms
The private literature on graph algorithms includes a large body of work on static graph algorithms (see e.g. [AU19, blocki2022privately, dinitz2024tight, DLRSSY22, ELRS22, kasiviswanathan2013analyzing, kalemaj2023node, LUZ24, mueller2022sok, NRS07, RS16, raskhodnikova2016differentially, Upadhyay13] and references therein). Aside from the problems we study, work in this area includes results on preserving graph cuts [AU19], the stochastic block model recovery [chen2023private], graph clustering [bun2021differentially, imola2023differentially], and many other areas.
Our paper focuses on the pure, $\varepsilon$-DP setting. In this setting, there has been a variety of recent works that we study for static graphs. The densest subgraph (DSG) problem has been extensively studied in the non-private streaming context [BHNT15, EHW16, mcgregor2015densest]. Our work builds on the results of [EHW16, mcgregor2015densest] which show that edge sampling approximately preserves the density of the densest subgraph. In the context of privacy, beyond the already cited work of [FichtenbergerHO21] that focused on density estimation only, all other works are in the static DP or LEDP setting [NV21, AHS21, DLRSSY22, dinitz2024tight, DLL23]. In the $\varepsilon$-DP setting, the best known lower bound on the additive error of the densest subgraph problem is $\Omega(\sqrt{\varepsilon^{-1}\log n})$ [AHS21, NV21]. Comparatively, the best known upper bounds in the central $\varepsilon$-DP setting are the $(2,O(\varepsilon^{-1}\log n))$-approximation [DLL23], $(2+\eta,O(\varepsilon^{-1}\log^{2}n))$-approximate [dinitz2024tight], and the $(1+\eta,O(\varepsilon^{-1}\log^{4}n))$-approximation [DLRSSY22] upper bounds. In the $\varepsilon$-LEDP setting, the best known upper bounds are the $(2+\eta,O(\eta^{-1}\log^{2}n))$-approximate [dinitz2024tight] and $O(2,O(\varepsilon^{-1}\log n))$-approximate [DLL23] bounds. Our paper presents the first continual release algorithm for releasing an actual approximate densest subgraph (as opposed to estimating its density) in edge-private insertion-only streams.
Related to the densest subgraph problem is the $k$-core decomposition of the graph for which some streaming results are known [Esfandiari2018, KingTY23, SGJ13]. In the private setting, this problem has been tackled by [DLRSSY22, DLL23, HSZ24] where the best error bound achieved in both the central $\varepsilon$-DP and $\varepsilon$-LEDP settings provide $(1,O(\varepsilon^{-1}\log n))$-approximations of the core number [DLL23]. This is tight against the recently shown lower bound of $\Omega(\varepsilon^{-1}\log n)$ [HSZ24]. In our paper, we provide the first continual release algorithm for approximating the core number of nodes in a graph.
To the best of our knowledge, the only differentially private algorithm that releases an implicit vertex cover is that of [gupta2010differentially] in the static $\varepsilon$-edge DP setting.
Appendix B The Size of a Maximum Matching
In this section, we recall the Micali-Vazirani algorithm [MV80] for finding a maximum matching in a graph. See also [vazirani2012simplification, vazirani1994theory] for a complete proof of correctness.
Theorem B.1 ([MV80]).
There is an algorithm that finds a maximum matching in a general graph $G=(V,E)$ in time $O(\lvert E\rvert\sqrt{\lvert V\rvert})$. Moreover, it uses $O(\lvert E\rvert)$ space.
The Micali-Vazirani algorithm repeatedly finds augmenting paths in each iteration. In each iteration, it finds a maximal set of disjoint minimum length augmenting paths with respect to the current matching and augments along all paths, both in $O(\lvert E\rvert)$ time. Then, one can argue that it suffices to terminate after $O(\sqrt{\lvert V\rvert})$ iterations [karzanov1973exact, HK71]. Each iteration can be implemented using $O(\lvert V\rvert+\lvert E\rvert)$ auxiliary space.
B.1 Coupled Node Sensitivity of the Greedy Matching Algorithm
Theorem B.2 (Section 1.4.2 in [blocki2022privately]).
The greedy matching algorithm $\textsc{GreedyMatching}_{\pi}$ that processes edges in the order specified by $\pi$ and outputs the size of a maximal matching has sensitivity 1 with respect to node neighboring graphs.
Appendix C Private Static Densest Subgraph
Theorem C.1 (Theorem 6.1 in [DLL23]).
Fix $\eta\in(0,1]$. There is an $\varepsilon$-edge DP densest subgraph algorithm that runs in polynomial time and $O(m+n)$ space and returns a subset $V^{\prime}\subseteq V$ of vertices that induces a $O\left(2,O(\varepsilon^{-1}\log n)\right)$-approximate densest subgraph with probability at least $1-O(1/\operatorname*{poly}(n))$.
Theorem C.2 (Theorem 5.1 in [DLRSSY22]).
Fix $\eta\in(0,1]$. There is an $\varepsilon$-edge DP densest subgraph algorithm that runs in $O((m+n)\log^{3}n)$ time and $O(m+n)$ space and returns a subset $V^{\prime}\subseteq V$ of vertices that induces a $O\left(1+\eta,O(\varepsilon^{-1}\log^{4}n)\right)$-approximate densest subgraph with probability at least $1-O(1/\operatorname*{poly}(n))$ for any constant $\eta>0$.
Appendix D The Density of the Densest Subgraph
We now recall the following theorem about computing the exact density of a densest subgraph in the static setting.
Theorem D.1.
Given an undirected graph $G=(V,E)$, there is an algorithm ExactDensity that computes the exact density of the densest subgraph. Moreover, the algorithm terminates in $O(\lvert E\rvert^{3}\log(\lvert V\rvert))$ time and uses $O(\lvert E\rvert)$ space.
We first describe the auxiliary flow graph describe in [boob2020flowless, chekuri2022densest]. Given an undirected graph $G=(V,E)$ and a value $\lambda>0$, we construct the following bipartite flow network $D=(N,A)$. We use the terminology of nodes and arcs to distinguish from the original graph $G$. $N$ consists of a source node $s$, a node $a_{e}$ for every edge $e\in E$, a node $a_{v}$ for every vertex $v\in V$, and a sink node $t$. There is a directed arc from $s$ to $a_{e}$ for every $e\in E$ with capacity 1, directed arcs from $a_{e}$ to $a_{v}$ and from $a_{e}$ to $a_{u}$ for every $e=\{u,v\}\in E$, both with infinite capacity, and a directed arc from $a_{v}$ to $t$ for every $v\in V$ with capacity $\lambda$.
For every $U\subseteq V$ in the original graph, the set of nodes
$N_{U}:=\{s\}\cup\{a_{uv}:u,v\in U\}\cup\{a_{v}:v\in U\}$
is an $(s,t)$-cut with capacity
$\lvert E\rvert-\lvert E[U]\rvert+\lambda\lvert U\rvert.$
Any minimum $(s,t)$-cut must be of this form (allowing for $U=\varnothing$). This leads to the following observation.
Proposition D.2 ([chekuri2022densest]).
Let $G=(V,E)$ be an undirected graph, $\lambda\geq\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$, and $D=(N,A)$ be the directed flow graph obtained from $G$. Let $N_{U}$ induce a minimum $(s,t)$-cut in $D$. Either the densest subgraph in $G$ has density at most $\lambda$ and $\delta(N_{U})$ has capacity $\lvert E\rvert$, or $G[U]$ has density strictly greater than $\lambda$ and $\delta(N_{U})$ has capacity strictly less than $\lvert E\rvert$.
Proof (Proposition D.2).
For $U=\varnothing$, the cut induced by $N_{\varnothing}=\{s\}$ has capacity $\lvert E\rvert\leq\lambda\lvert V\rvert$.
Let $U\subseteq V$. If the density of $G[U]$ is at most $\lambda$, then $\lvert E[U]\rvert\leq\lambda\lvert U\rvert$ so that the cut induced by $N_{U}$ has capacity at least $\lvert E\rvert$. If this holds for all $\varnothing\neq U\subseteq V$, then the cut induced by $N_{\varnothing}$ with capacity $\lvert E\rvert$ is a minimum cut and by the max-flow min-cut theorem, the maximum flow has value $\lvert E\rvert$.
Otherwise, assume that the density of $G[U^{*}]$ is strictly greater than $\lambda$. Then $\lvert E[U^{*}]\rvert>\lambda\lvert U^{*}\rvert$ so that the cut induced by $N_{U^{*}}$ has capacity strictly less than $\lvert E\rvert$. Another application of the max-flow min-cut theorem concludes the proof. ∎
Now we recall the classical result of [dinitz2006dinitz].
Theorem D.3 ([dinitz2006dinitz]).
Given directed graph $D=(N,A)$ with a source-sink pair $s,t\in N$ and non-negative arc capacities, there is an algorithm for computing an exact maximum $(s,t)$-flow. Moreover, the algorithm terminates in $O(\lvert N\rvert^{2}\lvert A\rvert)$ time while using $O(\lvert N\rvert+\lvert A\rvert)$ space.
We are now ready to show Theorem D.1.
Proof (Theorem D.1).
Without loss of generality, assume that $G$ is connected so that $\lvert E\rvert=\Omega(\lvert V\rvert)$.
Construct the auxiliary flow network $D=(N,A)$ above for some guess $\lambda\geq\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$ of the value of the densest subgraph. There are $2+\lvert V\rvert+\lvert E\rvert=O(\lvert E\rvert)$ nodes and $\lvert E\rvert+2\lvert E\rvert+\lvert V\rvert=O(\lvert E\rvert)$ arcs. We can check if the densest subgraph has value at most $\lambda$ by computing the maximum flow in the $D$ using Dinic’s algorithm Theorem D.3. This is justified by Proposition D.2. Moreover, each flow computation also retrieves the vertices inducing a minimum $(s,t)$-cut by computing the vertices reachable from $s$ in the residual graph. The time and space complexity of computing the residual flow is dominated by the time and space complexity of the flow algorithm. Thus each flow computation either declares that the densest subgraph has density “at most $\lambda$”, or produces some $U\subseteq V$ with density strictly more than $\lambda$.
The value of the densest subgraph is guaranteed to lie in the interval $[\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}},\nicefrac{{(\lvert V\rvert-1)}}{{2}}]\subseteq[\nicefrac{{1}}{{2}},\nicefrac{{\lvert V\rvert}}{{2}}]$, so we can binary search over $\lambda$ in $O(\log\nicefrac{{\lvert V\rvert}}{{\alpha}})$ time and constant space to produce an estimate of the optimal density within an additive error of $\alpha$. Note that the densest subgraph has density at least $\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$.
We first use the flow gadget to check if the densest subgraph has density strictly greater than $\nicefrac{{\lvert E\rvert}}{{\lvert V\rvert}}$. Thus at every iteration, we maintain a vertex set $U$ and and upper bound $h\geq 0$ such that the density $\rho(U)$ of $G[U]$ and the optimal density $\rho^{*}$ satisfies
$\rho(U)\leq\rho^{*}\leq h.$
Each iteration of the binary search takes time
$O(\lvert N\rvert^{2}\lvert A\rvert)=O(\lvert E\rvert^{3})$
Thus the total running time is
$O\left(\lvert E\rvert^{3}\log(\nicefrac{{\lvert V\rvert}}{{\alpha}})\right).$
Here $\alpha>0$ is the accepted level of additive error.
We remark that the density of a subgraph is a rational number whose numerator lies in $[\lvert E\rvert]$ and denominator in $[\lvert V\rvert]$. This means that the gap between the optimal density and the density of suboptimal subgraphs is at least $\nicefrac{{1}}{{\lvert V\rvert(\lvert V\rvert-1)}}\geq\nicefrac{{1}}{{\lvert V\rvert^{2}}}$. We can thus set $\alpha=\nicefrac{{1}}{{\lvert V\rvert^{2}}}$ so that the density of the maintained subgraph $U$ is optimal after termination. This brings the total running time to
$O\left(\lvert E\rvert^{3}\log(\lvert V\rvert^{3})\right)=O\left(\lvert E\rvert^{3}\log(\lvert V\rvert)\right)$
Faster max-flow or min-cut algorithms lead to better running time.
The auxiliary space is proportional to the size of the digraph and is thus
$O(\lvert N\rvert+\lvert A\rvert)=O(\lvert E\rvert).$
Appendix E Public Bound on Arboricity
Our edge-DP matching algorithm (Theorem 4.1), node-DP matching algorithms (Theorem 4.13, Theorem 4.14), and node-DP implicit vertex cover algorithm (Theorem 5.6) all require a public data-oblivious estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream. While our algorithms always guarantee privacy, their utility is only guaranteed when $\tilde{\alpha}\geq\alpha$. One way to remove this assumption is to first execute an insertion-only continual release algorithm for estimating the arboricity and then run the desired algorithm for a second pass given the estimated arboricity.
E.1 Notification of Failure
In the one-pass setting, the simple approach above fails since we require a bound on the maximum arboricity throughout the entire stream. However, it may still be desirable to notify the observer that the current arboricity of the graph exceeds the public bound and thus the utility is no longer guaranteed. For this purpose, we can run a continual release algorithm for estimating the arboricity alongside the desired algorithm. For the sake of simplicity, we explain how we can use our $k$-core algorithm (Theorem 7.1) to accomplish such a task.
Recall that the degeneracy of a graph is defined to be the least number $k$ such that every induced subgraph has a vertex of degree at most $k$. It is known that this quantity is equal to the value of the largest $k$-core, say denoted by $k_{\max}$. Using the definition of the arboricity $\alpha$ as the minimum number of forests that partition the edge set, we can show that
$\frac{1}{2}k_{\max}\leq\alpha\leq k_{\max}.$
Hence our $k$-core algorithm immediately yields a $(4+\eta,O(\varepsilon^{-1}\eta^{-2}\log^{3}(n)))$-approximation of the arboricity. When this estimated arboricity (approximately) exceeds the public bound, we can notify the observer that utility is no longer guaranteed by outputting “FAIL”. Note that algorithms with better approximation and space guarantees that directly estimate $\alpha$ most likely exist but we describe the approach here using Theorem 7.1 for the sake of simplicity.
E.2 Guessing the Arboricity
In the case of our matching algorithms (Section 4), we briefly sketch how to completely remove the dependence on a prior public $\tilde{\alpha}$ bound at the cost of additional space. Suppose we are given an $\varepsilon$-edge DP $(\beta,\zeta)$-approximation continual release algorithm $\mathcal{A}_{e}$ for computing the arboricity that uses $O(S)$ space (e.g. see previous section) or similarly an $\varepsilon$-node DP approximation algorithm $\mathcal{A}_{n}$. We can guess $\tilde{\alpha}=2^{p}$ for each $p=1,2,\dots,O(\log n)$ and execute $O(\log n)$ instances of our edge- or node-DP continual release matching algorithm (Theorem 4.1, Theorem 4.13) alongside $\mathcal{A}$. Thus we can identify the “correct” output from the parallel instances corresponding to the current arboricity $\alpha_{t}$ up to a $O\left(2\beta,O(\zeta)\right)$-approximation.
Plugging in our edge-DP matching algorithm (Theorem 4.1) and $\mathcal{A}_{e}$ yields an $\varepsilon$-edge DP continual release algorithm that returns a $\left(O(\beta\alpha_{t}+\zeta),O(\varepsilon^{-1}\operatorname*{poly}\log(n))\right)$-approximate estimate of the maximum matching size using $O(S+\varepsilon^{-1}\operatorname*{poly}\log(n))$ space. Here $\alpha_{t}$ is the true arboricity of the graph at time $t$. Similarly, implementing the above approach with $\mathcal{A}_{n}$ and our node-DP matching algorithm (Theorem 4.13) yields an $\varepsilon$-node DP continual release algorithm that returns a $\left(O(\beta\alpha_{t}+\zeta),O(\varepsilon^{-1}(\beta\alpha_{t}+\zeta)\operatorname*{poly}\log(n))\right)$-approximate estimate of the maximum matching size using $\tilde{O}(S+\varepsilon^{-1}n)$ space.
Appendix F Private Boosting of Randomized Algorithms via Generalized Select
The success probability of a DP algorithm can be privately boosted while losing privacy that scales linearly with the number of boosting iterations [gupta2010differentially]. We use a more sophisticated selection technique that enables us to lose only a constant factor of privacy [cohen2022generalized].
Let $\mathcal{M}_{i}:\mathcal{X}\to\mathbb{R}\times\mathcal{Y}$ for $i\in[\ell]$ be a finite collection of $\varepsilon$-DP mechanisms. Let $f:\mathcal{Y}\to\mathbb{R}$ be some objective function of the output. Suppose $\mathcal{M}_{i}(X)=(\tilde{f}(Y),Y)$ so that the first component is a noisy estimate of the objective value of the solution $Y$.
[…]
1 Input: privacy budget $\varepsilon$, parameter $\gamma$, repetitions $\tau$, $\ell$ $\varepsilon$-DP mechanisms $\mathcal{M}_{i}:\mathcal{X}\to\mathbb{R}\times\mathcal{Y}$, dataset $X$
2 Function PrivateSelection($\varepsilon$, $\gamma$, $\tau$, $\{\mathcal{M}_{i}\}_{i\in[\ell]}$, $X$)
3       Sample $p\in[0,1]$ according to cdf $\mathbf{P}[p\leq x]=x^{\gamma}$
4       $S\leftarrow\varnothing$ (set of candidate solutions)
5       for $i=1,\dots,\ell$ and $t=1,\dots,\tau$ do
6             Flip a coin $r_{i,t}\sim\operatorname{Ber}(p)$
7             if $r_{i,t}=1$ then
8                   $S\leftarrow S\cup\{\mathcal{M}_{i}(X)\}$
9
10
11      Output solution $(\tilde{f}(Y_{i,t}),Y_{i,t}))\leftarrow\min(S)$ ordered by $\tilde{f}(Y_{i,t})$
Algorithm F.1 PrivateSelection
Theorem F.1 (Theorem 1 in [cohen2022generalized]).
Suppose we execute Algorithm F.1
[…]
where each $\mathcal{M}_{i}:\mathcal{X}\to\mathbb{R}\times\mathcal{Y}$ is an $\varepsilon$-DP mechanism. Then the output is $(2+\gamma)\varepsilon$-DP.
We now describe a general result for boosting the success probability of a $\varepsilon$-DP algorithm
Theorem F.2.
Let $\mathcal{M}_{i}:\mathcal{X}\to\mathbb{R}\times\mathcal{Y}$ for $i\in[\ell]$ be a finite collection of $\varepsilon$-DP mechanisms, $\gamma>0$, and $n,\tau\in\mathbb{N}$. Let $f:\mathcal{Y}\to\mathbb{R}$ be some objective function of the output. Suppose $\mathcal{M}_{i}(X)=(\tilde{f}(Y),Y)$ so that the first component is a private estimate of the objective value of the solution $Y$.
Suppose the following holds.
  1. (a)
    There is some $i^{*}\in[\ell],B\in\mathbb{R},q\in(0,1)$ such that $\mathbf{P}[f(Y_{i^{*}})\leq B]\geq q$.
  2. (b)
    There is some increasing error function $\zeta:\mathbb{Z}_{\geq 0}\to\mathbb{R}$ such that $\mathbf{P}[\max_{i\in[\ell]}|\tilde{f}(Y_{i})-f(Y_{i})|\leq\zeta(n)]\geq 1-1/\operatorname*{poly}(n)$.
Then running PrivateSelection on the $(\mathcal{M}_{i},X)$’s with parameters
$\varepsilon,\gamma,\text{ and }\tau=\Theta\left(\frac{1}{q}n^{O(\nicefrac{{1}}{{\gamma}})}\log(n)\right)$
yields a $(2+\gamma)\varepsilon$-DP algorithm such that with probability at least $1-1/\operatorname*{poly}(n)$, the output solution $\bar{Y}$ satisfies
$f(\bar{Y})\leq B+2\zeta\left(\tau n\right).$
Moreover, we call each mechanism $\mathcal{M}_{i}$ at most $\tau$ times.
Proof.
The privacy guarantee comes directly from Theorem F.1. We need only show the correctness.
Let $i^{*}$ be as in Item a. The algorithm runs $\mathcal{M}_{i^{*}}$ for $T\sim\operatorname{Bin}(\tau,p)$ times where $\Pr[p\leq x]\leq x^{\gamma}$. First, we have $p>n^{-\Theta(\nicefrac{{1}}{{\alpha}})}$ with probability at least $1-1/\operatorname*{poly}(n)$ by definition. Conditioning on a fixed $p$ satisfying the above, we have $\mathop{{\mathbb{E}}}[T]=\tau p$ and an application of the multiplicative Chernoff bound (Theorem 3.1) implies that $T\geq\Omega(\tau p)$ with probability at least $1-\exp(-\Omega(\tau p))$. By choosing
$\tau=\Theta(q^{-1}p^{-1}\log(n))=\Theta(q^{-1}n^{\Theta(\nicefrac{{1}}{{\gamma}})}\log(n)),$
the probability above is at most $1-1/\operatorname*{poly}(n)$. Finally, running $M_{i^{*}}$ at least $\Theta(\tau p)=\Theta(q^{-1}\log(n))$ times means there is at least one solution $Y^{*}$ collected (but not necessarily output) by PrivateSelection satisfying
$f(Y^{*})\leq B.$
By Item b and a union bound, the maximum estimation error of every execution of some mechanism is at most
$\zeta=\zeta\left(\tau n\right)$
with probability at least $1-(1/\operatorname*{poly}(\tau n))\cdot\tau\geq 1-1/\operatorname*{poly}(n)$.
Let $\bar{Y}$ denote the final selected solution. Conditioning on the success of all three events, we have
$\tilde{f}(\bar{Y})$ $\leq\tilde{f}(Y^{*})\leq B+\zeta$ (135)
$f(\bar{Y})$ $\leq\tilde{f}(\bar{Y})+\zeta\leq B+2\zeta.$ (136)
Appendix G Missing Proofs from Section 7
In this section, we prove the approximation bound for Algorithm 7.1 via a slightly modified proof of [DLRSSY22, Theorem 4.7] using Lemma 7.3 and Lemma 7.4.
Lemma G.1 (Folklore).
Suppose we perform the following peeling procedure. Provided an input graph $G=(V,E)$, we iteratively remove (peel) nodes with degree $\leq d^{*}$ for some $d^{*}>0$ until no nodes with degree $\leq d^{*}$ remain. Then, $k(i)\leq d^{*}$ for each removed node $v$ and all remaining nodes $w$ (not peeled) have core number $k(w)>d^{*}$.
Lemma G.2.
Algorithm 7.1 returns a $\left(2+\eta,O\left(\varepsilon^{-1}\log^{3}n\right)\right)$-approximate $k$-core decomposition.
Proof.
Our proof follows almost verbatim from [DLRSSY22, Theorem 4.7] except we use Lemma 7.3 and Lemma 7.4 instead. Since we return the same core estimate as the equivalent vertex in $\mathcal{K}$, we only need to show the approximation returned from $\mathcal{K}$ gives the correct approximation. In the below proof, a group $g$ is defined as the $g$-th data structure in $\mathcal{K}$; hence the topmost level in the $g$-th group is the topmost level in the $g$-th level data structure in $\mathcal{K}$. We let $Z_{\ell}$ denote the set of nodes in levels $\geq\ell$ in any group $g$.
In this proof, when we refer to the level of a node $i$ in subgraph $j$, we mean the level of any vertex $i\in V$ in $\mathcal{K}$. Using notation and definitions from [DLRSSY22], let $\hat{k}(i)$ be the core number estimate of $i$ in $\mathcal{K}$ and $k(i)$ be the core number of $i$. First, we show that
$\text{if }\hat{k}(i)\leq(2+\eta)(1+\eta)^{g^{\prime}},\text{ then}k(i)\leq(1+\eta)^{g^{\prime}+1}+\frac{d\log^{3}n}{\varepsilon}$ (137)
with probability at least $1-\frac{1}{n^{c-1}}$ for any group $g^{\prime}\geq 0$, for any constant $c\geq 2$, and for appropriately large constant $d>0$. Recall $F-1$ is the topmost level of group $g^{\prime}$. In order for $(2+\lambda)(1+\eta)^{g^{\prime}}$ to be the estimate of $i$’s core number, $i$ must be in the topmost level of $g^{\prime}$ but below the topmost levels of all $g^{\prime\prime}>g^{\prime}$. Suppose we consider $i$’s level in group $g^{\prime}+1$ and let $\ell$ be $i$’s level in $g^{\prime}+1$. By Lemma 7.3, if $\ell<F-1$, then $\deg_{\mathcal{K}}(\ell,g^{\prime},i)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. Furthermore, each node $w$ at the same or lower level $\ell^{\prime}\leq\ell$ has $\deg_{\mathcal{K}}(\ell^{\prime},g^{\prime},w)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$, also by Lemma 7.3.
Suppose we perform the following iterative procedure: starting from level $level=0$ in $g^{\prime}$, remove all nodes in level $level$ during this turn and set $level\leftarrow level+1$ for the next turn. Using this procedure, the nodes in level 00 are removed in the first turn, the nodes in level $1$ are removed in the second turn, and so on until the graph is empty. Let $d_{level}(i)$ be the induced degree of any node $i$ after the removal in the $(level-1)$-st turn and prior to the removal in the $level$-th turn. Since we showed that $\deg_{\mathcal{K}}(level,g^{\prime},i)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ for any node $i$ at level $level<F-1$, node $i$ on level $level<F-1$ during the $level$-th turn has $d_{level}(v)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$. Thus, when $i$ is removed in the $level$-th turn, it has degree $\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$. Since all nodes removed before $i$ also had degree $\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ when they were removed, by Lemma G.1, node $i$ has core number $k(i)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ with probability $\geq 1-\frac{1}{n^{c}}$ for any $c\geq 1$. By the union bound over all nodes, this proves that $k(i)\leq(1+\eta)^{g^{\prime}}+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ for all $i\in[n]$ with $\hat{k}(i)\leq(2+\eta)(1+\eta)^{g^{\prime}}$ for all constants $c\geq 2$ with probability at least $1-\frac{1}{n^{c-1}}$.
Now we prove our lower bound on $\hat{k}(i)$. We prove that for any $g^{\prime}\geq 0$,
$\text{if }\hat{k}(i)\geq(1+\eta)^{g^{\prime}},\text{ then }k(i)\geq\frac{(1+\eta)^{g^{\prime}-2}-O\left(\frac{\log^{3}n}{\varepsilon}\right)}{{(2+\eta)(1+\eta)^{2}}}$ (138)
for all nodes $i$ in the graph with probability at least $1-\frac{1}{n^{c-1}}$ for any constant $c\geq 2$. We assume for contradiction that with probability $>\frac{1}{n^{c-1}}$ there exists a node $i$ where $\hat{k}(i)\geq(1+\eta)^{g^{\prime}}$ and $k(i)<\frac{(1+\eta)^{g^{\prime}-2}-O\left(\frac{\log^{3}n}{\varepsilon}\right)}{{(2+\eta)(1+\eta)^{2}}}$. This, in turn, implies that $k(i)\geq\frac{(1+\eta)^{g^{\prime}-2}-O\left(\frac{\log^{3}n}{\varepsilon}\right)}{{(2+\eta)(1+\eta)^{2}}}$ for all $i\in[n]$ where $\hat{k}(i)\geq(1+\eta)^{g^{\prime}}$ holds with probability $<1-\frac{1}{n^{c-1}}$. To consider this case, we use the pruning process defined in Lemma G.1. In the below proof, let $d_{S}(i)$ denote the induced degree of node $i$ in the subgraph induced by nodes in $S$. For a given subgraph $S$, we prune $S$ by repeatedly removing all nodes $i\in S$ whose $d_{S}(i)<\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$ for appropriately selected constant $a>0$. We consider levels from the same group $g^{\prime}$ since levels in groups lower than $g^{\prime}$ will also have smaller upper bound cutoffs, leading to an easier proof. Let $j$ be the number of levels below level $F-1$. We prove via induction that the number of nodes pruned from the subgraph induced by $Z_{F-1-j}$ must be at least
$\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{j-1}\left(\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\left(1-\frac{1}{(2+\eta)(1+\eta)^{2}}\right)\right).$ (139)
We first prove the base case when $j=1$. In this case, we know that for any node $i$ in level $F-1$ in group $g^{\prime}$, it holds that $\deg_{\mathcal{K}}(F-1,g^{\prime},i)\geq(1+\eta)^{g^{\prime}-2}-O\left(\frac{\log^{3}n}{\varepsilon}\right)$ with probability $\geq 1-\frac{1}{n^{c}}$ for any $n\geq 1$ by Lemma 7.4. Taking the union bound over all nodes in level $F-1$ shows that this bound holds for all nodes $i\in[n]$ with probability at least $1-\frac{1}{n^{c-1}}$. All below expressions hold with probability at least $1-\frac{1}{n^{c-1}}$; for simplicity, we omit this phrase when giving these expressions. In order to prune $i$ from the graph, we must prune at least
$\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)-\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$
$=\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\cdot\left(1-\frac{1}{{(2+\eta)(1+\eta)^{2}}}\right).$
neighbors of $i$ from $Z_{F-2}$. We must prune at least this many neighbors in order to reduce the degree of $i$ to below the cutoff for pruning a vertex (as we show more formally below). In the case when $(1+\eta)^{g^{\prime}}\leq 4(1+\eta)^{2}(\frac{a\log^{3}n}{\varepsilon})$,1919194444 is an arbitrarily chosen large enough constant. then our original approximation statement in the lemma is trivially satisfied because the core number is always non-negative and so even if $k(i)=0$, this is still within our additive approximation bounds. Hence, we only need to prove the case when $(1+\eta)^{g^{\prime}}>4(1+\eta)^{2}\left(\frac{a\log^{3}n}{\varepsilon}\right)$.
Then, if fewer than $\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\left(1-\frac{1}{(2+\eta)(1+\eta)^{2}}\right)$ neighbors of $i$ are pruned from the graph, then $i$ is not pruned from the graph. If $i$ is not pruned from the graph, then $i$ is part of a $\left(\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}\right)$-core (by Lemma G.1) and $k(i)\geq\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$ for all $i\in[n]$ where $\hat{k}(i)\geq(1+\eta)^{g^{\prime}}$ with probability $\geq 1-\frac{1}{n^{c-1}}$, a contradiction with our assumption. Thus, it must be the case that there exists at least one $i$ where at least $\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\left(1-\frac{1}{(2+\eta)(1+\eta)^{2}}\right)$ neighbors of $i$ are pruned in $Z_{T(g^{\prime})-1}$. For our induction hypothesis, we assume that at least the number of nodes as indicated in Equation 139 is pruned for $j$ and prove this for $j+1$.
Each node $w$ in levels $F-1-j$ and above has $d_{Z_{F-2-j}}(w)\geq(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}$ by Lemma 7.4 (recall that all $j$ levels below $F-1$ are in group $g^{\prime}$). For simplicity of expression, we denote $J\triangleq\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\left(1-\frac{1}{(2+\eta)(1+\eta)^{2}}\right)$. Then, in order to prune the $\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{j-1}J$ nodes by our induction hypothesis, we must prune at least
$\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{j-1}J\cdot\left(\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{2}\right)$ (140)
edges where we “charge” the edge to the endpoint that is pruned last. (Note that we actually need to prune at least $\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)\cdot\left(1-\frac{1}{{(2+\eta)(1+\eta)^{2}}}\right)$ edges per pruned node as in the base case but $\frac{\left((1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}\right)}{2}$ lower bounds this amount.) Each pruned node prunes less than $\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$ edges. Thus, using Equation 140, the number of nodes that must be pruned from $Z_{F-2-j}$ is
$\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{j-1}J\cdot\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{2\left(\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}\right)}=\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{j}J.$ (141)
Equation 141 proves our induction step. Using Equation 139, the number of nodes that must be pruned from $Z_{F-1-2\log_{{(2+\eta)(1+\eta)^{2}}/2}\left(n\right)}$ is greater than $n$ since $J\geq 1$ by our assumption that $(1+\eta)^{g^{\prime}-2}>4\left(\frac{a\log^{3}n}{\varepsilon}\right)$:
$\left(\frac{{(2+\eta)(1+\eta)^{2}}}{2}\right)^{2\log_{{(2+\eta)(1+\eta)^{2}}/2}\left(n\right)}\cdot J\geq n^{2}.$ (142)
Thus, at $j=2\log_{{(2+\eta)(1+\eta)^{2}}/2}\left(n\right)$, we run out of nodes to prune. We have reached a contradiction as we require pruning greater than $n$ nodes with probability at least $\frac{1}{n^{c-1}}\cdot\left(1-\frac{1}{n^{c-1}}\right)>0$ via the union bound over all nodes where $\hat{k}(i)\geq(1+\eta)^{g^{\prime}}$ and using our assumption that with probability $>\frac{1}{n^{c-1}}$ there exists an $i\in[n]$ where $k(i)<\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$. This contradicts with the fact that more than $n$ nodes can be pruned with 00 probability.
From Equation 137, we can first obtain the inequality $k(i)\leq\hat{k}(i)+\frac{d\log^{3}n}{\varepsilon}$ since this bounds the case when $\hat{k}(i)=(2+\eta)(1+\eta)^{g^{\prime}}$; if $\hat{k}(i)<(2+\lambda)(1+\eta)^{g^{\prime}}$ then the largest possible value for $\hat{k}(i)$ is $(2+\eta)(1+\eta)^{g^{\prime}-1}$ by our algorithm and we can obtain the tighter bound of $k(i)\leq(1+\eta)^{g^{\prime}}+\frac{d\log^{3}n}{\varepsilon}$. We can substitute $\hat{k}(i)=(2+\eta)(1+\eta)^{g^{\prime}}$ since $(1+\eta)^{g^{\prime}+1}<(2+\eta)(1+\eta)^{g^{\prime}}$ for all $\eta\in(0,1)$ and $\eta>0$. Second, from Equation 138, for any estimate $(2+\eta)(1+\eta)^{g}$, the largest $g^{\prime}$ for which this estimate has $(1+\eta)^{g^{\prime}}$ as a lower bound is $g^{\prime}=g+\lfloor\log_{(1+\eta)}(2+\eta)\rfloor\geq g+\log_{(1+\eta)}(2+\eta)-1$. Substituting this $g^{\prime}$ into $\frac{(1+\eta)^{g^{\prime}-2}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}$ results in
$\frac{(1+\eta)^{g+\log_{(1+\eta)}(2+\lambda)-3}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}=\frac{\frac{(2+\lambda)(1+\eta)^{g}}{(1+\eta)^{3}}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}=\frac{\frac{\hat{k}(i)}{(1+\eta)^{3}}-\frac{a\log^{3}n}{\varepsilon}}{{(2+\eta)(1+\eta)^{2}}}.$
Thus, we can solve $k(i)\geq\frac{\frac{\hat{k}(i)}{(1+\eta)^{3}}-\frac{a\log^{3}n}{\varepsilon}}{(2+\lambda)(1+\eta)^{4}}$ and $k(i)\leq\hat{k}(i)+\frac{d\log^{3}n}{\varepsilon}$ to obtain $k(i)-\frac{d\log^{3}n}{\varepsilon}\leq\hat{k}(i)\leq((2+\eta)(1+\eta)^{5})k(i)+\frac{a(2+\eta)(1+\eta)^{5}\log^{3}n}{\varepsilon}$. Simplifying, we obtain
$k(i)-$ $O(\varepsilon^{-1}\log^{3}n)\leq\hat{k}(i)\leq(2+\eta)(1+\eta)^{5}k(i)+O(\eta^{-1}\log n)$
which is consistent with the definition of a $(2+\eta^{\prime},O(\varepsilon^{-1}\log n))$-factor approximation algorithm for core number for any constant $\eta^{\prime}>0$ and appropriately chosen constant $\eta^{\prime}>\eta\in(0,1)$ that depend on $\eta^{\prime}$. ∎